### v1

# Characterizations of amenability for noncommutative dynamical systems and Fell bundles

Abstract.
We prove that for a locally compact group $G$, a $\mathrm{C}^{*}$-dynamical system $(A,G,\alpha)$ is amenable in the sense of Anantharaman-Delaroche if and only if, for every other system $(B,G,\beta)$, the diagonal system $(A\otimes_{\max}B,G,\alpha\otimes^{d}_{\max}\beta)$ has the weak containment property (wcp).
For Fell bundles over $G$, we construct a diagonal tensor product $\otimes^{d}_{\max}$ and show that a Fell bundle $\mathcal{A}$ has the positive approximation property of Exel and Ng (AP) precisely when $\mathcal{A}\otimes^{d}_{\max}\mathcal{B}$ has the wcp for every Fell bundle $\mathcal{B}$ over $G$. Equivalently, $\mathcal{A}$ has the AP if and only if the natural action of $G$ on the $\mathrm{C}^{*}$-algebra of kernels of $\mathcal{A}$ is amenable.
We show that the approximation properties introduced by Abadie and by Bédos–Conti are equivalent to the AP. We also study the permanence of the wcp, the AP, and the nuclearity of cross-sectional $\mathrm{C}^{*}$-algebras under restrictions, quotients, and other constructions.
Our results extend and unify previous characterizations of amenability for $\mathrm{C}^{*}$-dynamical systems and Fell bundles, and provide new tools to analyze structural properties of associated $\mathrm{C}^{*}$-algebras.
1. Introduction
Amenability is a central notion connecting group theory, operator algebras and noncommutative dynamics. Since the foundational work of Anantharaman-Delaroche on amenable actions of locally compact groups on $\mathrm{C}^{*}$- and $\mathrm{W}^{*}$-algebras [7, 8], a rich theory has developed relating amenability to approximation properties, tensorial behaviour and structural properties of crossed products. The goal of this paper is to provide new, unified characterizations of amenability both for $\mathrm{C}^{*}$-dynamical systems and for Fell bundles over locally compact groups, and to exploit those characterizations to obtain permanence and nuclearity results.
Our strategy is twofold. First, we give a tensorial (“diagonal”) criterion for amenability of actions: amenability of a system can be tested by forming diagonal maximal tensor products with arbitrary test-systems and checking the weak containment property (wcp). Second, we extend this viewpoint to Fell bundles by constructing a natural diagonal maximal tensor product for bundles and showing that Exel–Ng’s positive approximation property (AP) is equivalent to a diagonal tensorial wcp characterization and to amenability of a canonical kernel action. Along the way we compare several approximation properties appearing in the literature, prove the equivalence between Bédos–Conti’s approximation property (BCAP) and Exel–Ng’s AP, and establish permanence results (restrictions to subgroups and quotients to normal subgroups) as well as implications for nuclearity of cross-sectional algebras.
Below we state the main contributions in a compact, theorem-style form so the reader can immediately see the scope of the paper. After the statements we provide comments that highlight important subtleties and limitations (in particular concerning reductions to subgroups and quotient bundles) which are treated carefully in Section 5.
Theorem A.
Let $G$ be a locally compact group and let $\alpha$ be a $\mathrm{C}^{*}$-action of $G$ on a $\mathrm{C}^{*}$-algebra $A$. Using the universal unitary implementation $(A\subset\mathbb{B}(X_{\alpha}),U^{\alpha})$, denote by $A^{\prime}_{\alpha}$ the commutant of $A$ in $\mathbb{B}(X_{\alpha})$ and let $\beta$ be the continuous part of $\operatorname{Ad}(U^{\alpha})|_{A^{\prime}_{\alpha}}$. The following are equivalent:
  1. (1)
    $\alpha$ is amenable in the sense of Anantharaman-Delaroche (equivalently, the enveloping $\mathrm{W}^{*}$-action $\alpha^{\prime\prime}$ on $A^{\prime\prime}_{\alpha}$ is $\mathrm{W}^{*}$-amenable).
  2. (2)
    For every $\mathrm{C}^{*}$-action $\gamma$ of $G$, the diagonal maximal tensor action $\alpha\otimes_{\max}^{d}\gamma$ has the weak containment property (wcp), i.e. the canonical quotient
    $(A\otimes_{\max}B)\rtimes_{\alpha\otimes^{d}_{\max}\gamma}G\twoheadrightarrow(A\otimes_{\max}B)\rtimes_{{\mathop{\rm r}}\,(\alpha\otimes^{d}_{\max}\gamma)}G$
    is an isomorphism.
  3. (3)
    The diagonal action $\alpha\otimes_{\max}^{d}\beta$ has the wcp for the specific canonical test action $\beta$ defined above.
Theorem A gives a practical tensorial test for amenability: (2) is a full tensorial characterization while (3) shows that a single, canonical test-action suffices. Our proof removes the need to assume exactness of $G$ (previously required in some formulations) by exploiting a careful analysis of the universal unitary implementation and a von Neumann tensoring technique for equivariant completely positive maps (see Appendix A).
Theorem B.
Let $\mathcal{A}$ be a Fell bundle over a locally compact group $G$. There is a natural diagonal maximal tensor product $\otimes_{\max}^{d}$ for Fell bundles over the same group. With this notation, the following are equivalent:
  1. (1)
    $\mathcal{A}$ has Exel–Ng’s positive approximation property (AP).
  2. (2)
    For every Fell bundle $\mathcal{B}$ over $G$, the diagonal Fell bundle $\mathcal{A}\otimes_{\max}^{d}\mathcal{B}$ has the weak containment property (i.e. its full and reduced cross-sectional $\mathrm{C}^{*}$-algebras coincide).
  3. (3)
    The canonical action of $G$ on the $\mathrm{C}^{*}$-algebra of kernels of $\mathcal{A}$ (as in [3]) is amenable in the sense of Anantharaman-Delaroche.
Theorem B unifies Exel–Ng’s AP with a dynamical notion of amenability for the kernel action and extends the diagonal tensor viewpoint from actions to general Fell bundles.
Theorem C.
The following permanence and nuclearity statements hold for actions and Fell bundles over a locally compact group $G$:
  1. (1)
    (BCAP equivalence) The approximation property introduced by Bédos–Conti (BCAP), suitably extended to Fell bundles over locally compact groups, is equivalent to Exel–Ng’s AP (no nuclearity hypothesis on the unit fiber is required). This establishes BCAP as a conceptually and technically robust characterization of amenability for Fell bundles.
  2. (2)
    (Restriction to subgroups) If $\mathcal{A}$ has the AP and $H\leq G$ is a closed subgroup, then the restriction $\mathcal{A}|_{H}$ has the AP.
  3. (3)
    (Normal subgroup quotients) If $H\trianglelefteq G$ is a normal closed subgroup and ${\mathcal{B}}_{/H}$ denotes the quotient Fell bundle over $G/H$ constructed from $\mathcal{A}$ (see remarks below), then ${\mathcal{B}}_{/H}$ has the AP whenever $\mathcal{A}$ has the AP.
  4. (4)
    (Nuclearity interplay) If $A$ is nuclear and $\alpha$ is $\mathrm{C}^{*}$-amenable, then $A\rtimes_{{\mathop{\rm r}},\alpha}G$ is nuclear. Conversely, if $G$ is a SIN group (hence inner amenable) and $A\rtimes_{{\mathop{\rm r}},\alpha}G$ is nuclear, then $\alpha$ is amenable. A similar result holds for Fell bundles.
Item (1) of Theorem C deserves special emphasis: Bédos and Conti introduced BCAP in the discrete group setting and proved that AP implies BCAP (and the converse under nuclearity of the unit fiber). Our contribution is to extend BCAP to the locally compact setting and to show, without imposing nuclearity on the unit fiber, that BCAP and AP are equivalent. This equivalence both clarifies the landscape of approximation properties and provides a powerful, alternate toolkit to detect amenability in examples.
Remarks on reductions and quotient bundles
A recurring theme in Section 5 is the relationship between $\mathrm{C}^{*}$-amenability (i.e. AP) of a Fell bundle $\mathcal{B}$ over $G$ and the $\mathrm{C}^{*}$-amenability of its restriction $\mathcal{B}_{H}$ to a closed subgroup $H$ and of the bundle obtained by forming the (partial) quotient over $G/H$. The results in Section 5 show the following facts, proved there in full detail:
  • If $\mathcal{B}$ is $\mathrm{C}^{*}$-amenable then $\mathcal{B}_{H}$ is $\mathrm{C}^{*}$-amenable for every closed subgroup $H$ (Corollary 4.18).
  • When $H$ is open (or more generally, when its normalizer in $G$ is open) we use results from [17] about the existence of conditional expectations $C^{*}(\mathcal{B})\to C^{*}(\mathcal{B}_{H})$ and $C^{*}_{\mathop{\rm r}}(\mathcal{B})\to C^{*}_{\mathop{\rm r}}(\mathcal{B}_{H})$, so nuclearity and the wcp pass from $\mathcal{B}$ to $\mathcal{B}_{H}$ in that case.
  • If $H$ is normal we construct full and reduced completions of the partial cross-sectional bundle over $G/H$, denoted ${\mathcal{B}}_{/H}$ and ${\mathcal{B}}^{\mathop{\rm r}}_{/H}$ respectively, and prove isomorphisms $C^{*}(\mathcal{B})\cong C^{*}({\mathcal{B}}_{/H})$ and $C^{*}_{\mathop{\rm r}}(\mathcal{B})\cong C^{*}_{\mathop{\rm r}}({\mathcal{B}}^{\mathop{\rm r}}_{/H})$; moreover, under the hypothesis that $\mathcal{B}_{H}$ has the wcp the two completions coincide.
These constructions allow us to deduce many permanence results, but they also uncover a nuanced limitation: while the implication
 $\text{$\mathcal{B}$ is $\mathrm{C}^{*}$-amenable}\quad\Longrightarrow\quad\text{$\mathcal{B}_{H}$ and }{\mathcal{B}}_{/H}\ \text{are $\mathrm{C}^{*}$-amenable}$
holds (in fact the forward direction is proved generally in the paper), the converse
$\text{$\mathcal{B}_{H}$ and }{\mathcal{B}}_{/H}\ \text{are $\mathrm{C}^{*}$-amenable}\quad\Longrightarrow\quad\text{$\mathcal{B}$ is $\mathrm{C}^{*}$-amenable}$
is not established in full generality in this article. We prove this converse under additional hypotheses (for instance, when $G$ is inner amenable and the unit fibre $B_{e}$ is nuclear; see Corollary 5.8), and we know the converse holds for discrete $G$ (this will be treated in detail in a forthcoming paper). Thus, Section 5 contains both positive permanence results and a careful account of what is currently known about converses and the hypotheses required for them.
Methods and structure of the paper
The proofs combine three main ingredients:
  • an analysis of the universal unitary implementation of an action and its commutant (used to reduce amenability questions to equivariant conditional expectations);
  • a construction of diagonal maximal tensor products for Fell bundles based on Abadie’s maximal tensor product of bundles; and
  • a technical toolkit to form tensor products of equivariant completely positive maps between von Neumann algebras (developed in Appendix A).
The paper is organized as follows. Section 2 contains preliminaries on $\mathrm{C}^{*}$- and $\mathrm{W}^{*}$-actions and the proof of Theorem A. Section 3 constructs $\otimes_{\max}^{d}$ for Fell bundles and proves Theorem B. Section 4 establishes equivalence of BCAP and AP and collects permanence results, including those listed in Theorem C. Section 5 investigates restrictions, quotients and other permanence properties; here we give the most complete proofs of the forward implications and indicate precisely the extra hypotheses needed for the known converses. Two appendices contain auxiliary technical lemmas used throughout the paper.
Acknowledgements
We dedicate this work to the memory of Fernando Abadie. The second author warmly thanks the Departamento de Ciências Físicas e Matemáticas at Universidade Federal de Santa Catarina for hospitality and support.
2. Amenability of dynamical systems
2.1. Preliminaries
We start by presenting our notation and some basic facts. Some results are stated and proved in the Appendix.
In this article, we use group as an abbreviation of locally compact and Hausdorff group. A unitary representation $U$ of a group $G$ on (a Hilbert space) $X$ is a group homomorphism $t\mapsto U_{t}$ from $G$ to the unitary operators of $X$ such that, for all $x,y\in X,$ $(t\in G)\mapsto\langle U_{t}x,y\rangle$ is continuous. We convention that all of our inner products are linear in the second variable.
For us, a von Neumann algebra (or $\mathrm{W}^{*}$-algebra) is a $\mathrm{C}^{*}$-algebra that is linearly and isometrically isomorphic to the dual of a Banach space. Faithful normal representations need not be nondegenerate (i.e. unital), and even so we use them to construct faithful representations of the von Neumann tensor product $\bar{\otimes}$. By definition, $\mathrm{W}^{*}$-subalgebra means a (unital) $\operatorname{\rm w}^{*}$}-closed -subalgebra. The minimal and maximal tensor products of $\mathrm{C}^{*}$-algebras will be denoted $\otimes$} and $\otimes_{\max},$ respectively. The multiplier algebra and the center of a $\mathrm{C}^{*}$-algebra $A$ will be denoted $\mathcal{M}(A)$ and $Z(A)$, respectively.
When we say that $\gamma$ is a $\mathrm{W}^{*}$-action (respectively, a $\mathrm{C}^{*}$-action) of a group $G$ on a $\mathrm{W}^{*}$-algebra ($\mathrm{C}^{*}$-algebra) $M$, we mean that $\gamma\colon G\times M\to M$, $(t,m)\mapsto\gamma_{t}(m)$, is a set-theoretic action such that each $\gamma_{t}$ is a -isomorphism and, for all $m\in M$, the map $t\mapsto\gamma_{t}(m)$ is $\operatorname{\rm w}^{*}$}-continuous (norm continuous). A covariant representation of $\gamma$ is a triple $(X,\pi,U)$ where $\pi\colon M\to\mathbb{B}(X)$ is a -representation, $U$ is a unitary representation of $G$ on $X$, and for all $t\in G$ and $m\in M$, we have $\pi(\gamma_{t}(m))=U_{t}\pi(m)U_{t}^{*}=:\operatorname{Ad}(U)_{t}(\pi(m))$. If we qualify a covariant representation with a property (e.g., normal, unital, nondegenerate), we mean that the representation $\pi$ has that property. A unitary implementation of the $\mathrm{W}^{*}$-action (respectively, $\mathrm{C}^{*}$-action) $\gamma$ of $G$ on $M$ is a covariant representation $(X,\pi,U)$ of $\gamma$ that is faithful, unital, and normal (respectively, faithful and nondegenerate). In this case, we say that $(M\subseteq\mathbb{B}(X),U)$ is a unitary implementation of $\gamma$, and we may omit explicit reference to $\pi$ when no confusion can arise. A triple $(G,A,\alpha)$ is called a $\mathrm{C}^{*}$- (respectively, $\mathrm{W}^{*}$-) dynamical system if $\alpha$ is a $\mathrm{C}^{*}$-($\mathrm{W}^{*}$-)action of $G$ on $A$.
As in [7], we say that a $\mathrm{W}^{*}$-action $\gamma$ of $G$ on $M$ is amenable (or $\mathrm{W}^{*}$-amenable) if there exists an equivariant conditional expectation $P\colon L^{\infty}(G)\bar{\otimes}M\to M$, where we identify each $m\in M$ with $1\otimes m\in L^{\infty}(G)\bar{\otimes}M$ and the action of $G$ on $L^{\infty}(G)\bar{\otimes}M$ is the diagonal action induced by conjugation by the left regular representation $\lambda\colon G\to\mathbb{B}(L^{2}(G))$ and $\gamma$. It was first noticed in [7, 8] that the amenability of $\gamma$ is related to that of the restriction of $\gamma$ to the center of $M$, $\gamma|_{Z(M)}$. Anantharaman-Delaroche proved they are equivalent under separability conditions and the general case is a consequence of [11, Theorem 3.17].
The continuous part of $\gamma$, denoted $\gamma^{c}$, is the restriction of $\gamma$ to the $\mathrm{C}^{*}$-subalgebra $M^{c}\subseteq M$ formed by those $m\in M$ such that $t\mapsto\gamma_{t}(m)$ is norm continuous. Recall from [8] that $M^{c}$ is $\operatorname{\rm w}^{*}$}-dense in $M$ and notice that $\gamma^{c}$ is a $\mathrm{C}^{*}$-action. The separability assumptions made in the introduction of [8] are not necessary in the proof of [8, Lemme 2.1], so $\gamma$ is $\mathrm{W}^{*}$-amenable if and only if there exists an equivariant conditional expectation $P\colon(L^{\infty}(G)\bar{\otimes}Z(M))^{c}\to Z(M)^{c}$. We used Anantharan-Delaroche’s ideas to prove the extension lemma of Appendix B.
In Appendix A, we address the problem of constructing the (von Neumann) tensor product of possibly non-normal, equivariant, contractive completely positive (ccp) maps between $\mathrm{W}^{*}$-algebras. Our approach combines techniques from [8, Lemme 2.1], [20, Theorem 3], and [23, Theorem 4]. While these methods are likely familiar to experts, we were unable to find a precise reference in the literature.
The notion of amenability of a $\mathrm{C}^{*}$-action $\alpha$ of $G$ on $A$ requires the use of the enveloping $G$-$\mathrm{W}^{*}$-algebra of $\alpha,$ $A^{\prime\prime}_{\alpha}$, which we introduce here using the presentation of [11]. Let $A\rtimes_{\alpha}G$ be the (full) crossed-product of $\alpha$ and $A\rtimes_{\alpha}G\subseteq\mathbb{B}(X_{\alpha})$ the universal representation. This representation is the integrated form of the “universal” unitary implementation $(A\subseteq\mathbb{B}(X_{\alpha}),U^{\alpha})$. By definition, $A^{\prime\prime}_{\alpha}$ is the bicommutant of $A$ in $\mathbb{B}(X_{\alpha})$ and the $\mathrm{W}^{*}$-action $\alpha^{\prime\prime}$ of $G$ on $A^{\prime\prime}_{\alpha}$ is given by $(t,m)\mapsto U^{\alpha}_{t}mU^{\alpha}_{t^{-1}}\equiv\operatorname{Ad}(U^{\alpha})_{t}(m).$ We say that $\alpha$ is $\mathrm{C}^{*}$-amenable if $\alpha^{\prime\prime}$ is $\mathrm{W}^{*}$-amenable.
We recall here some facts we will need later on. Given a covariant representation $(X,\pi,U)$ of $\alpha$, when no confusion can arise, we write $A\rtimes G$ instead of $A\rtimes_{\alpha}G$. In general, $\pi\rtimes U\colon A\rtimes_{\alpha}G\to\mathbb{B}(X)$ will denote the integrated form of $(\pi,U)$. For all $t\in G$ and $f\in C_{c}(G,A)$ we have $U_{t}\pi\rtimes U(f)=\pi\rtimes U(\tilde{\alpha}_{t}(f))$ with $\tilde{\alpha}_{t}(f)(s)=\alpha_{t}(f({t^{-1}}s)).$ Similarly, for $a\in A$ we define $af$ as the pointwise product of $a$ and $f$ and we have $\pi(a)\pi\rtimes U(f)=\pi\rtimes U(af)$.
The reduced covariant representation $(L^{2}(G)\otimes X,1\pi,\lambda U)$ associated to $(X,\pi,U)$ is given by $1\pi(a)=1\otimes\pi(a)$ and $\lambda U_{t}=\lambda_{t}\otimes U_{t}$; where $\lambda\colon G\to\mathbb{B}(L^{2}(G))$ is the left regular representation. Recall that $1\pi\rtimes\lambda U$ factors through a representation $1\pi\rtimes_{\mathop{\rm r}}\lambda U$ of the reduced crossed product $A\rtimes_{{\mathop{\rm r}}\alpha}G$; that $1\pi\rtimes_{\mathop{\rm r}}\lambda U$ is faithful if and only if $\pi$ is faithful and, finally, that $\alpha$ has the wcp if and only if some $1\pi\rtimes\lambda U$ is faithful; in which case every such representation with $\pi$ faithful will be faithful as well.
Remark 2.1.
[…]
The key to the main result of this section is
Lemma 2.2.
Let $(X,\pi,U)$ be a nondegenerate covariant representation of a $\mathrm{C}^{*}$-action $\alpha$ of a group $G$ on a $\mathrm{C}^{*}$-algebra $A$ and set $M:=\pi(A)^{\prime}$ (the commutant). Let $\beta$ be the continuous part of $\operatorname{Ad}(U)|_{M}$, viewed as $\mathrm{C}^{*}$-action of $G$ on $M^{c}$. If the diagonal action $\alpha\otimes_{\max}^{d}\beta$ of $G$ on $A\otimes_{\max}M^{c}$ has the wcp, then $\operatorname{Ad}(U)|_{M}$ and $\operatorname{Ad}(U)|_{\pi(A)^{\prime\prime}}$ are $\mathrm{W}^{*}$-amenable.
Proof.
Notice that $Z(\pi(A)^{\prime\prime})=Z(\pi(A)^{\prime})=Z(M)$. So, as discussed previously, it suffices to show that there exists an equivariant conditional expectation $P\colon(L^{\infty}(G)\bar{\otimes}Z(M))^{c}\to Z(M)^{c}$.
Let $(Y,\rho,V)$ be a faithful and nondegenerate covariant representation of $\alpha\otimes_{\max}^{d}\beta$ and split it into covariant representations $(Y,\rho_{A},V)$ and $(Y,\rho_{c},V)$ of $\alpha$ and $\beta$, that is, $\rho_{A}(a)\rho_{c}(m)=\rho(a\otimes m)$ for $a\in A$ and $m\in M$. Since $\alpha\otimes_{\max}^{d}\beta$ has the wcp, $1\rho\rtimes\lambda V$ is faithful and
$C:=1\rho\rtimes\lambda V((A\otimes_{\max}M^{c})\rtimes G)\subseteq\mathbb{B}(L^{2}(G)\otimes Y)$
is isomorphic to the full crossed product. Let $\iota^{c}\colon M^{c}\to\mathbb{B}(X)$ be the inclusion and $\psi_{0}\colon C\to\mathbb{B}(X)$ the unique *-homomorphism such that $\psi_{0}\circ(1\rho\rtimes\lambda V)=(\pi\times\iota^{c})\rtimes U$; where $\pi\times\iota^{c}(a\otimes m)=\pi(a)m$.
Fix a ccp extension $\psi\colon\mathbb{B}(L^{2}(G)\otimes Y)\to\mathbb{B}(X)$ of $\psi_{0}$ and name $D$ the multiplicative domain of $\psi$ [10, Proposition 1.5.7 & Definition 1.5.8]. Recall that
$D$ $=\{T\in\mathbb{B}(L^{2}(G)\otimes Y)\colon\psi(T^{*}T)=\psi(T)^{*}\psi(T),\ \psi(TT^{*})=\psi(T)\psi(T)^{*}\}$
$=\{T\in\mathbb{B}(L^{2}(G)\otimes Y)\colon\forall\ S\in\mathbb{B}(L^{2}(G)\otimes Y),\ \psi(ST)=\psi(S)\psi(T),\ \psi(TS)=\psi(T)\psi(S)\}.$
From the first identity it follows that $C\subseteq D$. Now fix $t\in G$. To prove that $\psi(\lambda V_{t})=U_{t}$ we take $f\in C_{c}(G,A\otimes_{\max}M^{c})$ and $x\in X$. We have
$\psi(\lambda V_{t})(\pi\times\iota^{c})\rtimes U(f)x$ $=\psi(\lambda V_{t})\psi(1\rho\rtimes\lambda V(f))x=\psi\left((\lambda V_{t})1\rho\rtimes\lambda V(f)\right)x=\psi\left(1\rho\rtimes\lambda V(\tilde{\alpha}_{t}(f))\right)x$
$=(\pi\times\iota^{c})\rtimes U(\tilde{\alpha}_{t}(f))x=U_{t}(\pi\times\iota^{c})\rtimes U(f)x.$
Since $(\pi\times\iota^{c})\rtimes U$ is nondegenerate, the equalities above imply that $\psi(\lambda V_{t})=U_{t}$. In particular, $\psi$ is unital and this yields $\lambda V_{t}\in D$ (because $\lambda V_{t}$ and $U_{t}$ are unitaries). Notice that this implies that $\psi$ is equivariant with respect to $\operatorname{Ad}(\lambda V)$ and $\operatorname{Ad}(U)$.
We claim that $\psi\circ(1\rho)=\pi\times\iota^{c}$. Indeed, this is a straightforward consequence of the fact that for all $a\in A\otimes_{\max}M^{c}$, $f\in C_{c}(G,A\otimes_{\max}M^{c})$ and $x\in X$,
$\psi(1\rho(a))(\pi\times\iota^{c})\rtimes U(f)x$ $=\psi(1\rho(a))\psi(1\rho\rtimes\lambda V(f))x=\psi\left((1\rho(a))1\rho\rtimes\lambda V(f)\right)x$
$=\psi\left(1\rho\rtimes\lambda V(af)\right)x=(\pi\times\iota^{c})\rtimes U(af)x=(\pi\times\iota^{c}(a))(\pi\times\iota^{c})\rtimes U(f)x.$
The restriction of $\psi$ to the image of $1\rho$ is a representation, so $1\rho(A\otimes_{\max}M^{c})\subseteq D$. Moreover, for all $a\in A$, $\psi(1\otimes\rho_{A}(a))=\psi(1\rho(a\otimes 1))=\pi(a)$. Thus, for all $m\in M^{c}$, $\psi(1\otimes\rho_{c}(m))\pi(a)=\psi(1\otimes\rho(a\otimes m))=\pi(a)\iota^{c}(m)=\iota^{m}(c)\pi(a)$. Replacing $a$ with an approximate unit of $A$ and taking limit in the weak operator topology (wot) it follows that $\psi(1\otimes\rho_{c}(m))=\iota^{c}(m)$. This implies that the image of $1\otimes\rho_{c}$ is contained in $D$.
Let $\overline{\rho_{c}}\colon M\to\mathbb{B}(Y)$ be the extension of $\rho_{c}$ given by Lemma B.1. Then $\rho_{A}(A)\subseteq\rho_{c}(M^{c})^{\prime}=\overline{\rho_{c}}(M)^{\prime}$ and $\overline{\rho_{c}}(M)\subseteq\overline{\rho_{c}}(M)^{\prime\prime}\subseteq\rho_{A}(A)^{\prime}$. Since $\rho_{c}$ is a *-homomorphism and $\overline{\rho_{c}}$ an extension of $\rho_{c}$, $M^{c}$ is contained in the multiplicative domain of $\rho$ and it follows that for all $T\in Z(M)$ and $m\in M^{c}$, $\overline{\rho_{c}}(T)\rho_{c}(m)=\overline{\rho_{c}}(Tm)=\overline{\rho_{c}}(mT)=\rho_{c}(m)\overline{\rho_{c}}(T)$. Hence, $\overline{\rho_{c}}(Z(M))\subseteq\rho_{A}(A)^{\prime}\cap\rho_{c}(M)^{\prime}.$
Viewing $\overline{\rho_{c}}|_{Z(M)}$ as a ucp equivariant function from $Z(M)$ to $N:=\rho_{A}(A)^{\prime}\cap\rho_{c}(M)^{\prime}$, we can use Theorem A.7 to get a ucp and $\operatorname{Ad}(\lambda)\bar{\otimes}(\gamma|_{Z(M)})-\operatorname{Ad}(\lambda)\bar{\otimes}(\operatorname{Ad}(V)|_{N})$ equivariant map
$\mu\colon L^{\infty}(G)\bar{\otimes}Z(M)\to L^{\infty}(G)\bar{\otimes}N\subseteq\mathbb{B}(L^{2}(G)\otimes Y)$
such that, for all $f\in L^{\infty}(G)$ and $m\in Z(M)$, $\mu(f\otimes m)=f\otimes\overline{\rho_{c}}(m)$.
We claim that $\psi(\mu(L^{\infty}(G)\bar{\otimes}Z(M)))\subseteq Z(M)$. To prove this we use the fact that
$Z(M)=M\cap M^{\prime}\equiv\pi(A)^{\prime}\cap M^{\prime}=(\pi\times\iota^{c}(A\otimes_{\max}M^{c}))^{\prime}.$
Take $T\in L^{\infty}(G)\bar{\otimes}Z(M)$ and $a\in A\otimes_{\max}M^{c}$. Since $1\rho(a)\in D$; $\psi(1\rho(a))=\pi\times\iota^{c}(a)$ and $L^{\infty}(G)\bar{\otimes}N\subseteq[1\rho(A\otimes_{\max}M^{c})]^{\prime}$ we get $\psi(\mu(T))\pi\times\iota^{c}(a)=\psi(\mu(T)1\rho(a))=\psi(1\rho(a)\mu(T))=\pi\times\iota^{c}(a)\psi(\mu(T)),$ so $\psi(\mu(T))\in Z(M).$
Since both $\mu$ and $\psi$ are equivariant and norm continuous, $\psi(\mu((L^{\infty}(G)\bar{\otimes}Z(M))^{c}))\subseteq Z(M)^{c}$ and we get a ccp equivariant map
$P$ $\colon(L^{\infty}(G)\bar{\otimes}Z(M))^{c}\to Z(M)^{c}$ $T$ $\mapsto\psi(\mu(T))$
For $m\in Z(M)^{c}\subseteq L^{\infty}(G)\bar{\otimes}Z(M)$ we have $m\in M^{c}$ and $P(1\otimes m)=\psi(1\otimes\rho_{c}(m))=\iota^{c}(m)=m.$ This shows that $P$ is an equivariant conditional expectation. ∎
[…]
Theorem 2.3.
Let $\alpha$ be a $\mathrm{C}^{*}$-action of a group $G$ on a $\mathrm{C}^{*}$-algebra $A$. Using the universal unitary implementation $(A\subseteq\mathbb{B}(X_{\alpha}),U^{\alpha})$, let $A^{\prime}_{\alpha}$ be the commutant of $A$ in $\mathbb{B}(X_{\alpha})$, and let $\beta$ be the continuous part of $\operatorname{Ad}(U^{\alpha})|_{A^{\prime}}$. In this setting, the following claims are equivalent.
  1. (1)
    $\alpha$ is $\mathrm{C}^{*}$-amenable.
  2. (2)
    For every $\mathrm{C}^{*}$-action $\gamma$ of $G$, the diagonal $\mathrm{C}^{*}$-action $\alpha\otimes_{\max}^{d}\gamma$ of $G$ has the wcp.
  3. (3)
    The $\mathrm{C}^{*}$-action $\alpha\otimes_{\max}^{d}\beta$ has the wcp.
Proof.
Since (2) trivially implies (3), it suffices to show that (3) implies (1), but this is a straightforward consequence of Lemma 2.2. ∎
Remark 2.4.
[…]
2.2. SIN groups and amenability
A group $G$ has small invariant neighborhoods (SIN) if there exists a basis $\mathcal{U}$ of compact neighborhoods of $e\in G$, each of which is invariant under conjugation: for all $(t,U)\in G\times\mathcal{U}$, $tUt^{-1}=U$. Every discrete group is SIN, and every SIN group is inner amenable and unimodular (compute the measure of $tUt^{-1}=U$). The purpose of this section is to provide an alternative proof, for SIN groups, of the following result of McKee and Pourshahami.
Theorem 2.5 (Corollary 6.6 of [19]).
If $G$ is an inner amenable group and $\alpha$ is a $\mathrm{C}^{*}$-action of $G$ on $A$ such that $A\rtimes_{{\mathop{\rm r}},\alpha}G$ is nuclear, then $\alpha$ is amenable.
In Section 3.1 we combine results of several authors to obtain more general statements (valid even for Fell bundles) than those presented in this section. Here, however, we provide self-contained proofs.
We use the letter $\lambda$ to denote the left regular representation of $G$, while $\rho\colon G\to\mathbb{B}(L^{2}(G))$ denotes the right regular representation. We also define $\omega\colon G\to\mathbb{B}(L^{2}(G))$ by setting $\omega_{t}:=\lambda_{t}\rho_{t}$.
Let $\alpha$ and $\beta$ be $\mathrm{C}^{*}$-actions of $G$ on the $\mathrm{C}^{*}$-algebras $A$ and $B$, respectively. We write $\otimes$ for the minimal tensor product of $\mathrm{C}^{*}$-algebras, and $\alpha\otimes^{d}\beta$ for the diagonal action of $G$ on $A\otimes B$.
The natural maps
$\iota^{{\mathop{\rm r}}\alpha}$ $\colon A\to\mathcal{M}(A\rtimes_{{\mathop{\rm r}}\alpha}G),$ $\quad u^{{\mathop{\rm r}}\alpha}$ $\colon G\to\mathcal{M}(A\rtimes_{{\mathop{\rm r}}\alpha}G),$
$\iota^{\alpha}$ $\colon A\to\mathcal{M}(A\rtimes_{\alpha}G),$ $\quad u^{\alpha}$ $\colon G\to\mathcal{M}(A\rtimes_{\alpha}G),$
form the covariant pairs $(\iota^{{\mathop{\rm r}}\alpha},u^{{\mathop{\rm r}}\alpha})$ and $(\iota^{\alpha},u^{\alpha})$. For all $a\in A$, $s\in G$, and $f\in C_{c}(G,A)$, we have
$\iota^{{\mathop{\rm r}}\alpha}(a)f=\iota^{\alpha}(a)f=af,\qquad u^{{\mathop{\rm r}}\alpha}_{s}f=u^{\alpha}_{s}f=\tilde{\alpha}_{s}(f),$
where $af$ denotes the pointwise product of $a$ and $f$, and
$\tilde{\alpha}_{s}(f)(t)=\alpha_{t}\big(f(t^{-1}s)\big).$
Proposition 2.6.
The covariant pair $(\iota^{{\mathop{\rm r}}\alpha}\otimes\iota^{{\mathop{\rm r}}\beta},u^{{\mathop{\rm r}}\alpha}\otimes^{d}u^{{\mathop{\rm r}}\beta})$ of $\alpha\otimes^{d}\beta$, defined by
$\iota^{{\mathop{\rm r}}\alpha}\otimes\iota^{{\mathop{\rm r}}\beta}\colon A\otimes B$ $\longrightarrow\mathcal{M}\big((A\rtimes_{{\mathop{\rm r}}\alpha}G)\otimes(B\rtimes_{{\mathop{\rm r}}\beta}G)\big),$ $(a\otimes b)$ $\longmapsto\iota^{{\mathop{\rm r}}\alpha}(a)\otimes\iota^{{\mathop{\rm r}}\beta}(b),$
$u^{{\mathop{\rm r}}\alpha}\otimes^{d}u^{{\mathop{\rm r}}\beta}\colon G$ $\longrightarrow\mathcal{M}\big((A\rtimes_{{\mathop{\rm r}}\alpha}G)\otimes(B\rtimes_{{\mathop{\rm r}}\beta}G)\big),$ $t$ $\longmapsto u^{{\mathop{\rm r}}\alpha}_{t}\otimes u^{{\mathop{\rm r}}\beta}_{t},$
integrates to a representation that factors through a faithful representation of $(A\otimes B)\rtimes_{{\mathop{\rm r}}\,(\alpha\otimes^{d}\beta)}G$.
Proof.
Take unitary implementations $(X,\pi_{A},U)$ and $(Y,\pi_{B},V)$ of $\alpha$ and $\beta$, respectively. Then $\iota^{{\mathop{\rm r}}\alpha}$ can be viewed as
$A\to\mathbb{B}(L^{2}(G)\otimes X\otimes L^{2}(G)\otimes Y),\qquad a\mapsto 1\otimes\pi_{A}(a)\otimes 1\otimes 1,$
with $u^{{\mathop{\rm r}}\alpha}_{r}=\lambda_{r}\otimes U_{r}\otimes 1\otimes 1$. Similarly,
$\iota^{{\mathop{\rm r}}\beta}(b)=1\otimes 1\otimes 1\otimes\pi_{B}(b),\qquad u^{{\mathop{\rm r}}\beta}_{r}=1\otimes 1\otimes\lambda_{r}\otimes V_{r}.$
With this notation, we obtain
$\iota^{{\mathop{\rm r}}\alpha}\otimes\iota^{{\mathop{\rm r}}\beta}(a\otimes b)=1\otimes\pi_{A}(a)\otimes 1\otimes\pi_{B}(b),\qquad(u^{{\mathop{\rm r}}\alpha}\otimes^{d}u^{{\mathop{\rm r}}\beta})_{t}=\lambda_{t}\otimes U_{t}\otimes\lambda_{t}\otimes V_{t}.$
In the notation of Section 2.1, this means
$(\iota^{{\mathop{\rm r}}\alpha}\otimes\iota^{{\mathop{\rm r}}\beta},\,u^{{\mathop{\rm r}}\alpha}\otimes^{d}u^{{\mathop{\rm r}}\beta})=(1\otimes(\pi_{A}\otimes 1\otimes\pi_{B}),\,\lambda\otimes(U\otimes\lambda\otimes V)).$
The conclusion follows immediately from the fact that $\pi_{A}\otimes 1\otimes\pi_{B}$ is a faithful representation of $A\otimes B$. ∎
Proposition 2.7.
If $G$ is SIN, then the covariant pair $(\iota^{{\mathop{\rm r}}\alpha}\otimes_{\max}\iota^{{\mathop{\rm r}}\beta},u^{{\mathop{\rm r}}\alpha}\otimes_{\max}^{d}u^{{\mathop{\rm r}}\beta})$ of $\alpha\otimes_{\max}^{d}\beta$, given by
$\iota^{{\mathop{\rm r}}\alpha}\otimes_{\max}\iota^{{\mathop{\rm r}}\beta}\colon A\otimes_{\max}B$ $\longrightarrow\mathcal{M}\big((A\rtimes_{{\mathop{\rm r}}\alpha}G)\otimes_{\max}(B\rtimes_{{\mathop{\rm r}}\beta}G)\big),$ $(a\otimes b)$ $\longmapsto\iota^{{\mathop{\rm r}}\alpha}(a)\otimes\iota^{{\mathop{\rm r}}\beta}(b),$
$u^{{\mathop{\rm r}}\alpha}\otimes_{\max}^{d}u^{{\mathop{\rm r}}\beta}\colon G$ $\longrightarrow\mathcal{M}\big((A\rtimes_{{\mathop{\rm r}}\alpha}G)\otimes_{\max}(B\rtimes_{{\mathop{\rm r}}\beta}G)\big),$ $t$ $\longmapsto u^{{\mathop{\rm r}}\alpha}_{t}\otimes u^{{\mathop{\rm r}}\beta}_{t},$
integrates to a faithful representation of $(A\otimes_{\max}B)\rtimes_{\alpha\otimes_{\max}^{d}\beta}G$.
Proof.
Let $(X,\pi,U)$ be a covariant representation of $\alpha\otimes_{\max}^{d}\beta$ such that the integrated form $\pi\rtimes U$ is faithful and nondegenerate. Decompose $\pi$ into representations $\pi_{A}\colon A\to\mathbb{B}(X)$ and $\pi_{B}\colon B\to\mathbb{B}(X)$, so that $\pi(a\otimes b)=\pi_{A}(a)\pi_{B}(b)$. Then $(X,\pi_{A},U)$ and $(X,\pi_{B},U)$ are covariant representations.
Using the notation of Section 2.1, we construct the covariant representations
$(L^{2}(G)\otimes X,\,1\pi_{A},\,\lambda U)\quad\text{and}\quad(L^{2}(G)\otimes X,\,1\pi_{B},\,\lambda U),$
which induce faithful representations
$1\pi_{A}\rtimes_{\mathop{\rm r}}\lambda U$ $\colon A\rtimes_{{\mathop{\rm r}}\alpha}G\to\mathbb{B}(L^{2}(G)\otimes X),$ $1\pi_{B}\rtimes_{\mathop{\rm r}}\lambda U$ $\colon B\rtimes_{{\mathop{\rm r}}\beta}G\to\mathbb{B}(L^{2}(G)\otimes X).$
We conjugate the second by the unitary $W\in\mathbb{B}(L^{2}(G)\otimes X)$ defined by
$(W\xi)(r)=\Delta(r)^{-1/2}U_{r}\xi(r^{-1}).$
Note that $W=W^{*}=W^{-1}$. Define $\overline{1\pi_{B}}(b):=W(1\pi_{B})(b)W$ and $\overline{\lambda U}_{r}:=W(\lambda U)_{r}W$. Then
$\overline{1\pi_{B}}(b)\xi(r)=\pi_{B}(\beta_{r}(b))\xi(r),\qquad\overline{\lambda U}_{r}=\rho_{r}\otimes 1.$
Straightforward computations show that for all $r,s\in G$, $a\in A$, and $b\in B$,
$(\lambda U_{r})\,\overline{\lambda U}_{s}$ $=\overline{\lambda U}_{s}\,(\lambda U_{r}),$ $(\lambda U_{r})\,\overline{1\pi_{B}}(b)$ $=\overline{1\pi_{B}}(b)\,(\lambda U_{r}),$
$(1\pi_{A})(a)\,\overline{1\pi_{B}}(b)$ $=\overline{1\pi_{B}}(b)\,(1\pi_{A})(a),$ $(1\pi_{A})(a)\,\overline{\lambda U}_{s}$ $=\overline{\lambda U}_{s}\,(1\pi_{A})(a).$
Hence, the images of $1\pi_{A}\rtimes_{\mathop{\rm r}}\lambda U$ and $\overline{1\pi_{B}}\rtimes_{\mathop{\rm r}}\overline{\lambda U}:=\operatorname{Ad}(W)\circ(1\pi_{B}\rtimes_{\mathop{\rm r}}\lambda U)$ commute, and we obtain a representation
$\kappa:=(1\pi_{A}\rtimes_{\mathop{\rm r}}\lambda U)\otimes_{\max}(\overline{1\pi_{B}}\rtimes_{\mathop{\rm r}}\overline{\lambda U})\colon(A\rtimes_{{\mathop{\rm r}}\alpha}G)\otimes_{\max}(B\rtimes_{{\mathop{\rm r}}\beta}G)\to\mathbb{B}(L^{2}(G)\otimes X).$
Extending $\kappa$ to the multiplier algebra and composing with $\iota^{{\mathop{\rm r}}\alpha}\otimes_{\max}\iota^{{\mathop{\rm r}}\beta}$ and $u^{{\mathop{\rm r}}\alpha}\otimes_{\max}^{d}u^{{\mathop{\rm r}}\beta}$, we obtain the covariant representation $(L^{2}(G)\otimes X,\Pi,V)$ with $V_{r}=\omega_{r}\otimes U_{r}$.
Let $\mathcal{U}$ be a basis of compact invariant neighborhoods of $e$. For each $u\in\mathcal{U}$ let $\mathbb{1}_{u}\in C_{c}(G)$ be the characteristic function of $u$, $m_{u}$ its measure, and set $\xi_{u}:=(m_{u})^{-1/2}\mathbb{1}_{u}$, so that $\|\xi_{u}\|_{2}=1$.
For all $a\in A$, $b\in B$, $x,y\in X$, and $r\in G$ we have
$\langle\Pi(a\otimes b)V_{r}(\xi_{u}\otimes x),\xi_{u}\otimes y\rangle$ $=\int_{G}\langle\pi_{A}(a)\pi_{B}(\beta_{s}(b))U_{r}x,y\rangle\,\overline{\omega_{r}\xi_{u}}(s)\,\xi_{u}(s)\,ds$
$=(m_{u})^{-1}\int_{u}\langle\pi_{A}(a)\pi_{B}(\beta_{s}(b))U_{r}x,y\rangle\,ds$
$=(m_{u})^{-1}\int_{u}\langle\pi\circ(\operatorname{id}_{A}\otimes\beta_{s})(a\otimes b)\,U_{r}x,y\rangle\,ds.$
Now let $f\in C_{c}(G,A\otimes_{\max}B)$. Since the image of $f$ is compact,
$\lim_{u}\sup\{\|(\operatorname{id}_{A}\otimes\beta_{s})(f(t))-f(t)\|\colon s\in u,\ t\in G\}=0.$
Writing $m_{f}$ for the measure of ${\mathop{\rm supp}}(f)$, we obtain
$\lim_{u}\Big|\langle\Pi\rtimes V(f)(\xi_{u}\otimes x),\xi_{u}\otimes y\rangle-\langle\pi\rtimes U(f)x,y\rangle\Big|\\ \leq\lim_{u}\int_{G}(m_{u})^{-1}\int_{u}\big|\langle\pi\circ(\operatorname{id}_{A}\otimes\beta_{s})(f(r))U_{r}x,y\rangle-\langle\pi(f(r))U_{r}x,y\rangle\big|\,ds\,dr\\ \leq\lim_{u}m_{f}\|x\|\|y\|\sup\{\|(\operatorname{id}_{A}\otimes\beta_{s})(f(t))-f(t)\|\colon s\in u,\ t\in G\}=0.$
Since $\|\xi_{u}\otimes x\|=\|x\|$, it follows immediately that $\|\pi\rtimes U(f)\|\leq\|\Pi\rtimes V(f)\|$. By continuity, the same inequality holds for all $f\in(A\otimes_{\max}B)\rtimes_{\alpha\otimes_{\max}^{d}\beta}G$, and hence
$\Pi\rtimes V=\kappa\circ\big((\iota^{{\mathop{\rm r}}\alpha}\otimes_{\max}\iota^{{\mathop{\rm r}}\beta})\rtimes(u^{{\mathop{\rm r}}\alpha}\otimes_{\max}^{d}u^{{\mathop{\rm r}}\beta})\big)$
is a faithful representation. ∎
Proposition 2.8.
If $A$ is nuclear and $\alpha$ is $\mathrm{C}^{*}$-amenable, then $A\rtimes_{\mathop{\rm r}}G$ is nuclear. The converse holds if $G$ is SIN.
Proof.
The forward implication is already known [11, Theorem 7.2], but we provide a proof here for the convenience of the reader. Take any $\mathrm{C}^{*}$-algebra $B$ and equip it with the trivial action $\beta$ of $G$. Then $\alpha\otimes_{\max}^{d}\beta=\alpha\otimes^{d}\beta$ (in particular, $\alpha$) has the wcp, and
$(A\rtimes_{{\mathop{\rm r}}\alpha}G)\otimes_{\max}B$ $=(A\rtimes_{\alpha}G)\otimes_{\max}B=(A\otimes_{\max}B)\rtimes_{\alpha\otimes_{\max}^{d}\beta}G=(A\otimes B)\rtimes_{\alpha\otimes^{d}\beta}G$
$=(A\otimes B)\rtimes_{{\mathop{\rm r}}\alpha\otimes^{d}\beta}G=(A\rtimes_{{\mathop{\rm r}}\alpha}G)\otimes B,$
showing that $A\rtimes_{{\mathop{\rm r}}\alpha}G$ is nuclear.
For the converse, let $B$ be any $\mathrm{C}^{*}$-algebra equipped with the trivial action $\beta$ of $G$. If $\kappa\colon A\otimes_{\max}B\to A\otimes B$ denotes the natural quotient map, then
$\iota^{{\mathop{\rm r}}\alpha}\otimes_{\max}\iota^{{\mathop{\rm r}}\beta}=(\iota^{{\mathop{\rm r}}\alpha}\otimes\iota^{{\mathop{\rm r}}\beta})\circ\kappa$
is faithful. Hence $\kappa$ is faithful, and it follows that $A$ is nuclear.
Now let $\beta$ be an arbitrary $\mathrm{C}^{*}$-action of $G$ on a $\mathrm{C}^{*}$-algebra $B$. By Propositions 2.6 and 2.7 we have
$(\iota^{{\mathop{\rm r}}\alpha}\otimes\iota^{{\mathop{\rm r}}\beta},u^{{\mathop{\rm r}}\alpha}\otimes^{d}u^{{\mathop{\rm r}}\beta})=(\iota^{{\mathop{\rm r}}\alpha}\otimes_{\max}\iota^{{\mathop{\rm r}}\beta},u^{{\mathop{\rm r}}\alpha}\otimes_{\max}^{d}u^{{\mathop{\rm r}}\beta}).$
Therefore $\alpha\otimes_{\max}^{d}\beta$ has the wcp, and Theorem 2.3 implies that $\alpha$ is amenable. ∎
We suspect that the conclusion of Proposition 2.7 also holds for inner amenable groups. If this is indeed the case, then the proof of Proposition 2.8 would immediately yield Theorem 2.5.
3. Amenability of Fell bundles and $\mathrm{C}^{*}$-partial actions
When we say that $\mathcal{B}=\{B_{t}\}_{t\in G}$ is a Fell bundle we mean that the disjoint union $\bigsqcup_{t\in G}B_{t}$ forms a $\mathrm{C}^{*}$-algebraic bundle over $G$ in the sense of [16, Ch. VIII, 16.2], with fibers $B_{t}$. Following [3], we denote by $C_{c}(\mathcal{B})$, $L^{1}(\mathcal{B})$, $C^{*}(\mathcal{B})$, $C^{*}_{\mathop{\rm r}}(\mathcal{B})$, and $\mathbb{k}(\mathcal{B})$ the *-algebra of continuous cross-sections with compact support, the $L^{1}$ cross-sectional algebra, the full cross-sectional $\mathrm{C}^{*}$-algebra, the reduced cross-sectional $\mathrm{C}^{*}$-algebra, and the $\mathrm{C}^{*}$-algebra of kernels of $\mathcal{B}$, respectively.
Recall that the algebra of kernels with compact support, $\mathbb{k}_{c}(\mathcal{B})$, consists of compactly supported continuous functions $k\colon G\times G\to\mathcal{B}$ with $k(s,t)\in B_{st^{-1}}$. It is a dense *-subalgebra of $\mathbb{k}(\mathcal{B})$, fixed under the canonical action $\beta$ of $G$ on $\mathbb{k}(\mathcal{B})$, which is given
$\beta_{r}(k)(s,t)\;=\;k(sr,tr)\,\Delta(r),\qquad(r,s,t\in G,\;k\in\mathbb{k}_{c}(\mathcal{B})).$
As usual, the semidirect product bundle associated to a $\mathrm{C}^{*}$-action $\alpha$ of $G$ on $A$ is $A\times G$, with operations
$(a,r)(b,s)=(a\,\alpha_{r}(b),rs),\qquad(a,r)^{*}=(\alpha_{r^{-1}}(a)^{*},r^{-1}).$
It is customary to write $a\delta_{r}$ instead of $(a,r)$, so that the fiber $A\times\{r\}$ is denoted $A\delta_{r}$.
In general, semidirect product bundles can be constructed from twisted partial actions on $\mathrm{C}^{*}$-algebras [13]. However, we shall only use $\mathrm{C}^{*}$-partial actions here, as in [3].
The reduced cross-sectional $\mathrm{C}^{*}$-algebra of a Fell bundle was constructed by Exel and Ng in [15]. For a Fell bundle $\mathcal{B}$, we denote by
$\lambda^{\mathcal{B}}\colon C^{*}(\mathcal{B})\longrightarrow C^{*}_{\mathop{\rm r}}(\mathcal{B})$
the canonical surjective -homomorphism given by the reduced representation [15, Definition 2.7]. We say that $\mathcal{B}$ has the weak containment property (wcp) if $\lambda^{\mathcal{B}}$ is injective, that is, if the full and reduced cross-sectional $\mathrm{C}^{*}$-algebras of $\mathcal{B}$ coincide. Exel and Ng referred to this property as amenability of the Fell bundle, but in order to avoid confusion with other notions of amenability used in this article, we adopt the terminology wcp instead (see Definition 3.10 and Remark 3.11).
Whenever we say that a Fell bundle has an approximation property in the sense of Exel and Ng, we mean one of those introduced in [15, Definition 3.6]; by [15, Theorem 3.9], each of them implies the wcp. The $L^{1}$-approximation property, introduced by Abadie in [1, Definition 4.9(3)], also implies the wcp [1, Theorem 4.11], and is formally weaker than the Exel–Ng approximation properties [1, Proposition 4.10]. In what follows, we shall prove that all these approximation properties are in fact equivalent.
Abadie constructed a maximal tensor product of Fell bundles, denoted $\bigotimes_{\max}$, using the maximal tensor product of $\mathrm{C}^{*}$-algebras [1, Proposition 3.4 and Definition 3.8]. Given Fell bundles $\mathcal{A}=\{A_{t}\}_{t\in G}$ and $\mathcal{B}=\{B_{t}\}_{t\in H}$, their tensor product
$\mathcal{A}\bigotimes_{\max}\mathcal{B}=\{A_{s}\otimes_{\max}B_{t}\}_{(s,t)\in G\times H}$
is a Fell bundle over $G\times H$, where each fiber is the maximal tensor product $A_{s}\otimes_{\max}B_{t}$ (of right Hilbert modules over $A_{e}$ and $B_{e}$). The operations are given by $(a\otimes b)(a^{\prime}\otimes b^{\prime})=aa^{\prime}\otimes bb^{\prime}$ and $(a\otimes b)^{*}=a^{*}\otimes b^{*}$, and the topology is the unique Banach bundle topology for which
$f\oslash g\colon G\times H\to\mathcal{A}\bigotimes_{\max}\mathcal{B},\qquad(s,t)\mapsto f(s)\otimes g(t),$
is continuous for all $f\in C_{c}(\mathcal{A})$ and $g\in C_{c}(\mathcal{B})$. When $G=H$, we view $G$ as the diagonal of $G\times G$ and define the diagonal tensor product as the reduction of $\mathcal{A}\bigotimes_{\max}\mathcal{B}$ to $G$, that is
$\mathcal{A}\otimes_{\max}^{d}\mathcal{B}:=\{A_{t}\otimes_{\max}B_{t}\}_{t\in G},$
so that for $f\in C_{c}(\mathcal{A})$ and $g\in C_{c}(\mathcal{B})$ one has
$(f\oslash^{d}g)(t):=f(t)\otimes g(t)\equiv(f\oslash g)(t,t).$
Before stating our main theorem for Fell bundles, we recall the relevant approximation properties.
We use [16, VIII 3.8] to regard the multiplier algebra $\mathcal{M}(B_{e})$ of the fiber over the unit $e\in G$ as the $\mathrm{C}^{*}$-algebra of multipliers of $\mathcal{B}$ of order $e$. As shown in [1], the natural actions $\mathcal{M}(B_{e})\times\mathcal{B}\to\mathcal{B}$ and $\mathcal{B}\times\mathcal{M}(B_{e})\to\mathcal{B}$ are continuous. Given $\xi,\eta\in C_{c}(G,\mathcal{M}(B_{e}))$, we define a continuous map
$\Phi_{\xi\eta}\colon\mathcal{B}\longrightarrow\mathcal{B},\qquad b\in B_{t}\;\mapsto\;\Phi_{\xi\eta}(b)\equiv\xi\cdot b\cdot\eta:=\int_{G}\xi(s)^{*}\,b\,\eta(s^{-1}t)\,ds\;\in B_{t};$
which is linear on each fiber. With the norm
$\|\xi\|_{2}\;:=\;\Big\|\int_{G}\xi(s)^{*}\xi(s)\,ds\Big\|^{1/2}\quad\text{on }L^{2}(G,\mathcal{M}(B_{e})),$
one has $\|\Phi_{\xi\eta}(b)\|\leq\|b\|\,\|\xi\|_{2}\,\|\eta\|_{2}.$ When $\xi=\eta$ we simply write $\Phi_{\xi}=\Phi_{\xi\xi}$.
Let $\mathrm{CLB}(\mathcal{B})$ be the set of continuous maps $\Phi\colon\mathcal{B}\to\mathcal{B}$ that preserve fibers, are linear on fibers, and satisfy
$\|\Phi\|:=\sup\{\,\|\Phi(b)\|\colon b\in\mathcal{B},\,\|b\|\leq 1\}<\infty.$
A net $\{\Phi_{i}\}_{i\in I}\subset\mathrm{CLB}(\mathcal{B})$ converges to $\Phi\in\mathrm{CLB}(\mathcal{B})$ on compact slices [15, Definition 3.4] if
$\lim_{i}\|\Phi_{i}\circ f-\Phi\circ f\|_{\infty}=0\qquad\text{for all }f\in C_{c}(\mathcal{B}).$
Replacing $\|\cdot\|_{\infty}$ by $\|\cdot\|_{1}$ gives the notion of $L^{1}$-convergence on compact slices. Since ${\mathop{\rm supp}}(\Phi_{i}\circ f-\Phi\circ f)\subset{\mathop{\rm supp}}(f)$, uniform convergence on compact slices implies $L^{1}$-convergence.
Definition 3.1 (Approximation properties for Fell bundles).
Let $\mathcal{B}=\{B_{t}\}_{t\in G}$ be a Fell bundle.
  1. (1)
    $\mathcal{B}$ has the $L^{1}$-approximation property (Abadie) if there exist a constant $M>0$ and nets $\{\xi_{i}\}_{i\in I},\{\eta_{i}\}_{i\in I}\subset C_{c}(G,\mathcal{M}(B_{e}))$ such that $\|\xi_{i}\|_{2}\|\eta_{i}\|_{2}\leq M$ for all $i$, and $\{\Phi_{\xi_{i}\eta_{i}}\}_{i\in I}$ converges in the $L^{1}$-sense on compact slices to the identity map $\operatorname{id}_{\mathcal{B}}$.
  2. (2)
    $\mathcal{B}$ has the approximation property (Exel–Ng) if the nets can be chosen in $C_{c}(G,B_{e})$ and the convergence is uniform on compact slices.
    1. (a)
      $\mathcal{B}$ has the positive approximation property if, in addition, $\xi_{i}=\eta_{i}$ for all $i$.
    2. (b)
      $\mathcal{B}$ has the positive $1$-approximation property if, moreover, the constant $M$ can be chosen to be $1$.
With these definitions in place and abbreviating approximation property to AP, we have the following chain of (trivial) implications:
$\text{positive $1$-AP}\;\Rightarrow\;\text{positive AP}\;\Rightarrow\;\text{AP}\;\Rightarrow\;L^{1}\text{-AP}.$
We are now ready to state the main result of this section.
Theorem 3.2.
Let $\mathcal{B}=\{B_{t}\}_{t\in G}$ be a Fell bundle. The following conditions are equivalent:
  1. (1)
    $\mathcal{B}$ has the positive $1$-approximation property of Exel and Ng.
  2. (2)
    $\mathcal{B}$ has the positive approximation property of Exel and Ng.
  3. (3)
    $\mathcal{B}$ has the approximation property of Exel and Ng.
  4. (4)
    $\mathcal{B}$ has the $L^{1}$-approximation property of Abadie.
  5. (5)
    For every Fell bundle $\mathcal{A}=\{A_{t}\}_{t\in G}$, the diagonal tensor product $\mathcal{A}\otimes_{\max}^{d}\mathcal{B}$ has the wcp.
  6. (6)
    The canonical action $\beta$ of $G$ on $\mathbb{k}(\mathcal{B})$ is $\mathrm{C}^{*}$-amenable.
Remark 3.3.
In Definition 4.13 we introduce two more approximation properties, and in Corollary 4.17 we show that they are equivalent to the ones above.
[…]
Lemma 3.4.
Let $\mathcal{B}=\{B_{t}\}_{t\in G}$ be a Fell bundle, $\{\Phi_{i}\}_{i\in I}$ a net in $\mathrm{CLB}(\mathcal{B})$, and $\Phi\in\mathrm{CLB}(\mathcal{B})$. Assume that $M:=\sup_{i\in I}\|\Phi_{i}\|<\infty$ and that there exists $\Gamma\subseteq C_{c}(\mathcal{B})$ such that for $\Gamma^{\prime}:={\mathop{\rm span}}(\Gamma)$ one has $C_{c}(G)\Gamma^{\prime}\subseteq\Gamma^{\prime}$ and, for each $t\in G$, the set $\{f(t):f\in\Gamma^{\prime}\}$ is dense in $B_{t}$. If $\lim_{i}\|\Phi_{i}\circ f-\Phi\circ f\|_{\infty}=0$ for all $f\in C_{c}(\mathcal{B})$, then $\{\Phi_{i}\}_{i\in I}$ converges to $\Phi$ uniformly on compact slices. Moreover, if in the limit we use $\|\cdot\|_{1}$, then $\{\Phi_{i}\}_{i\in I}$ $L^{1}$-converges to $\Phi$ on compact slices.
Proof.
By [16, II 14.6], $\Gamma^{\prime}$ is dense in $C_{c}(\mathcal{B})$ for the inductive limit topology, hence also dense for both $\|\cdot\|_{\infty}$ and $\|\cdot\|_{1}$. Let $\|\cdot\|$ denote either of these norms. For $f\in C_{c}(\mathcal{B})$, $n\in\mathbb{N}$, $z_{1},\ldots,z_{n}\in\mathbb{C}$, $g_{1},\ldots,g_{n}\in\Gamma$, and $i\in I$, one has
$\|\Phi_{i}\circ f-\Phi\circ f\|\;\leq\;(M+\|\Phi\|)\Big\|f-\sum_{j=1}^{n}z_{j}g_{j}\Big\|+\sum_{j=1}^{n}|z_{j}|\;\|\Phi_{i}\circ g_{j}-\Phi\circ g_{j}\|.$
A standard approximation argument then yields $\lim_{i}\|\Phi_{i}\circ f-\Phi\circ f\|=0$. ∎
We continue the proof of Theorem 3.2 with the following
Lemma 3.5.
Let $\mathcal{A}=\{A_{t}\}_{t\in G}$ and $\mathcal{B}=\{B_{t}\}_{t\in G}$ be Fell bundles. If at least one of them has the $L^{1}$-approximation property, then so does $\mathcal{A}\otimes_{\max}^{d}\mathcal{B}$.
Proof.
Assume $\mathcal{A}$ has the $L^{1}$-approximation property. For $\xi,\eta\in C_{c}(G,\mathcal{M}(A_{e}))$ define
$\tilde{\xi}(t):=\xi(t)\otimes 1\in\mathcal{M}(A_{e}\otimes_{\max}B_{e})\quad(t\in G).$
Then $\tilde{\xi}\in C_{c}(G,\mathcal{M}(A_{e}\otimes_{\max}B_{e}))$ and for $c\in A_{t}$, $b\in B_{t}$ we have
$\Phi_{\tilde{\xi}\tilde{\eta}}(c\otimes b)=\int_{G}\xi(s)^{*}c\eta(s^{-1}t)\otimes b\,ds=\Phi_{\xi\eta}(c)\otimes b.$
Moreover, $\|\tilde{\xi}\|=\|\xi\|$ in $L^{2}(G,\mathcal{M}(A_{e}\otimes_{\max}B_{e}))$. For $f\in C_{c}(\mathcal{A})$ and $g\in C_{c}(\mathcal{B})$ one checks
$\Phi_{\tilde{\xi}\tilde{\eta}}(f\oslash^{d}g)=(\Phi_{\xi\eta}\circ f)\oslash^{d}g,$
and hence
$\|\Phi_{\tilde{\xi}\tilde{\eta}}(f\oslash^{d}g)-f\oslash^{d}g\|_{1}\;\leq\;\|\Phi_{\xi\eta}\circ f-f\|_{1}\,\|g\|_{\infty}.$
Let $\{\xi_{i}\},\{\eta_{i}\}\subseteq C_{c}(G,\mathcal{M}(A_{e}))$ witness the $L^{1}$-approximation property of $\mathcal{A}$. Then $\|\tilde{\xi}_{i}\|\|\tilde{\eta}_{i}\|=\|\xi_{i}\|\|\eta_{i}\|\leq M$, and for all $f\in C_{c}(\mathcal{A})$, $g\in C_{c}(\mathcal{B})$,
$\lim_{i}\|\Phi_{\tilde{\xi}_{i}\tilde{\eta}_{i}}(f\oslash^{d}g)-f\oslash^{d}g\|_{1}=0.$
Taking $\Gamma:=\{f\oslash^{d}g:f\in C_{c}(\mathcal{A}),\,g\in C_{c}(\mathcal{B})\}$ and applying Lemma 3.4, we conclude that $\mathcal{A}\otimes_{\max}^{d}\mathcal{B}$ has the $L^{1}$-approximation property. ∎
[…]
Lemma 3.6.
Let $\mathcal{A}=\{A_{t}\}_{t\in G}$, $\mathcal{B}=\{B_{t}\}_{t\in G}$, and $\mathcal{C}=\{C_{t}\}_{t\in G}$ be Fell bundles. If $\mathcal{A}$ is the semidirect product bundle of a $\mathrm{C}^{*}$-action and $\mathcal{B}$ and $\mathcal{C}$ are equivalent, then $\mathcal{A}\otimes_{\max}^{d}\mathcal{B}$ and $\mathcal{A}\otimes_{\max}^{d}\mathcal{C}$ are equivalent in the same sense.
Proof.
Let $\mathcal{X}=\{X_{t}\}_{t\in G}$ be a $\mathcal{B}$$\mathcal{C}$ equivalence bundle, and let $\mathbb{L}(\mathcal{X})=\{\mathbb{L}(X_{t})\}_{t\in G}$ denote its linking bundle as in [6, Section 3.1]. Since $B_{e}$ and $C_{e}$ are hereditary $\mathrm{C}^{*}$-subalgebras of $\mathbb{L}(X_{e})$, we have
$A_{e}\otimes_{\max}B_{e}\;\subseteq\;A_{e}\otimes_{\max}\mathbb{L}(X_{e})\;\supseteq\;A_{e}\otimes_{\max}C_{e},$
so that $\mathcal{A}\otimes_{\max}^{d}\mathcal{B}$ and $\mathcal{A}\otimes_{\max}^{d}\mathcal{C}$ can be naturally regarded as Fell subbundles of $\mathcal{A}\otimes_{\max}^{d}\mathbb{L}(\mathcal{X})$.
Moreover, $\mathcal{X}$ sits as a Banach subbundle of $\mathbb{L}(\mathcal{X})$, and hence
$\mathcal{A}\otimes_{\max}^{d}\mathcal{X}:=\Big\{\overline{\mathop{\rm span}}\{a\otimes x:a\in A_{t},\;x\in X_{t}\}\Big\}_{t\in G}=\{A_{t}\otimes_{\max}X_{t}\}_{t\in G}$
is a Banach subbundle of $\mathcal{A}\otimes_{\max}^{d}\mathbb{L}(\mathcal{X})$. By restricting the multiplication and involution from $\mathcal{A}\otimes_{\max}^{d}\mathbb{L}(\mathcal{X})$, this subbundle acquires a natural structure of an $\mathcal{A}\otimes_{\max}^{d}\mathcal{B}$$\mathcal{A}\otimes_{\max}^{d}\mathcal{C}$ equivalence bundle.
To check fullness, set
$S:=\overline{\mathop{\rm span}}\Big\{{}_{\mathcal{A}\otimes_{\max}^{d}\mathcal{B}}\langle a\otimes x,\;b\otimes y\rangle:a\in A_{t},\,b\in A_{t^{-1}},\,x\in X_{t},\,y\in X_{t^{-1}},\,t\in G\Big\}.$
A direct computation shows that
$S=\overline{\mathop{\rm span}}\{ab^{*}\otimes{}_{\mathcal{B}}\langle x,y\rangle:a\in A_{t},\;b\in A_{t^{-1}},\;x\in X_{t},\;y\in X_{t^{-1}},\;t\in G\}.$
Now suppose $\mathcal{A}$ is the semidirect product bundle associated with a $\mathrm{C}^{*}$-action $\alpha$ of $G$ on a $\mathrm{C}^{*}$-algebra $A$. Given $z\delta_{t}\in A\delta_{t}\cong A_{t}$, one can write $z=z_{1}z_{2}$ with $z_{1},z_{2}\in A$, and then
$z\delta_{e}=(z_{1}\delta_{t})\,(\alpha_{t^{-2}}(z_{2})^{*}\delta_{t^{-1}})^{*}.$
This implies that $S=A_{e}\otimes_{\max}B_{e}$. A similar computation shows that
$\overline{\mathop{\rm span}}\Big\{\langle a\otimes x,b\otimes y\rangle_{\mathcal{A}\otimes_{\max}^{d}\mathcal{C}}:a\in A_{t},\,b\in A_{t^{-1}},\,x\in X_{t},\,y\in X_{t^{-1}},\,t\in G\Big\}=A_{e}\otimes_{\max}C_{e}.$
Thus conditions (7R) and (7L) of [6, Definitions 2.2] are satisfied. All the remaining conditions follow automatically from the fact that the operations on $\mathcal{A}\otimes_{\max}^{d}\mathcal{X}$ were inherited from the Fell bundle $\mathcal{A}\otimes_{\max}^{d}\mathbb{L}(\mathcal{X})$. Hence $\mathcal{A}\otimes_{\max}^{d}\mathcal{B}$ and $\mathcal{A}\otimes_{\max}^{d}\mathcal{C}$ are equivalent. ∎
Corollary 3.7.
[…]
Proof.
Let $\alpha$ be a $\mathrm{C}^{*}$-action of $G$ on a $\mathrm{C}^{*}$-algebra $A$, and let $\mathcal{A}$ denote its semidirect product bundle. Recall that $\beta$ is the natural action of $G$ on $\mathbb{k}(\mathcal{B})$, and let $\mathcal{C}$ be its semidirect product bundle.
By [5, Theorem 3.4], $\mathcal{B}$ is equivalent to $\mathcal{C}$, and hence, by Lemma 3.6, $\mathcal{A}\otimes_{\max}^{d}\mathcal{B}$ is equivalent to $\mathcal{A}\otimes_{\max}^{d}\mathcal{C}$. Since $\mathcal{A}\otimes_{\max}^{d}\mathcal{B}$ has the wcp (equivalently, it is $\mathop{\rm r}$}-amenable in the sense of [6, Definition 4.9]), [6, Corollary 4.20] implies that $\mathcal{A}\otimes_{\max}^{d}\mathcal{C}$ also has the wcp.
Now, $\mathcal{A}\otimes_{\max}^{d}\mathcal{C}$ is precisely the semidirect product bundle of the diagonal action $\alpha\otimes_{\max}^{d}\beta$. Thus, the reduced representation of $\mathcal{A}\otimes_{\max}^{d}\mathcal{C}$ is the canonical map
$C^{*}(\mathcal{A}\otimes_{\max}^{d}\mathcal{C})\;=\;(A\otimes_{\max}^{d}\mathbb{k}(\mathcal{B}))\rtimes G\;\longrightarrow\;(A\otimes_{\max}^{d}\mathbb{k}(\mathcal{B}))\rtimes_{{\mathop{\rm r}}}G.$
Therefore, $\alpha\otimes_{\max}^{d}\beta$ has the wcp, and the desired conclusion follows from Theorem 2.3. ∎
[…]
Lemma 3.8.
Let $\mathcal{A}=\{A_{t}\}_{t\in G}$ be a Fell subbundle of $\mathcal{B}=\{B_{t}\}_{t\in G}$ such that $A_{e}$ is hereditary in $B_{e}$. Suppose there exists a net $\{\xi_{i}\}_{i\in I}\subseteq C_{c}(G,\mathcal{M}(B_{e}))$ such that $\{\Phi_{\xi_{i}}\}_{i\in I}$ converges to $\operatorname{id}_{\mathcal{B}}$ on compact slices (as is the case when $\mathcal{B}$ has the positive approximation property of Exel and Ng). Then $\mathcal{A}$ has the positive $1$-approximation property.
Proof.
It suffices to show that for every finite set $S\subseteq C_{c}(\mathcal{A})$ and $\varepsilon>0$ there exists $\eta\in C_{c}(G,A_{e})$ with $\|\eta\|_{2}\leq 1$ such that
$\|\Phi_{\eta}(b)-b\|<\varepsilon\qquad\text{for all }b\in S(G):=\bigcup_{f\in S}f(G).$
Fix such $S$ and $\varepsilon$. Since every approximate unit of $A_{e}$ is a strong approximate unit for $\mathcal{A}$, there exists a positive element $a\in A_{e}$ with $\|a\|<1$ such that
$\|aba-b\|<\tfrac{\varepsilon}{4}\qquad\forall\,b\in S(G).$
It follows that
$\|a^{2}ba^{2}-b\|\;\leq\;\|a(aba-b)a\|+\|aba-b\|\;<\;\tfrac{\varepsilon}{2}.$
Choose $f\in C_{c}(\mathcal{A})$ with $f(e)=a^{2}$, and for each $g\in S$ define $g^{\prime}\in C_{c}(\mathcal{A})$ by $g^{\prime}(s)=ag(s)a$. Set $S^{\prime}:=\{g^{\prime}:g\in S\}\cup\{f\}$. By assumption, there exists $\xi\in C_{c}(G,\mathcal{M}(B_{e}))$ such that
$\|\Phi_{\xi}(b)-b\|<\min\{\tfrac{\varepsilon}{2},\,1-\|a\|^{2}\}\qquad\forall\,b\in S^{\prime}(G).$
Now define $\eta\in C_{c}(G,A_{e})$ by $\eta(t)=a\xi(t)a$. Then for all $t\in G$ and $b\in B_{t}$,
$\Phi_{\eta}(b)=\int_{G}a\xi(s)^{*}\,aba\,\xi(s)\,a\,ds=a\,\Phi_{\xi}(aba)\,a.$
For $b\in S(G)$ set $b^{\prime}:=aba\in S^{\prime}(G)$. Since $\|b^{\prime}-b\|<\varepsilon/2$, we obtain
$\|\Phi_{\eta}(b)-b\|\;\leq\;\|a\Phi_{\xi}(b^{\prime})a-a^{2}ba^{2}\|+\|a^{2}ba^{2}-b\|\;<\;\|a\|^{2}\|\Phi_{\xi}(b^{\prime})-b^{\prime}\|+\tfrac{\varepsilon}{2}\;<\;\varepsilon.$
Finally, note that
$\langle\eta,\eta\rangle=a\Phi_{\xi}(a^{2})a,$
and since $a^{2}\in S^{\prime}(G)$, we have
$\|\eta\|^{2}\;\leq\;\|\Phi_{\xi}(a^{2})\|\;\leq\;\|\Phi_{\xi}(a^{2})-a^{2}\|+\|a^{2}\|<(1-\|a\|^{2})+\|a\|^{2}=1.$
Thus $\eta$ has the desired properties. ∎
Lemma 3.9.
The positive $1$-approximation property of Exel and Ng is preserved under strong equivalence of Fell bundles.
Proof.
Let $\mathcal{A}=\{A_{t}\}_{t\in G}$ and $\mathcal{B}=\{B_{t}\}_{t\in G}$ be Fell bundles and let $\mathcal{X}=\{X_{t}\}_{t\in G}$ be a strong $\mathcal{A}\mathcal{B}$ equivalence bundle with inner products $_{\mathcal{A}}\langle\cdot$$\cdot\rangle$ and $\langle\cdot,\cdot\rangle_{\mathcal{B}}$. Since strong equivalence is an equivalence relation [5, Theorem A.1], we may assume, without loss of generality, that $\mathcal{A}$ has the positive $1$-approximation property. Let $\{\xi_{i}\}_{i\in I}\subseteq C_{c}(G,A_{e})$ be a net as in the definition of this property.
Let $\Gamma$ denote the set of cross-sections of $\mathcal{B}$ of the form $r\mapsto\langle x,f(r)y\rangle_{\mathcal{B}}$, with $x,y\in X_{e}$ and $f\in C_{c}(\mathcal{A})$. Clearly $C_{c}(G)\Gamma\subseteq\Gamma$, hence $\Gamma^{\prime}:={\mathop{\rm span}}\Gamma$ satisfies $C_{c}(G)\Gamma^{\prime}\subseteq\Gamma^{\prime}$. For all $r\in G$, using [5, Notation 2.1] and the definition of strong equivalence, one has
$B_{r}=B_{r}B_{r}^{*}B_{r}=B_{r}\langle X_{r},X_{r}\rangle_{\mathcal{B}}=\langle X_{r}B_{r^{-1}},X_{r}\rangle_{\mathcal{B}}\subseteq\langle X_{e},{}_{\mathcal{A}}\langle X_{r},X_{r}\rangle X_{r}\rangle_{\mathcal{B}}\subseteq\langle X_{e},A_{r}X_{e}\rangle_{\mathcal{B}},$
so that $\{u(r):u\in\Gamma^{\prime}\}={\mathop{\rm span}}\{\langle x,ay\rangle_{\mathcal{B}}:x,y\in X_{e},\,a\in A_{r}\}$ is dense in $B_{r}$.
By combining the Axiom of Choice with Lemma 3.4, we get that it suffices to show that for every finite $U\subseteq\Gamma$ and $\varepsilon>0$, the set
$S(U,\varepsilon):=\{\eta\in C_{c}(G,B_{e}):\|\eta\|_{2}\leq 1,\ \|\Phi_{\eta}(f)-f\|_{\infty}\leq\varepsilon\ \forall f\in U\}$
is nonempty.
Fix $U\subseteq\Gamma$ finite and $\varepsilon>0$. Since $X_{e}$ is an $A_{e}$$B_{e}$ equivalence module and every approximate unit of $B_{e}$ is a strong approximate unit of $\mathcal{B}$, [6, Lemma 2.8] yields the existence of $x_{1},\dots,x_{n}\in X_{e}$ such that, for $b:=\sum_{k=1}^{n}\langle x_{k},x_{k}\rangle_{\mathcal{B}}$, we have $0<\|b\|<1$ and
$\|bu(r)b-u(r)\|<\varepsilon/2\quad\forall(r,u)\in G\times U.$
By assumption, for sufficiently large $i\in I$ we have
$\|\Phi_{\xi_{i}}({}_{\mathcal{A}}\langle x_{k}u(r),x_{l}\rangle)-{}_{\mathcal{A}}\langle x_{k}u(r),x_{l}\rangle\|<\tfrac{\varepsilon}{2n^{2}}$
for all $k,l=1,\dots,n$, all $r\in G$, and all $u\in U$.
Define $\eta\in C_{c}(G,B_{e})$ by $\eta(s)=\sum_{k=1}^{n}\langle x_{k},\xi_{i}(s)x_{k}\rangle_{\mathcal{B}}$. Then
$\langle\eta,\eta\rangle=\int_{G}\sum_{k,l=1}^{n}\langle\xi_{i}(s)x_{k},{}_{\mathcal{A}}\langle x_{k},x_{l}\rangle\xi_{i}(s)x_{l}\rangle_{\mathcal{B}}\,ds.$
Viewing $X^{n}$ as a $\mathbb{M}_{n}(A_{e})$$B_{e}$ Hilbert bimodule, one checks
$\sum_{k,l}\langle\xi_{i}(s)x_{k},{}_{\mathcal{A}}\langle x_{k},x_{l}\rangle\xi_{i}(s)x_{l}\rangle_{\mathcal{B}}\leq\sum_{k=1}^{n}\langle x_{k},\xi_{i}(s)^{*}\xi_{i}(s)x_{k}\rangle_{\mathcal{B}},$
whence
$\langle\eta,\eta\rangle\leq\|\xi_{i}\|_{2}^{2}\,b,$
and therefore $\|\eta\|_{2}\leq\|\xi_{i}\|_{2}\leq 1$.
Now let $(r,u)\in G\times U$, say $u(r)=\langle x_{u},f_{u}(r)y_{u}\rangle_{\mathcal{B}}$. A direct computation gives
$\Phi_{\eta}(u(r))=\sum_{k,l=1}^{n}\langle x_{k},\Phi_{\xi_{i}}({}_{\mathcal{A}}\langle x_{k}u(r),x_{l}\rangle)x_{l}\rangle_{\mathcal{B}}.$
Combining this with $\|x_{k}\|^{2}=\|\langle x_{k},x_{k}\rangle_{\mathcal{B}}\|\leq\|b\|<1$, we obtain
$\|\Phi_{\eta}(u(r))-u(r)\|$ $\leq\|bu(r)b-u(r)\|+\Big\|\sum_{k,l}\langle x_{k},(\Phi_{\xi_{i}}-\operatorname{id})({}_{\mathcal{A}}\langle x_{k}u(r),x_{l}\rangle)x_{l}\rangle_{\mathcal{B}}\Big\|$
$<\tfrac{\varepsilon}{2}+\sum_{k,l=1}^{n}\|x_{k}\|\|x_{l}\|\,\tfrac{\varepsilon}{2n^{2}}$
$<\tfrac{\varepsilon}{2}+\tfrac{\varepsilon}{2}=\varepsilon.$
Thus $S(U,\varepsilon)\neq\emptyset$, and the lemma follows. ∎
Proof of Theorem 3.2.
[…]
Definition 3.10.
A Fell bundle $\mathcal{B}$ is called $\mathrm{C}^{*}$-amenable if it satisfies any (and hence all) of the equivalent conditions listed in Theorem 3.2.
A $\mathrm{C}^{*}$-partial action is called $\mathrm{C}^{*}$-amenable if its associated semidirect product Fell bundle is $\mathrm{C}^{*}$-amenable.
Remark 3.11.
In [15], Exel and Ng defined a Fell bundle to be amenable if the canonical map
$\lambda^{\mathcal{B}}\colon C^{*}(\mathcal{B})\longrightarrow C^{*}_{\mathop{\rm r}}(\mathcal{B})$
is injective. Their notion of amenability therefore coincides with what we call in this paper the weak containment property (wcp). To avoid ambiguity, we reserve the term amenability for the stronger concept introduced in Definition 3.10, which is expressed in terms of approximation properties and aligns more closely with the classical notion of amenability for group actions in the sense of Anantharaman–Delaroche.
Remark 3.12.
Assume $\alpha$ is a $\mathrm{C}^{*}$-partial action of $G$ on $A$, let $\mathcal{B}$ denote its semidirect product bundle and let $\beta$ be the natural action of $G$ on $\mathbb{k}(\mathcal{B})$. Then $\beta$ is a Morita enveloping action of $\alpha$ [3, Theorem 6.1], and Morita enveloping actions are unique up to Morita equivalence [3, Proposition 6.3]. In particular, if $\alpha$ is already a $\mathrm{C}^{*}$-action, then $\beta$ is Morita equivalent to $\alpha$. Since $\mathrm{C}^{*}$-amenability is invariant under Morita equivalence of $\mathrm{C}^{*}$-actions [11, Proposition 3.20], Definition 3.10 is consistent with the usual notion of amenability for $\mathrm{C}^{*}$-actions. More generally, a $\mathrm{C}^{*}$-partial action $\alpha$ is $\mathrm{C}^{*}$-amenable if and only if any (equivalently, all) of its Morita enveloping actions are $\mathrm{C}^{*}$-amenable.
Each of the equivalent conditions in Theorem 3.2 can be reformulated in terms of $\alpha$. For instance, for $\mathrm{C}^{*}$-actions the positive approximation property of Exel and Ng coincides with the AP of [11, Definition 4.3]. (To get the equivalence, one may apply Lemma 3.4 with $\Gamma\subset C_{c}(\mathcal{B})$ consisting of sections of the form $t\mapsto f(t)a\delta_{t}$, where $f\in C_{c}(G)$ and $a\in A$.) On the other hand, some other approximation properties – such as the QAP of [11, Definition 4.1] or the weak versions of the AP and the QAP – do not correspond to the translated versions of the properties in Theorem 3.2.
Remark 3.13.
Our definition of amenability for Fell bundles naturally induces a notion of amenability for twisted $\mathrm{C}^{*}$-actions, and even for twisted $\mathrm{C}^{*}$-partial actions: given such a twisted partial action, we assign to it the corresponding (semidirect product) Fell bundle, and we then say that the twisted partial action is amenable if the associated Fell bundle is $\mathrm{C}^{*}$-amenable.
More generally, given any construction that associates to an object $X$ a Fell bundle $F(X)$, it is natural to declare $X$ $\mathrm{C}^{*}$-amenable whenever $F(X)$ is $\mathrm{C}^{*}$-amenable.
There are a number of interesting consequences we think worth mentioning.
Corollary 3.14 (Lemma 3.5).
If $\mathcal{A}$ and $\mathcal{B}$ are Fell bundles over the same group and one of them is $\mathrm{C}^{*}$-amenable, then $\mathcal{A}\otimes_{\max}^{d}\mathcal{B}$ is $\mathrm{C}^{*}$-amenable. In particular, amenability is preserved under taking diagonal tensor products with arbitrary Fell bundles.
The positive AP can be weakened to an equivalent (unbounded) condition.
Corollary 3.15.
A Fell bundle $\mathcal{B}=\{B_{t}\}_{t\in G}$ is $\mathrm{C}^{*}$-amenable if and only if there exists a net $\{\xi_{i}\}_{i\in I}\subseteq C_{c}(G,\mathcal{M}(B_{e}))$ such that for all $f\in C_{c}(\mathcal{B})$,
$\lim_{i}\|\Phi_{\xi_{i}}\circ f-f\|_{\infty}=0.$
Proof.
The existence of such a net is clearly weaker than the positive $1$-approximation property. For the converse, apply Lemma 3.8 with $\mathcal{A}=\mathcal{B}$. ∎
Using the corollary above one could give a (quite long but) direct proof of the following.
Corollary 3.16.
The $\mathrm{C}^{*}$-amenability of Fell bundles (respectively, $\mathrm{C}^{*}$-partial actions) is preserved under equivalence of Fell bundles (of $\mathrm{C}^{*}$-partial actions, see [3, Definition 4.4]).
Proof.
Let $\mathcal{A}$ and $\mathcal{B}$ be weakly equivalent Fell bundles over the same group $G$, and let $\alpha$ and $\beta$ denote the respective actions on the $\mathrm{C}^{*}$-algebras of kernels. By [5, Theorem 4.2], $\alpha$ is Morita equivalent to $\beta$. Combining this with [11, Proposition 3.20] and Theorem 3.2, we obtain the chain of equivalences
[…]
The claim for $\mathrm{C}^{*}$-partial actions follows since, if two $\mathrm{C}^{*}$-partial actions are Morita equivalent, then their semidirect product bundles are equivalent [6, Section 2.2.2]. ∎
Corollary 3.17.
If $\mathcal{A}=\{A_{t}\}_{t\in G}$ is a Fell subbundle of $\mathcal{B}=\{B_{t}\}_{t\in G}$ such that $A_{e}$ is hereditary in $B_{e}$, and if $\mathcal{B}$ is $\mathrm{C}^{*}$-amenable, then $\mathcal{A}$ is $\mathrm{C}^{*}$-amenable.
Proof.
Since $\mathcal{B}$ is $\mathrm{C}^{*}$-amenable, it has the positive approximation property of Exel and Ng. By Lemma 3.8, this implies that any Fell subbundle $\mathcal{A}\subseteq\mathcal{B}$ with $A_{e}$ hereditary in $B_{e}$ inherits the positive $1$-approximation property. Hence $\mathcal{A}$ is $\mathrm{C}^{*}$-amenable by Theorem 3.2. ∎
The next result provides a systematic way of producing amenable $\mathrm{C}^{*}$-partial actions: it shows that restricting an amenable $\mathrm{C}^{*}$-action to a (possibly non-invariant) $\mathrm{C}^{*}$-ideal yields an amenable $\mathrm{C}^{*}$-partial action. Moreover, up to Morita equivalence of $\mathrm{C}^{*}$-partial actions, this construction accounts for all amenable $\mathrm{C}^{*}$-partial actions. Indeed, by [3], every $\mathrm{C}^{*}$-partial action is Morita equivalent to one that is globalizable, i.e. the restriction of a global $\mathrm{C}^{*}$-action.
Corollary 3.18.
Let $\alpha=(\{A_{t}\}_{t\in G},\{\alpha_{t}\}_{t\in G})$ be a $\mathrm{C}^{*}$-action that is $\mathrm{C}^{*}$-amenable, and let $I$ be a $\mathrm{C}^{*}$-ideal of $A$ (not necessarily invariant). Then the restricted $\mathrm{C}^{*}$-partial action $\alpha|_{I}$ (as in [6, Proposition 4.2]) is also $\mathrm{C}^{*}$-amenable.
Proof.
The semidirect product bundle $\mathcal{B}_{\alpha|_{I}}$ of $\alpha|_{I}$ is a hereditary Fell subbundle of $\mathcal{B}_{\alpha}$. The claim then follows directly from Corollary 3.17. ∎
Remark 3.19.
The corollary above reflects a more general phenomenon for Fell bundles. By [6, 5], every Fell bundle is strongly equivalent to one arising from a $\mathrm{C}^{*}$-partial action, and every such partial action is Morita equivalent to the restriction of a global $\mathrm{C}^{*}$-action. Since $\mathrm{C}^{*}$-amenability is preserved under strong equivalence of Fell bundles and under passage to enveloping (global) actions, it follows that amenability of Fell bundles and $\mathrm{C}^{*}$-partial actions is entirely governed by the amenability of global $\mathrm{C}^{*}$-actions.
3.1. Amenability and nuclearity
In this section we relate the amenability of a Fell bundle to the nuclearity of its cross-sectional $\mathrm{C}^{*}$-algebra. We start by giving an alternative proof of a Theorem of Abadie (see [1, Corollary 4.13] or [2, Corolário 3.29]). Fix a Fell bundle $\mathcal{B}=\{B_{t}\}_{t\in G}$.
Theorem 3.20.
If $\mathcal{B}$ is $\mathrm{C}^{*}$-amenable and $B_{e}$ is nuclear, then $C^{*}(\mathcal{B})$ is nuclear.
Proof.
Let $\beta$ be the natural action of $G$ on $\mathbb{k}(\mathcal{B})$ and $\mathcal{C}$ its semidirect product bundle. Then $\mathcal{B}$ is equivalent to $\mathcal{C}$ and $C^{*}(\mathcal{C})=\mathbb{k}(\mathcal{B})\rtimes_{\beta}G$. By [3, Corollary 5.2], $\mathbb{k}(\mathcal{B})$ is nuclear. Since $\beta$ is $\mathrm{C}^{*}$-amenable, [11, Theorem 7.2] implies that $C^{*}(\mathcal{C})$ is nuclear. Using [6, Theorem 4.5] we get that $C^{*}(\mathcal{B})$ is nuclear. ∎
Theorem 3.21.
Let $\mathcal{B}$ be a Fell bundle over a locally compact group $G$.
  1. (1)
    If $C^{*}_{\mathop{\rm r}}(\mathcal{B})$ is nuclear, then $B_{e}$ is nuclear.
  2. (2)
    If $G$ inner amenable and $C^{*}_{\mathop{\rm r}}(\mathcal{B})$ is nuclear, then $\mathcal{B}$ is $\mathrm{C}^{*}$-amenable.
Proof.
By [6, Proposition 4.13], $C^{*}_{\mathop{\rm r}}(\mathcal{B})$ is Morita equivalent to $C^{*}_{\mathop{\rm r}}(\mathcal{C})=\mathbb{k}(\mathcal{B})\rtimes_{{\mathop{\rm r}}\beta}G$. Hence, $\mathbb{k}(\mathcal{B})\rtimes_{{\mathop{\rm r}}\beta}G$ is nuclear and [18, Theorem 3.4] implies that $\mathbb{k}(\mathcal{B})$ is nuclear (unfortunately, in that article nuclear $\mathrm{C}^{*}$-algebras are called amenable). Corollary 5.2 of [3] implies $B_{e}$ is nuclear. Finally, if $G$ is inner amenable, then [19, Corollary 6.6] implies $\beta$ is $\mathrm{C}^{*}$-amenable. ∎
4. An approximation property of Bédos and Conti
A new approximation property for Fell bundles over discrete groups was recently introduced in [9]. In this section we extend it to Fell bundles over locally compact groups and prove that it is equivalent to $\mathrm{C}^{*}$-amenability. We then use this equivalence to establish certain permanence properties of $\mathrm{C}^{*}$-amenability.
Our approach follows the definitions and dilation theorems of [12], where the setting is that of Fell bundles over discrete groups. Throughout, let $\mathcal{B}=\{B_{t}\}_{t\in G}$ be a Fell bundle over a locally compact group $G$.
Recall from [6, Lemma 2.8] that for each $n\in\mathbb{N}$ and $t=(t_{1},\ldots,t_{n})\in G^{n}$, the set
$\mathbb{M}_{t}(\mathcal{B}):=\bigl\{(b_{i,j})_{i,j=1}^{n}:b_{i,j}\in B_{t_{i}^{-1}t_{j}}\text{ for all }i,j\bigr\}$
is a $\mathrm{C}^{*}$-algebra under the usual matrix operations. Given a *-representation $T\colon\mathcal{B}\to\mathbb{B}(X)$, one obtains a natural *-representation
$T^{t}\colon\mathbb{M}_{t}(\mathcal{B})\longrightarrow\mathbb{M}(\mathbb{B}(X))=\mathbb{B}(X^{n}),\qquad(b_{i,j})\longmapsto(T_{b_{i,j}})_{i,j=1}^{n}.$
Moreover, if $T|_{B_{e}}$ is faithful, then by the $\mathrm{C}^{*}$-identity each restriction $T|_{B_{t}}$ is isometric, and hence $T^{t}$ is faithful.
The following definition extends [12, Definition A.2] to Fell bundles over locally compact groups.
Definition 4.1.
A map $\phi\colon\mathcal{B}\to\mathbb{B}(X)$ is said to be completely positive (cp) if:
  1. (1)
    $X$ is a Hilbert space and the restriction $\phi|_{B_{t}}$ is linear for each fiber $B_{t}$.
  2. (2)
    For every $n\in\mathbb{N}$ and $t=(t_{1},\dots,t_{n})\in G^{n}$, the map
    $\phi^{t}\colon\mathbb{M}_{t}(\mathcal{B})\to\mathbb{M}_{n}(\mathbb{B}(X)),\qquad(b_{i,j})\mapsto(\phi(b_{i,j})),$
    is positive.
  3. (3)
    For all $x,y\in X$, the function $b\mapsto\langle x,\phi(b)y\rangle$ is continuous on $\mathcal{B}$.
The support of $\phi$ is
${\mathop{\rm supp}}(\phi)\;:=\;\overline{\{t\in G:\phi(B_{t})\neq\{0\}\}}.$
If $\|\phi|_{B_{e}}\|\leq 1$, then $\phi$ is called ccp (contractive completely positive).
Remark 4.2.
A basic way to construct a cp map on $\mathcal{B}$ is as follows. Take a *-representation $T\colon\mathcal{B}\to\mathbb{B}(Y)$, a Hilbert space $X$, and an operator $V\in\mathbb{B}(X,Y)$. Then the map $S\colon\mathcal{B}\to\mathbb{B}(X)$ defined by
$S_{b}:=V^{*}T_{b}V,\qquad b\in\mathcal{B},$
is completely positive, since for every $n\in\mathbb{N}$ and $t\in G^{n}$ we have
$S^{t}=(V^{n})^{*}\,T^{t}\,V^{n},$
where $V^{n}\colon X^{n}\to Y^{n}$ is the diagonal amplification of $V$.
Theorem A.4 and Remark A.5 of [12] give the following.
Theorem 4.3 (Stinespring’s Dilation Theorem).
If $\phi\colon\mathcal{B}\to\mathbb{B}(X)$ is cp, then there exist a *-representation $T\colon\mathcal{B}\to\mathbb{B}(Y)$ and an operator $V\in\mathbb{B}(X,Y)$ such that
$Y=\overline{\mathop{\rm span}}\,T(\mathcal{B})VX\qquad\text{and}\qquad\phi(b)=V^{*}T_{b}V\quad\text{for all }b\in\mathcal{B}.$
Moreover, any such $T$ is nondegenerate, and for every approximate unit $\{b_{i}\}_{i\in I}$ we have
$V^{*}V=\text{{sot}-}\lim_{i}\phi(b_{i}),\qquad\|V\|^{2}=\|\phi|_{B_{e}}\|=\lim_{i}\|\phi(b_{i})\|.$
Proof.
Let $G^{d}$ be the group $G$ with the discrete topology, and set $\mathcal{B}^{d}:=\{B_{t}\}_{t\in G^{d}}$. As sets, $\mathcal{B}$ and $\mathcal{B}^{d}$ coincide. By [12, Theorem A.4], there exist a *-representation $T^{d}\colon\mathcal{B}^{d}\to\mathbb{B}(Y)$ and an operator $V\in\mathbb{B}(X,Y)$ such that $Y=\overline{\mathop{\rm span}}\,T^{d}(\mathcal{B}^{d})VX$ and $\phi(b)=V^{*}T^{d}_{b}V$ for all $b\in\mathcal{B}^{d}$.
Define $T\colon\mathcal{B}\to\mathbb{B}(Y)$ by $T_{b}:=T^{d}_{b}$. To check that $T$ is wot-continuous, it suffices (since $Y=\overline{\mathop{\rm span}}\,T^{d}(\mathcal{B}^{d})VX$) to verify that for all $c,d\in\mathcal{B}$ and $x,y\in X$, the function
[…]
is continuous in $b$. This follows directly from the continuity of $\phi$. ∎
Corollary 4.4 (Integrated form of cp maps).
If $\phi\colon\mathcal{B}\to\mathbb{B}(X)$ is cp, there exists a unique cp map $\tilde{\phi}\colon C^{*}(\mathcal{B})\to\mathbb{B}(X)$ such that
$\langle x,\tilde{\phi}(f)y\rangle=\int_{G}\langle x,\phi(f(t))y\rangle\,dt,\qquad f\in C_{c}(\mathcal{B}),\;x,y\in X.$
Moreover, $\|\tilde{\phi}\|=\|\phi|_{B_{e}}\|$.
Proof.
By Stinespring’s theorem, $\phi(b)=V^{*}T_{b}V$ for a *-representation $T\colon\mathcal{B}\to\mathbb{B}(Y)$ and $V\in\mathbb{B}(X,Y)$. Let $\tilde{T}\colon C^{*}(\mathcal{B})\to\mathbb{B}(Y)$ be the integrated form of $T$, and set $\tilde{\phi}(f)=V^{*}\tilde{T}(f)V$. This $\tilde{\phi}$ is cp, satisfies the required identity, is unique by density of $C_{c}(\mathcal{B})$, and $\|\tilde{\phi}\|=\|V\|^{2}=\|\phi|_{B_{e}}\|$. ∎
Remark 4.5.
Let $A,B,C$ be $\mathrm{C}^{*}$-algebras and $\pi\colon A\to B$ a surjective *-homomorphism. Suppose $\psi\colon A\to C$ is (completely) positive. Then there exists a (completely) positive map $\hat{\psi}\colon B\to C$ with $\hat{\psi}\circ\pi=\psi$ iff $\ker(\pi)\subseteq\ker(\psi)$.
The forward implication is clear. Conversely, if $\ker(\pi)\subseteq\ker(\psi)$, then surjectivity of $\pi$ yields the existence of a unique linear map $\hat{\psi}\colon B\to C$ with $\hat{\psi}\circ\pi=\psi$. For $b\in B^{+}$, take $a\in A^{+}$ with $\pi(a)=b$, so $\hat{\psi}(b)=\psi(a)\geq 0$, proving positivity. If $\psi$ is cp, then for all $n$ the map
$\hat{\psi}^{(n)}\colon\mathbb{M}_{n}(B)\to\mathbb{M}_{n}(C),\qquad(b_{i,j})\mapsto(\hat{\psi}(b_{i,j})),$
is positive since it is $\widehat{\psi^{(n)}}$. Finally, using an approximate unit of $A$, one obtains $\|\hat{\psi}\|=\|\psi\|$.
Corollary 4.6.
If $\phi\colon\mathcal{B}\to\mathbb{B}(X)$ is cp with compact support, then its integrated form $\tilde{\phi}\colon C^{*}(\mathcal{B})\to\mathbb{B}(X)$ factors through a cp map $\tilde{\phi}_{\mathop{\rm r}}\colon C^{*}_{\mathop{\rm r}}(\mathcal{B})\to\mathbb{B}(X)$ via the reduced representation $\lambda^{\mathcal{B}}\colon C^{*}(\mathcal{B})\to C^{*}_{\mathop{\rm r}}(\mathcal{B})$ of [15]. Moreover, $\|\tilde{\phi}_{\mathop{\rm r}}\|=\|\phi|_{B_{e}}\|$.
Proof.
Let $T\colon\mathcal{B}\to\mathbb{B}(Y)$ be a dilation of $\phi$ with $V\in\mathbb{B}(X,Y)$. Choose a compact neighborhood $N$ of $e$, $f\in C_{c}(G)$ with $f(t^{-1}s)=1$ for $(t,s)\in{\mathop{\rm supp}}(\phi)\times N$, and $g\in C_{c}(G)^{+}$ supported in $N$ with $\int_{G}g=1$. Define $P,Q\colon X\to L^{2}(G)\otimes Y$ by $Px(t)=f(t)Vx$, $Qx(t)=g(t)Vx$.
Recall from [15] that the integrated form of $S\colon\mathcal{B}\to\mathbb{B}(L^{2}(G)\otimes Y)$, $b\mapsto\lambda_{t}\otimes T_{b}$, factors through a *-representation $\hat{S}\colon C^{*}_{\mathop{\rm r}}(\mathcal{B})\to\mathbb{B}(L^{2}(G)\otimes Y)$ with $\hat{S}\circ\lambda^{\mathcal{B}}=\tilde{S}$. For $h\in C_{c}(\mathcal{B})$, $x,y\in X$, and $t\in G$,
$\int_{G}g(s)f(t^{-1}s)\,ds\,\langle x,\phi(h(t))y\rangle=\langle x,\phi(h(t))y\rangle,$
because $\langle x,\phi(h(t))y\rangle\neq 0$ implies $t\in{\mathop{\rm supp}}(\phi)$. Hence
$\langle x,P^{*}\hat{S}_{\lambda^{\mathcal{B}}(h)}Qy\rangle=\int_{G}\!\int_{G}g(s)f(t^{-1}s)\langle x,\phi(h(t))y\rangle\,dsdt=\langle x,\tilde{\phi}_{h}y\rangle,$
[…]
4.1. Completely positive maps between Fell bundles
Let $\mathcal{A}=\{A_{t}\}_{t\in G}$ and $\mathcal{B}=\{B_{t}\}_{t\in G}$ be Fell bundles over the same group.
Definition 4.7.
We say that $\phi\colon\mathcal{A}\to\mathcal{B}$ is completely positive (cp) if the following conditions hold.
  1. (1)
    $\phi$ is continuous.
  2. (2)
    For all $t\in G,$ $\phi(A_{t})\subseteq B_{t}$ and $\phi|_{A_{t}}$ is linear.
  3. (3)
    For all $n\in\mathbb{N}$ and $t\in G^{n}$, $\phi^{t}\colon\mathbb{M}_{t}(\mathcal{A})\to\mathbb{M}_{t}(\mathcal{B})$ given by $\phi^{t}((a_{i,j})_{i,j=1}^{n})=(\phi(a_{i,j}))_{i,j=1}^{n}$, is positive.
The support of $\phi$, ${\mathop{\rm supp}}(\phi),$ is the closure of $\{t\in G\colon\phi(A_{t})\neq\{0\}\}$. If $\|\phi|_{B_{e}}\|\leq 1$, we say that $\phi$ is ccp (contractive and completely positive).
Remark 4.8.
For a continuous function $\phi\colon\mathcal{A}\to\mathcal{B}$ such that for all $t\in G$, $\phi(A_{t})\subseteq B_{t}$ and $\phi|_{A_{t}}$ is linear, the following are equivalent:
  1. (1)
    $\phi$ is cp.
  2. (2)
    For every *-representation $T\colon\mathcal{B}\to\mathbb{B}(X),$ $T\circ\phi$ is cp.
  3. (3)
    There exists a *-representation $T\colon\mathcal{B}\to\mathbb{B}(X)$ such that $T|_{B_{e}}$ is faithful and $T\circ\phi$ is cp.
Remark 4.9.
If $\phi\colon\mathcal{A}\to\mathcal{B}$ is cp, then $\phi|_{A_{e}}$ is cp and, consequently, it is bounded. Combining the previous Remark with the dilation Theorem one gets that for all $a\in\mathcal{A},$ $\|\phi(a)\|\leq\|\phi|_{A_{e}}\|\|a\|$.
Theorem 4.10.
If $\phi\colon\mathcal{A}\to\mathcal{B}$ is cp, then there exists a cp map $\tilde{\phi}\colon C^{*}(\mathcal{A})\to C^{*}(\mathcal{B})$ such that $\tilde{\phi}_{f}=\phi\circ f$, for all $f\in C_{c}(\mathcal{A})$. Besides, $\|\tilde{\phi}\|\leq\|\phi|_{A_{e}}\|$. In case $\phi$ has compact support, there exists a (“wrong way”) cp map $\hat{\phi}\colon C^{*}_{\mathop{\rm r}}(\mathcal{A})\to C^{*}(\mathcal{B})$ such that $\hat{\phi}\circ\lambda^{\mathcal{A}}=\tilde{\phi}$. Moreover, $\|\hat{\phi}\|\leq\|\phi|_{A_{e}}\|$.
Proof.
Let $T\colon\mathcal{B}\to\mathbb{B}(X)$ be a *-representation with faithful integrated form $\tilde{T}\colon C^{*}(\mathcal{B})\to\mathbb{B}(X).$ Then $T\circ\phi$ is cp and it has an integrated form $\widetilde{T\circ\phi}$. For all $f\in C_{c}(\mathcal{A})$, $\|\phi\circ f\|_{C^{*}(\mathcal{B})}=\|\tilde{T}_{\phi\circ f}\|=\|\widetilde{T\circ\phi}_{f}\|\leq\|T\circ\phi|_{B_{e}}\|\|f\|_{C^{*}(\mathcal{A})}$. Hence, $C_{c}(\mathcal{A})\to C^{*}(\mathcal{B}),$ $f\mapsto\phi\circ f,$ is continuous with respect to the $\mathrm{C}^{*}$-norms. The unique continuous extension to $\tilde{\phi}\colon C^{*}(\mathcal{A})\to C^{*}(\mathcal{B})$ is cp because $\tilde{T}\circ\tilde{\phi}=\widetilde{T\circ\phi}$ is cp.
In case $\phi$ has compact support, by Corollary 4.6, for all $f\in C^{*}(\mathcal{A})$ we have $\|\tilde{\phi}_{f}\|=\|\tilde{T}\circ\tilde{\phi}_{f}\|=\|\widetilde{T\circ\phi}_{f}\|\leq\|T\circ\phi|_{A_{e}}\|\|\lambda^{\mathcal{A}}_{f}\|\leq\|\phi|_{A_{e}}\|\|\lambda^{\mathcal{A}}_{f}\|.$ The rest of the proof follows from Remark 4.5. ∎
4.1.1. Characterization of completely positive maps between semidirect product bundles
Assume that $\alpha$ and $\beta$ are $\mathrm{C}^{*}$-actions of $G$ on $A$ and $B$, respectively. To specify a function $P\colon\mathcal{A}\to\mathcal{B}$ between the respective semidirect product bundles such that $P(A\delta_{t})\subseteq B\delta_{t}$, for all $t\in G$, is to specify a set of functions $\{P_{t}\colon A\to B\}_{t\in G}$ ($P(a\delta_{t})=P_{t}(a)\delta_{t}$). In Corollary 4.12 we give a general example where $P_{t}=P_{e}$ for all $t\in G,$ but in that case $P$ has compact support if and only if $P=0.$
Lemma 4.11.
With the notation above, $P$ is cp if and only if each one of the $P_{t}^{\prime}$s is linear and for all $n\in\mathbb{N},$ $t_{1},\ldots,t_{n}\in G$ and $a_{1},\ldots,a_{n}\in A,$ $(\beta_{t_{i}}(P_{t_{i}^{-1}t_{j}}(\alpha_{t_{i}}^{-1}(a_{i}^{*}a_{j}))))_{i,j=1}^{n}$ is a positive matrix of $\mathbb{M}_{n}(B).$
Proof.
Fix $n\in\mathbb{N}$ and $t=(t_{1},\ldots,t_{n})\in G^{n}$. It is easy to check that
$\pi_{\alpha}^{t}\colon$ $\mathbb{M}_{n}(A)\to\mathbb{M}_{t}(\mathcal{A})$ $\pi_{\alpha}^{t}((a_{i,j})_{i,j=1}^{n})$ $=(\alpha_{t_{i}}^{-1}(a_{i,j})\delta_{t_{i}^{-1}t_{j}})_{i,j=1}^{n}$
is a *-isomorphism.
Each positive $a\in\mathbb{M}_{n}(A)$ is a sum of $n$ elements of the form $(a_{i}^{*}a_{j})_{i,j=1}^{n}$ and
$(\pi_{\beta}^{t})^{-1}\circ P^{t}\circ\pi_{\alpha}^{t}((a_{i}^{*}a_{j})_{i,j=1}^{n})=(\pi_{\beta}^{t})^{-1}(P_{t_{i}^{-1}t_{j}}(\alpha_{t_{i}}^{-1}(a_{i}^{*}a_{j}))\delta_{t_{i}^{-1}t_{j}})_{i,j=1}^{n}=(\beta_{t_{i}}(P_{t_{i}^{-1}t_{j}}(\alpha_{t_{i}}^{-1}(a_{i}^{*}a_{j}))))_{i,j=1}^{n}.$
Hence, $P^{t}$ is positive if and only if for all $n\in\mathbb{N},$ $t_{1},\ldots t_{n}\in G$ and $a_{1},\ldots,a_{n}\in A$, the last matrix on the right above is positive. ∎
Corollary 4.12.
If $P\colon A\to B$ is an equivariant cp map, then $\phi\colon\mathcal{A}\to\mathcal{B}$ given by $\phi(a\delta_{s})=P(a)\delta_{s}$, is cp. In addition, $\|\phi|_{A\delta_{e}}\|=\|P\|$.
Proof.
With the notation of Lemma 4.11, $P_{t}=P$ for all $t\in G$ and
$(\beta_{t_{i}}(P_{t_{i}^{-1}t_{j}}(\alpha_{t_{i}}^{-1}(a_{i}^{*}a_{j}))))_{i,j=1}^{n}=(P(a_{i}^{*}a_{j}))_{i,j=1}^{n}$
is a positive matrix. For all $a\in A$, $\|\phi(a\delta_{e})\|=\|P(a)\|\leq\|a\|=\|a\delta_{e}\|$. Hence, $\|\phi|_{A\delta_{e}}\|=\|P\|$. ∎
4.2. Ccp maps and amenability
Definition 4.13.
A Fell bundle $\mathcal{B}=\{B_{t}\}_{t\in G}$ has the approximation property of Bédos and Conti (BCAP) if there exists a net of ccp maps with compact support $\{\Phi_{i}\colon\mathcal{B}\to\mathcal{B}\}_{i\in I}$ that converges to $\operatorname{id}_{\mathcal{B}}$ uniformly on compact slices. The BCAP1 differs only in the convergence condition: it requires $L^{1}$-convergence on compact slices.
In the definition above the nets are bounded, so BCAP implies BCAP1. To prove that both are equivalent to the amenability of the bundle we show two results that should be true if the equivalence holds (Propositions 4.14 and 4.16). For Fell bundles over discrete groups, the Proposition below follows from the results of [9].
Proposition 4.14.
The following claims hold for every Fell bundle $\mathcal{B}=\{B_{t}\}_{t\in G}$.
  1. (1)
    If $\mathcal{B}$ $\mathrm{C}^{*}$-amenable, then $\mathcal{B}$ has the BCAP.
  2. (2)
    If $\mathcal{B}$ has the BCAP1, then $\mathcal{B}$ has the wcp.
Proof.
Let $\{\xi_{i}\}_{i\in I}\subseteq C_{c}(G,B_{e})$ be a net witnessing the positive $1$-approximation property of $\mathcal{B}$. By definition, $\{\Phi_{\xi_{i}}\colon\mathcal{B}\to\mathcal{B}\}_{i\in I}$ is a net of continuous maps that converges to $\operatorname{id}_{\mathcal{B}}$ on compact slices. To prove that each $\phi_{\xi_{i}}$ is cp we take a *-representation $T\colon\mathcal{B}\to\mathbb{B}(X)$ with $T|_{B_{e}}$ faithful. Define $V_{i}\colon X\to L^{2}(G)\times X$ by $V_{i}x(t)=T_{\xi_{i}(t)}a$ and $S\colon\mathcal{B}\to\mathbb{B}(L^{2}(G)\otimes Y)$ by $S(b\in B_{t})=\lambda_{t}\otimes T_{b}$. For all $x,y\in X,$ $t\in G$ and $b\in B_{t}$ we have
$\langle x,V_{i}^{*}S_{b}V_{i}y\rangle$ $=\langle V_{i}x,S_{b}V_{i}y\rangle=\int_{G}\langle T_{\xi_{i}(s)}x,T_{b}T_{\xi_{i}({t^{-1}}s)}y\rangle ds=\int_{G}\langle x,T_{\xi_{i}(s)^{*}b\xi_{i}({t^{-1}}s)}y\rangle ds$
$=\langle x,T\circ\Phi_{\xi_{i}}(b)y\rangle.$
Then $V_{i}^{*}S_{b}V_{i}=T\circ\Phi_{\xi_{i}}(b)$ and Remark 4.8 implies that $\Phi_{\xi_{i}}$ is cp. Notice that $\|V_{i}\|\leq 1$, so $\|\Phi_{\xi_{i}}|_{B_{e}}\|=\|T\circ\Phi_{\xi_{i}}|_{B_{e}}\|=\|V_{i}^{*}S|_{B_{e}}V_{i}\|\leq 1$.
Say $\Phi_{\xi_{i}}(b)\neq 0$ with $b\in B_{t}$. By the definition of $\Phi_{\xi_{i}}(b)$, there exists $s\in{\mathop{\rm supp}}(\xi_{i})$ such that ${t^{-1}}s\in{\mathop{\rm supp}}(\xi_{i})$. This yields ${\mathop{\rm supp}}(\Phi_{\xi_{i}})\subseteq{\mathop{\rm supp}}(\xi_{i}){\mathop{\rm supp}}(\xi_{i})^{-1}$; which completes the proof of the first claim.
Assume that $\mathcal{B}$ has the BCAP1 and take a net $\{\Phi_{i}\}_{i\in I}$ witnessing such property. By Theorem 4.10, for all $f\in C_{c}(\mathcal{B})$ we have $\|\Phi_{i}\circ f\|_{C^{*}(\mathcal{B})}\leq\|f\|_{C^{*}_{\mathop{\rm r}}(\mathcal{B})}$. In addition, $\|\Phi_{i}\circ f-f\|_{C^{*}(\mathcal{B})}\leq\|\Phi_{i}\circ f-f\|_{1}$. Hence, $\|f\|_{C^{*}(\mathcal{B})}=\lim_{i}\|\Phi_{i}\circ f\|_{C^{*}(\mathcal{B})}\leq\|f\|_{C^{*}_{\mathop{\rm r}}(\mathcal{B})}\leq\|f\|_{C^{*}(\mathcal{B})}$ and this implies that $\lambda^{\mathcal{B}}\colon C^{*}(\mathcal{B})\to C^{*}_{\mathop{\rm r}}(\mathcal{B})$ is an isometry. ∎
Lemma 4.15.
If $\mathcal{A}=\{A_{t}\}_{t\in G}$, $\mathcal{B}=\{B_{t}\}_{t\in G}$ and $\mathcal{C}=\{C_{t}\}_{t\in H}$ are Fell bundles and $\Phi\colon\mathcal{A}\to\mathcal{B}$ is a cp map, then there exists a cp map $\Phi\bigotimes_{\max}\operatorname{id}_{\mathcal{C}}\colon\mathcal{A}\bigotimes_{\max}\mathcal{C}\to\mathcal{B}\bigotimes_{\max}\mathcal{C}$ such that $\Phi\bigotimes_{\max}\operatorname{id}_{\mathcal{C}}(a\otimes c)=\Phi(a)\otimes c$. In addition, $\|\Phi\bigotimes_{\max}\operatorname{id}_{\mathcal{C}}|_{A_{e}\otimes_{\max}B_{e}}\|=\|\Phi|_{A_{e}}\|$.
Proof.
The integrated form $\tilde{\Phi}\colon C^{*}(\mathcal{A})\to C^{*}(\mathcal{B})$ induces a cp map
$\tilde{\Phi}\otimes_{\max}\operatorname{id}\colon C^{*}(\mathcal{A})\otimes_{\max}C^{*}(\mathcal{C})\to C^{*}(\mathcal{B})\otimes_{\max}C^{*}(\mathcal{C})$
such that for all $f\in C_{c}(\mathcal{A})$ and $g\in C_{c}(\mathcal{C})$, $(\tilde{\Phi}\otimes_{\max}\operatorname{id})(f\otimes g)=(\Phi\circ f)\otimes g$. By [1, Proposition 4.2], there exist $\mathrm{C}^{*}$-isomorphisms
$\mu$ $\colon C^{*}(\mathcal{A})\otimes_{\max}C^{*}(\mathcal{C})\to C^{*}(\mathcal{A}\bigotimes_{\max}\mathcal{C})$ $\nu$ $\colon C^{*}(\mathcal{B})\otimes_{\max}C^{*}(\mathcal{C})\to C^{*}(\mathcal{B}\bigotimes_{\max}\mathcal{C})$
such that for all $f\in C_{c}(\mathcal{A})$, $g\in C_{c}(\mathcal{B})$ and $h\in C_{c}(\mathcal{C})$, $\mu(f\otimes h)=f\oslash h$ and $\nu(g\otimes h)=g\oslash h$. The cp map
$\Psi:=\nu\circ(\tilde{\Phi}\otimes_{\max}\operatorname{id})\circ\mu^{-1}\colon C^{*}(\mathcal{A}\bigotimes_{\max}\mathcal{C})\to C^{*}(\mathcal{B}\bigotimes_{\max}\mathcal{C})$
is the unique continuous linear function such that for all $f\in C_{c}(\mathcal{A})$ and $g\in C_{c}(\mathcal{C})$, $\Psi(f\oslash g)=(\Phi\circ f)\oslash g$.
Take a nondegenerate *-representation $T\colon\mathcal{B}\bigotimes_{\max}\mathcal{C}\to\mathbb{B}(X)$ whose restriction to any fiber is isometric. If $\tilde{T}\colon C^{*}(\mathcal{B}\bigotimes_{\max}\mathcal{C})\to\mathbb{B}(X)$ is the integrated form of $T$, then $\tilde{T}\circ\Psi$ is cp and it has a Stinespring dilation $\tilde{S}\colon C^{*}(\mathcal{A}\bigotimes_{\max}\mathcal{C})\to\mathbb{B}(Y)$ with associated map $V\in\mathbb{B}(X,Y)$. Naturally, we name $S\colon\mathcal{A}\bigotimes_{\max}\mathcal{C}\to\mathbb{B}(Y)$ the disintegrated form of $\tilde{S}$. Using [16, VIII 11.8] and the fact that $\Psi(f\oslash g)=(\Phi\circ f)\oslash g$ get get $V^{*}S_{f(s)\otimes g(t)}V=T_{\Phi(f(s))\otimes g(t)}$, for all $f\in C_{c}(\mathcal{A})$, $g\in C_{c}(\mathcal{C})$ and $(s,t)\in G\times H$. This implies $V^{*}S(A_{s}\otimes_{\max}C_{t})V\subseteq T(B_{s}\otimes_{\max}C_{t})$, then there exists a function $\Phi\bigotimes_{\max}\operatorname{id}_{\mathcal{C}}\colon\mathcal{A}\bigotimes_{\max}\mathcal{C}\to\mathcal{B}\bigotimes_{\max}\mathcal{C}$ such that $(\Phi\bigotimes_{\max}\operatorname{id}_{\mathcal{C}})(A_{s}\otimes_{\max}C_{t})\subseteq B_{s}\otimes_{\max}C_{t}$ and $T\circ\Phi(z)=V^{*}S_{z}V,$ for all $(s,t)\in G\times H$ and $z\in\mathcal{A}\bigotimes_{\max}\mathcal{C}$. This implies $(\Phi\bigotimes_{\max}\operatorname{id}_{\mathcal{C}})(a\otimes c)=\Phi(a)\otimes c$ and $\|\Phi(z)\|\leq\|V\|^{2}\|z\|=\|\tilde{\Phi}\|\|z\|=\|\Phi|_{A_{e}}\|\|z\|$. We have $(\Phi\bigotimes_{\max}\operatorname{id}_{\mathcal{C}})(f\oslash g)=(\Phi\circ f)\oslash g$, so [16, II 13.16] implies $\Phi\bigotimes_{\max}\operatorname{id}_{\mathcal{C}}$ is continuous and it is cp because $T\circ\Phi(z)=V^{*}S_{z}V$ (see Remark 4.8). Finally, $\|(\Phi\bigotimes_{\max}\operatorname{id}_{\mathcal{C}})|_{A_{e}\otimes_{\max}C_{e}}\|=\|\Phi|_{A_{e}}\otimes_{\max}\operatorname{id}_{C_{e}}\|=\|\Phi|_{A_{e}}\|$. ∎
Proposition 4.16.
If $\mathcal{A}=\{A_{t}\}_{t\in G}$ and $\mathcal{B}=\{B_{t}\}_{t\in G}$ are Fell bundles and at least one of them has the BCAP1, then $\mathcal{A}\otimes_{\max}^{d}\mathcal{B}$ has the BCAP1.
Proof.
Let $\{\Phi_{i}\}_{i\in I}$ be a net of maps witnessing the BCAP1 of $\mathcal{A}$ and define
$\Phi_{i}\otimes_{\max}^{d}\operatorname{id}_{\mathcal{B}}\colon\mathcal{A}\otimes_{\max}^{d}\mathcal{B}\to\mathcal{A}\otimes_{\max}^{d}\mathcal{B}$
as the restriction of $\Phi_{i}\bigotimes_{\max}\operatorname{id}_{\mathcal{B}}$ to the reduction $\mathcal{A}\otimes_{\max}^{d}\mathcal{B}=(\mathcal{A}\bigotimes_{\max}\mathcal{B})_{G}$. Then $\{\Phi_{i}\otimes_{\max}^{d}\operatorname{id}_{\mathcal{B}}\}_{i\in I}$ is a net of cp maps and $\forall\,i\in I$, $\|(\Phi_{i}\otimes_{\max}^{d}\operatorname{id}_{\mathcal{B}})|_{A_{e}\otimes_{\max}B_{e}}\|=\|(\Phi_{i}\bigotimes_{\max}\operatorname{id}_{\mathcal{B}})|_{A_{e}\otimes_{\max}B_{e}}\|=\|\Phi_{i}\|\leq 1$.
For all $f\in C_{c}(\mathcal{A})$ and $g\in C_{c}(\mathcal{B})$, $(\Phi_{i}\otimes_{\max}^{d}\operatorname{id}_{\mathcal{B}})\circ(f\oslash^{d}g)=(\Phi_{i}\circ f)\otimes_{\max}^{d}g$. Hence,
$\|(\Phi_{i}\otimes_{\max}^{d}\operatorname{id}_{\mathcal{B}})\circ(f\oslash^{d}g)-f\oslash^{d}g\|_{1}\leq\|\Phi_{i}\circ f-f\|_{1}\|g\|_{\infty}$
and taking limit in $i$ it follows that $\lim_{i}\|(\Phi_{i}\otimes_{\max}^{d}\operatorname{id}_{\mathcal{B}})\circ(f\oslash^{d}g)-f\oslash^{d}g\|_{1}=0$. Taking $\Gamma:={\mathop{\rm span}}\{f\oslash^{d}g\colon f\in C_{c}(\mathcal{A}),\ g\in C_{c}(\mathcal{B})\}$ and using Lemma 3.4 we get that $\mathcal{A}\otimes_{\max}^{d}\mathcal{B}$ has the BCAP1. ∎
Corollary 4.17.
For Fell bundles, $\mathrm{C}^{*}$-amenability is equivalent to the BCAP and to the BCAP1.
Proof.
By Proposition 4.14, $\mathrm{C}^{*}$-amenability implies the BCAP and right after Definition 4.13 we noticed that BCAP implies BCAP1. Assume that the Fell bundle $\mathcal{B}$ over $G$ has the BCAP1. Combining Propositions 4.14 and 4.16 we get that for any other Fell bundle $\mathcal{A}$ over $G$, $\mathcal{A}\otimes_{\max}^{d}\mathcal{B}$ has the wcp, Theorem 3.2 implies $\mathcal{B}$ is $\mathrm{C}^{*}$-amenable. ∎
Corollary 4.18.
If the Fell bundle $\mathcal{B}=\{B_{t}\}_{t\in G}$ is $\mathrm{C}^{*}$-amenable and $H\subseteq G$ is a closed subgroup, then the reduction $\mathcal{B}_{H}=\{B_{t}\}_{t\in H}$ is $\mathrm{C}^{*}$-amenable.
Proof.
Let $\{\Phi_{i}\colon\mathcal{B}\to\mathcal{B}\}_{i\in I}$ be a net of ccp maps witnessing the BCAP of $\mathcal{B}$. Notice that $\Phi_{i}(\mathcal{B}_{H})\subseteq\mathcal{B}_{H}$ and that for all $n\in\mathbb{N}$ and $t\in H^{n},$ $\mathbb{M}_{t}(\mathcal{B})=\mathbb{M}_{t}(\mathcal{B}_{H})$. By [16, II 14.8], $\{f|_{H}\colon f\in C_{c}(\mathcal{B})\}=C_{c}(\mathcal{B}_{H})$. Then $\{\mathcal{B}_{H}\to\mathcal{B}_{H},b\mapsto\Phi_{i}(b)\}_{i\in I}$ is a net witnessing the BCAP of $\mathcal{B}_{H}$. ∎
Corollary 4.19.
If $\alpha=(\{A_{t}\}_{t\in G},\{\alpha_{t}\}_{t\in G})$ is a $\mathrm{C}^{*}$-amenable $\mathrm{C}^{*}$-partial action and $H\subseteq G$ is a closed subgroup, then $\alpha|_{H}:=(\{A_{t}\}_{t\in H},\{\alpha_{t}\}_{t\in H})$ is a $\mathrm{C}^{*}$-amenable $\mathrm{C}^{*}$-partial action.
Proof.
If $\mathcal{B}_{\alpha}$ denotes the semidirect product bundle of $\alpha$, then $(\mathcal{B}_{\alpha})_{H}=\mathcal{B}_{\alpha|_{H}}$ and the rest follows from the previous Corollary. ∎
Conditional expectations for Fell bundles over discrete groups were defined by Exel [14]. Recently, in [4], F. Abaide extended this notion to to Fell bundles over locally compact groups and proved that all conditional expectations are completely positive and contractive. The following result implies that if $E\colon\mathcal{B}\to\mathcal{A}$ is a conditional expectation and $\mathcal{B}$ has the approximation property, then $\mathcal{A}$ has the approximation property.
Corollary 4.20.
If $\mathcal{A}=\{A_{t}\}_{t\in G}$ is a Fell subbundle of $\mathcal{B}=\{B_{t}\}_{t\in G}$, $\mathcal{B}$ is $\mathrm{C}^{*}$-amenable and there exists net of ccp maps $\{\Phi_{i}\colon\mathcal{B}\to\mathcal{A}\}_{i\in I}$ such that $\{\Phi_{i}|_{\mathcal{A}}\}_{i\in I}$ converges to $\operatorname{id}_{\mathcal{A}}$ on compact slices, then $\mathcal{A}$ is $\mathrm{C}^{*}$-amenable.
Proof.
It suffices to show that given $F\subseteq C_{c}(\mathcal{A})$ finite and $\varepsilon>0$, there exists a ccp map $\Psi\colon\mathcal{A}\to\mathcal{A}$ with compact support such that for all $f\in F$, $\|\Phi\circ f-f\|_{\infty}<\varepsilon$. Fix $F$ and $\varepsilon$. There exists a ccp map with compact support $\Psi\colon\mathcal{B}\to\mathcal{B}$ such that $\forall\ f\in F$, $\|\Psi\circ f-f\|_{\infty}<\varepsilon/2.$ There is also an $i\in I$ such that for all $f\in F$, $\|\Phi_{i}\circ f-f\|_{\infty}<\varepsilon/2$. Notice that $\Psi_{i}:=\Phi_{i}\circ\Psi|_{\mathcal{A}}\colon\mathcal{A}\to\mathcal{A}$ is ccp and for all $f\in F$,
$\|\Psi_{i}\circ f-f\|_{\infty}\leq\|\Phi_{i}\circ(\Psi\circ f-f)\|+\|\Phi_{i}\circ f-f\|_{\infty}\leq\|\Psi\circ f-f\|_{\infty}+\|\Phi_{i}\circ f-f\|_{\infty}<\varepsilon.$
Moreover, ${\mathop{\rm supp}}(\Psi_{i})\subseteq{\mathop{\rm supp}}(\Psi)$ is compact. ∎
Corollary 4.21.
If $\alpha$ is a $\mathrm{C}^{*}$-amenable action of $G$ on $A$, $B\subseteq A$ is an invariant $\mathrm{C}^{*}$-subalgebra and any of the conditions below holds, then $\alpha|_{B}$ is $\mathrm{C}^{*}$-amenable.
  1. (1)
    There exists an equivariant conditional expectation $E\colon A\to B$ or, more generally, there exists a net of ccp equivariant maps $\{E_{i}\colon A\to B\}_{i\in I}$ such that $\{E_{i}|_{B}\}_{i\in I}$ converges pointwise in norm to $\operatorname{id}_{B}$
  2. (2)
    $B$ is hereditary in $A$.
Proof.
Let $\mathcal{A}$ and $\mathcal{B}$ be the semidirect product bundles of $\alpha$ and $\beta$, respectively. Define for each $i\in I$, $\Phi_{i}\colon\mathcal{A}\to\mathcal{B}$ by $\Phi_{i}(a\delta_{t})=E_{i}(a)\delta_{t}$. Corollary 4.12 implies that $\Phi_{i}$ is ccp. Let $\Gamma\subseteq C_{c}(\mathcal{B})$ be the set of sections of the form $t\mapsto g(t)b\delta_{t}$ with $g\in C_{c}(G)$ and $b\in B$. For all $f\in\Gamma$, $\lim_{i}\|\Phi_{i}\circ f-f\|_{\infty}=0$. Using Lemma 3.4 we get that $\{\Phi_{i}|_{\mathcal{B}}\}_{i\in I}$ converges uniformly to $\operatorname{id}_{\mathcal{B}}$ on compact slices. Corollary 4.20 implies $\mathcal{B}$ is amenable. The claim for the case where $B$ is hereditary in $A$ follows from Lemma 3.8. ∎
5. Reductions and quotient bundles
In this section we use a fixed Fell bundle $\mathcal{B}=\{B_{t}\}_{t\in G}$ over a group $G$ and a closed subgroup $H\subseteq G$; any additional hypothesis will be stated explicitly. As we saw in Corollary 4.18, the reduction $\mathcal{B}_{H}:=\{B_{t}\}_{t\in H}$ is $\mathrm{C}^{*}$-amenable if $\mathcal{B}$ is. Here we study when the wcp, or the nuclearity of the cross-sectional $\mathrm{C}^{*}$-algebra, passes from $\mathcal{B}$ to $\mathcal{B}_{H}$. In addition, for $H$ normal in $G$, we also study similar problems for the $\mathrm{C}^{*}$-completion of the partial cross-sectional bundle over $G/H$ derived from $\mathcal{B}$ (defined in [16]).
We will make extensive use of integration (and disintegration) of *-representations of Fell bundles, the regional topology, weak containment ($\preceq$), weak equivalence of *-representations ($\sim$) and other constructions of [16], like the induction of *-representations. Given a *-representation $T$ of $\mathcal{B}$, we put a $\tilde{\ }$ to indicate the integrated form of $T,$ e.g. $\tilde{T}$. For example, $\lambda\tilde{\otimes}T$ is the integrated form of the tensor product representation $\lambda\otimes T$. Similarly, if $S$ is a *-representation of $\mathcal{B}_{H}$, $\operatorname{Ind}_{B_{H}\uparrow\mathcal{B}}(S)$ is the *-representation of $\mathcal{B}$ induced by $S$ and $\widetilde{\operatorname{Ind}}_{B_{H}\uparrow\mathcal{B}}(S)$ its integrated form.
5.1. Reductions to open subgroups and full cross-sectional $\mathrm{C}^{*}$-algebras
Assume $H$ is open in $G$ and think $C_{c}(\mathcal{B}_{H})\subseteq C_{c}(\mathcal{B})$ by extending each $f\in C_{c}(\mathcal{B}_{H})$ as 0 outside $H.$ This gives $L^{1}(\mathcal{B}_{H})\subseteq L^{1}(\mathcal{B})$ and, passing to enveloping $\mathrm{C}^{*}$-algebras, we get a *-homomorphism $\iota\colon C^{*}(\mathcal{B}_{H})\to C^{*}(\mathcal{B})$. To prove $\iota$ is injective we take a nondegenerate *-representation $T\colon C^{*}(\mathcal{B}_{H})\to\mathbb{B}(X)$ with faithful integrated form $\tilde{T}\colon C^{*}(\mathcal{B}_{H})\to\mathbb{B}(X).$ By [17, Theorem 2.1], $T$ is $\mathcal{B}$-positive and it can be induced to a *-representation $S\colon\mathcal{B}\to\mathbb{B}(Y)$ by the process of [16, Ch. XI]. Since $G/H$ is discrete, we can use [16, XI 14.21] to deduce that $T$ is a subrepresentation of $S|_{\mathcal{B}_{H}}$. Consequently, $\tilde{T}|_{C_{c}(\mathcal{B}_{H})}$ is a subrepresentation of $\tilde{S}\circ\iota|_{C_{c}(\mathcal{B}_{H})}$ and this implies that $\iota$ is an isometry because, for all $f\in C_{c}(\mathcal{B}_{H}),$ $\|\iota(f)\|\leq\|f\|=\|\tilde{T}_{f}\|\leq\|\tilde{S}_{\iota(f)}\|\leq\|\iota(f)\|.$ From now on we think $C^{*}(\mathcal{B}_{H})\subseteq C^{*}(\mathcal{B})$ and omit any mention of $\iota$.
Notice that for all $f,g\in C_{c}(\mathcal{B}_{H})$ and $h\in C_{c}(\mathcal{B})$ we have $f*h*g=f*h|_{H}*g.$ Since $C_{c}(\mathcal{B}_{H})$ contains an approximate unit of $C^{*}(\mathcal{B}_{H}),$ the restriction map $E_{c}\colon C_{c}(\mathcal{B})\to C_{c}(\mathcal{B}_{H}),$ $E_{c}(h):=h|_{H},$ is a contraction with respect to the $\mathrm{C}^{*}$-norm. By Tomiyama’s Theorem [10, Theorem 1.5.10], the continuous extension $E\colon C^{*}(\mathcal{B})\to C^{*}(\mathcal{B}_{H})$ of $E_{c}$ is a conditional expectation.
Remark 5.1.
If $C^{*}(\mathcal{B})$ is nuclear, then so is $C^{*}(\mathcal{B}_{H})$.
5.2. Reductions to open subgroups and reduced cross-sectional $\mathrm{C}^{*}$-algebras
We continue assuming that $H$ is open in $G$ and denote $\lambda^{H}$ and $\lambda^{G}$ the respective left regular representations. To prove that $C_{c}(\mathcal{B}_{H})\subseteq C_{c}(\mathcal{B})$ induces and inclusion of $\mathrm{C}^{*}$-algebras $C^{*}_{\mathop{\rm r}}(\mathcal{B}_{H})\subseteq C^{*}_{\mathop{\rm r}}(\mathcal{B})$ we take a nondegenerate *-representation $T\colon\mathcal{B}\to\mathbb{B}(X)$ with $T|_{B_{e}}$ faithful. By [15, Corollary 2.15], there exists a faithful *-representation $\pi\colon C^{*}_{\mathop{\rm r}}(\mathcal{B})\to\mathbb{B}(L^{2}(G,X))$ such that, for all $f\in C_{c}(\mathcal{B}),$ $\pi(f)=[\lambda^{G}\tilde{\otimes}T]_{f}.$ For all $f\in C_{c}(\mathcal{B}_{H})\subseteq C_{c}(\mathcal{B}),$ $\pi(f)=[\lambda^{G}|_{H}\tilde{\otimes}T|_{\mathcal{B}_{H}}]_{f}$ and $\lambda^{G}|_{H}$ is a direct sum of copies of $\lambda^{H}$ because $G$ is the disjoint union of open cosets $Ht$. Then $\|f\|_{C^{*}_{\mathop{\rm r}}(\mathcal{B})}=\|[\lambda^{G}\tilde{\otimes}T]_{f}\|=\|[\lambda^{H}\tilde{\otimes}T|_{\mathcal{B}_{H}}]_{f}\|=\|f\|_{C^{*}_{\mathop{\rm r}}(\mathcal{B}_{H})},$ which implies that we may identify $C^{*}_{\mathop{\rm r}}(\mathcal{B}_{H})$ with the closure of $C_{c}(\mathcal{B}_{H})$ in $C^{*}_{\mathop{\rm r}}(\mathcal{B})$.
Let $P\in\mathbb{B}(L^{2}(G,X))$ be the orthogonal projection onto $L^{2}(H,X).$ For all $f\in C_{c}(\mathcal{B})$ we have $P(\lambda^{G}\tilde{\otimes}T)_{f}P=P(\lambda^{G}\tilde{\otimes}T)_{f|_{H}}P$ and
$\|f|_{H}\|_{C^{*}_{\mathop{\rm r}}(\mathcal{B}_{H})}=\|(\lambda^{H}\tilde{\otimes}T|_{\mathcal{B}_{H}})_{f|_{H}}\|=\|P(\lambda^{G}\tilde{\otimes}T)_{f|_{H}}P\|\leq\|(\lambda^{G}\tilde{\otimes}T)_{f}\|=\|f\|_{C^{*}_{\mathop{\rm r}}(\mathcal{B})}.$
Then the continuous extension $E_{\mathop{\rm r}}\colon C^{*}_{\mathop{\rm r}}(\mathcal{B})\to C^{*}_{\mathop{\rm r}}(\mathcal{B}_{H})$ of $E_{c}$ is a conditional expectation.
Remark 5.2.
If $C^{*}_{\mathop{\rm r}}(\mathcal{B})$ is nuclear, then so it is $C^{*}_{\mathop{\rm r}}(\mathcal{B}_{H})$.
Remark 5.3.
If $\mathcal{B}$ has the wcp, then $\mathcal{B}_{H}$ has the wcp because the closure of $C_{c}(\mathcal{B}_{H})$ in $C^{*}(\mathcal{B})$ is canonically isomorphic to the closure of $C_{c}(\mathcal{B}_{H})$ in $C^{*}_{\mathop{\rm r}}(\mathcal{B}).$
5.3. The universal completion of the partial cross-sectional bundle
Assume that $H$ is a closed and normal subgroup of $G$. Let $\mathcal{L}=\{L^{1}(\mathcal{B}_{u})\}_{u\in G/H}$ be the partial cross-sectional bundle over $G/H$ derived from $\mathcal{B}$, and ${\mathcal{B}}_{/H}=\{\dot{B}_{u}\}_{u\in G/H}$ the bundle $\mathrm{C}^{*}$-completion of $\mathcal{L}$, see [16, VIII 6 & VIII 16.7]. By construction, each fiber $L^{1}(\mathcal{B}_{u})$ is the $\|\ \|_{1}$-completion of $C_{c}(\mathcal{B}_{u})$, where $\mathcal{B}_{u}=\{B_{t}\}_{t\in u}$ is the reduction of $\mathcal{B}$ to the coset $u\in G/H$. The natural map $C_{c}(\mathcal{B})\to C_{c}(\mathcal{L}),$ $f\mapsto\tilde{f}$ with $\tilde{f}(u)=f|_{u}$, extends to a *-isomorphim $\Phi\colon L^{1}(\mathcal{B})\to L^{1}(\mathcal{L})$ (see [16, VIII 6.7]). The cross-sectional $\mathrm{C}^{*}$-algebras of $\mathcal{B}$ and $\mathcal{L}$ are the envelping $\mathrm{C}^{*}$-algebras of $L^{1}(\mathcal{B})$ and $L^{1}(\mathcal{L})$, so we may think $C^{*}(\mathcal{B})=C^{*}(\mathcal{L})$. In addition, the map $\rho\colon\mathcal{L}\to{\mathcal{B}}_{/H}$ of [16, VIII 17.7] is a dense embedding in the sense of [16, XI 12.6], so $\{\rho\circ f\colon f\in C_{c}(\mathcal{L})\}$ is dense in $C_{c}({\mathcal{B}}_{/H})$ with respect to the inductive limit topology (use [16, II 14.6]). The construction of the bundle $\mathrm{C}^{*}$-completion implies that all *-representations of $\mathcal{L}$ have the form $T\circ\rho$ for a *-representation $T$ of ${\mathcal{B}}_{/H}.$ Then the continuous *-homomorphism $L^{1}(\mathcal{L})\to L^{1}({\mathcal{B}}_{/H})$ mapping $f\in C_{c}(\mathcal{L})$ to $\rho\circ f$ induces an isomorphism between $C^{*}(\mathcal{L})$ and $C^{*}({\mathcal{B}}_{/H})$. This yields $C^{*}(\mathcal{B})=C^{*}({\mathcal{B}}_{/H})$ or, more precisely, there exists an isomorphism
(5.4) $\Omega\colon C^{*}(\mathcal{B})\to C^{*}({\mathcal{B}}_{/H})\textrm{ such that }\forall f\in C_{c}(\mathcal{B}),\ \Omega(f)=\rho\circ\tilde{f}$.
We claim that the fiber $\dot{B}_{H}$ of ${\mathcal{B}}_{/H}$ over the unit $H$ of $G/H$ is $C^{*}(\mathcal{B}_{H})$. For the bundle $\mathcal{L}$, [16, XI 11.6] says that $C^{*}(\mathcal{B}_{H})=\dot{B}_{H}$ if and only if every *-representation of $L^{1}(\mathcal{B}_{H})$ is $\mathcal{L}$-positive. Using [16, XI 12.7] we get that this last claim is equivalent to say that every *-representation of $\mathcal{B}_{H}$ is $\mathcal{B}$-positive, which is the case [17, Theorem 2.1].
5.4. The reduced completion of the partial cross-sectional bundle
We want to construct a reduced version ${\mathcal{B}}^{\mathop{\rm r}}_{/H}$ of ${\mathcal{B}}_{/H}$ so that $C^{*}_{\mathop{\rm r}}(\mathcal{B})$ is isomophic to $C^{*}_{\mathop{\rm r}}({\mathcal{B}}^{\mathop{\rm r}}_{/H})$ ($H$ is normal and closed in $G$). To do this we take a nondegenerate *-representation $T\colon\mathcal{B}\to\mathbb{B}(X)$ with $\phi:=T|_{B_{e}}$ faithful, and let $S\colon\mathcal{L}\to\mathbb{B}(Y)$ be the *-representation corresponding to $\operatorname{Ind}_{B_{e}\uparrow\mathcal{B}}(\phi)$ via the process of [16, VIII 15.9]. Using $S$ and the construction of [16, VIII 16.7] we can produce a Fell bundle ${\mathcal{B}}^{\mathop{\rm r}}_{/H}=\{\dot{B}^{\mathop{\rm r}}_{u}\}_{u\in G/H}$ and a dense embedding $\rho^{\mathop{\rm r}}\colon\mathcal{L}\to{\mathcal{B}}^{\mathop{\rm r}}_{/H}$ such that, for all $f\in\mathcal{L},$ $\|\rho^{\mathop{\rm r}}(f)\|=\|S_{f}\|$. By construction, for all $r\in G$; $f\in C_{c}(\mathcal{B}_{rH})$ and $\xi\in Y,$ $S_{f}\xi=\int_{N}\operatorname{Ind}_{B_{e}\uparrow\mathcal{B}}(\phi)_{f(rt)}\xi dt$. Notice that we may replace $f(rt)$ with $\tilde{f}(rN)(rt)$. This ultimately implies that $\tilde{S}\circ\Phi|_{C_{c}(\mathcal{B})}=\widetilde{\operatorname{Ind}}_{B_{e}\uparrow\mathcal{B}}(\phi)|_{C_{c}(\mathcal{B})}$.
The fiber $\dot{B}^{\mathop{\rm r}}_{H}$ of ${\mathcal{B}}^{\mathop{\rm r}}_{/H}$ over the unit $H$ of $G/H$ is the completion of $L^{1}(\mathcal{B}_{H})$ with respect to the seminorm $f\mapsto\|S_{f}\|.$ Hence, to prove $\dot{B}^{\mathop{\rm r}}_{H}=C^{*}_{\mathop{\rm r}}(\mathcal{B}_{H})$ it suffices to show that for all $f\in C_{c}(\mathcal{B}_{H})$ we have $\|f\|_{C^{*}_{\mathop{\rm r}}(\mathcal{B}_{H})}=\|S_{f}\|.$ By Fell’s absorption principle [15, 17], $\operatorname{Ind}_{B_{e}\uparrow\mathcal{B}}(\phi)\equiv\operatorname{Ind}_{B_{e}\uparrow\mathcal{B}}(T|_{B_{e}})$ is weakly equivalent ($\sim$) to $\lambda^{G}\otimes T.$ The process of restricting *-representations is continuous with respect to the regional topology [16, VIII 21.20] and, by definition, it is a homeomorphism with respect to the operation of taking integrated forms. This implies
$S|_{L^{1}(\mathcal{B}_{H})}\sim[(\lambda^{G}\otimes T)|_{\mathcal{B}_{H}}]^{\tilde{\ }}=[(\lambda^{G}|_{H})\otimes(T|_{\mathcal{B}_{H}})]^{\tilde{\ }}\sim\lambda^{H}\tilde{\otimes}(T|_{\mathcal{B}_{H}})$
where the last equivalence holds because $\lambda^{G}|_{H}\sim\lambda^{H}$ and because the process of taking tensor products with representations of the base group is continuous with respect to the regional topology. By [15, Corollary 2.15], for all $f\in L^{1}(\mathcal{B}_{H})$ we have $\|S_{f}\|=\|\lambda^{H}\tilde{\otimes}(T|_{\mathcal{B}_{H}})(f)\|=\|f\|_{C^{*}_{\mathop{\rm r}}(\mathcal{B}_{H})}.$ Then $\dot{B}^{\mathop{\rm r}}_{H}=C^{*}_{\mathop{\rm r}}(\mathcal{B}_{H})$.
Now we want to prove that $C^{*}_{\mathop{\rm r}}(\mathcal{B})=C^{*}_{\mathop{\rm r}}({\mathcal{B}}^{\mathop{\rm r}}_{/H}).$ We know $\lambda^{G}\tilde{\otimes}T\sim\operatorname{Ind}_{B_{e}\uparrow\mathcal{B}}(\phi),$ so there exists a faithful *-representation
$\mu$ $\colon C^{*}_{\mathop{\rm r}}(\mathcal{B})\to\overline{\widetilde{\operatorname{Ind}}_{B_{e}\uparrow\mathcal{B}}(\phi)(C_{c}(\mathcal{B}))}$
[…]
By induction in stages,
$\operatorname{Ind}_{B_{e}\uparrow\mathcal{B}}(\phi)=\operatorname{Ind}_{\mathcal{B}_{H}\uparrow\mathcal{B}}(\operatorname{Ind}_{B_{e}\uparrow\mathcal{B}_{H}}(T|_{B_{e}})).$
The integrated form of $\operatorname{Ind}_{B_{e}\uparrow\mathcal{B}_{H}}(T|_{B_{e}})\sim\lambda^{H}\otimes(T|_{\mathcal{B}_{H}})$ is a *-representation $R$ of $L^{1}(\mathcal{B}_{H})$ for which there exists a faithful *-representation $R^{\prime}$ of $\dot{B}^{\mathop{\rm r}}_{H}=C^{*}_{\mathop{\rm r}}(\mathcal{B}_{H})$ such that $R^{\prime}\circ\rho|_{L^{1}(\mathcal{B}_{H})}=R,$ see [15, Corollary 2.15]. By [16, XI 12.6-12.7], $S=\operatorname{Ind}_{L^{1}(\mathcal{B}_{H})\uparrow\mathcal{L}}(R)$ and $\operatorname{Ind}_{\dot{B}^{\mathop{\rm r}}_{H}\uparrow{\mathcal{B}}^{\mathop{\rm r}}_{/H}}(R^{\prime})\circ\rho^{\mathop{\rm r}}=\operatorname{Ind}_{L^{1}(\mathcal{B}_{H})\uparrow\mathcal{L}}(R).$ From Proposition 3.1 of [17] we get the existence of a faithful *-representation
$\nu\colon C^{*}_{\mathop{\rm r}}({\mathcal{B}}^{\mathop{\rm r}}_{/H})\to\overline{\widetilde{\operatorname{Ind}}_{\dot{B}^{\mathop{\rm r}}_{H}\uparrow{\mathcal{B}}^{\mathop{\rm r}}_{/H}}(R)(C_{c}({\mathcal{B}}^{\mathop{\rm r}}_{/H}))}$
such that $\nu(f)=\widetilde{\operatorname{Ind}}_{\dot{B}^{\mathop{\rm r}}_{H}\uparrow{\mathcal{B}}^{\mathop{\rm r}}_{/H}}(R^{\prime})(f),$ for all $f\in C_{c}({\mathcal{B}}^{\mathop{\rm r}}_{/H}).$
For all $f\in C_{c}(\mathcal{B})$ we have
[…]
The function $\Omega^{\mathop{\rm r}}\colon C_{c}(\mathcal{B})\to C_{c}({\mathcal{B}}^{\mathop{\rm r}}_{/H}),$ $f\mapsto\rho^{\mathop{\rm r}}\circ\tilde{f}$, is a *-homomorphism because $\nu\circ\Omega^{\mathop{\rm r}}=\mu$. Using [16, II 14.6] one can show that the image of $\Omega^{\mathop{\rm r}}$ is dense with respect to the inductive limit topology, so $\Omega^{\mathop{\rm r}}(C_{c}(\mathcal{B}))$ is dense in $C^{*}({\mathcal{B}}^{\mathop{\rm r}}_{/H})$ and $\Omega^{\mathop{\rm r}}$ extends to an isomorphism
(5.5) $\Omega^{\mathop{\rm r}}\colon C^{*}_{\mathop{\rm r}}(\mathcal{B})\to C^{*}_{\mathop{\rm r}}({\mathcal{B}}^{\mathop{\rm r}}_{/H}).$
Notation 5.6.
The fiber over $u\in G/H$ of ${\mathcal{B}}^{\mathop{\rm r}}_{/H}$ will be denoted $C^{*}_{\mathop{\rm r}}(\mathcal{B}_{u}).$
5.5. Weak containment, nuclearity and amenability
We go back to our original assumption of $H$ being closed in $G$ (it is not normal anymore, unless specifically stated).
Theorem 5.7.
If the normaliser $N$ of $H$ in $G$ is open (e.g. $G/H$ is discrete), then the following claims are true.
  1. (1)
    If $\mathcal{B}$ has the wcp, then $\mathcal{B}_{H}$ has the wcp.
  2. (2)
    If $C^{*}(\mathcal{B})$ is nuclear, then $C^{*}(\mathcal{B}_{H})$ is nuclear.
  3. (3)
    If $C^{*}_{\mathop{\rm r}}(\mathcal{B})$ is nuclear, then $C^{*}_{\mathop{\rm r}}(\mathcal{B}_{H})$ is nuclear.
Proof.
The first claim is [17, Corollary 5.6]. Since $(\mathcal{B}_{N})_{H}=\mathcal{B}_{H}$ and we have conditional expectations, $E\colon C^{*}(\mathcal{B})\to C^{*}(\mathcal{B}_{N})$ and $E_{\mathop{\rm r}}\colon C^{*}_{\mathop{\rm r}}(\mathcal{B})\to C^{*}_{\mathop{\rm r}}(\mathcal{B}_{N})$, we may assume that $N=G$ (i.e. $H$ is normal in $G$). We go a step further and use a trick of [16] to reduce the proof to $H=\{e\}$.
If $C^{*}(\mathcal{B})=C^{*}({\mathcal{B}}_{/H})$ is nuclear, then $C^{*}_{\mathop{\rm r}}({\mathcal{B}}_{/H})$ is nuclear and Theorem 3.21 implies that the fiber $C^{*}(\mathcal{B}_{H})$ of ${\mathcal{B}}_{/H}$ over $H\in G/H$ is nuclear. Replacing ${\mathcal{B}}_{/H}$ with ${\mathcal{B}}^{\mathop{\rm r}}_{/H}$ and recalling that $C^{*}_{\mathop{\rm r}}(\mathcal{B})=C^{*}_{\mathop{\rm r}}({\mathcal{B}}^{\mathop{\rm r}}_{/H})$, a similar argument implies that if $C^{*}_{\mathop{\rm r}}(\mathcal{B})$ is nuclear, then the fiber $C^{*}_{\mathop{\rm r}}(\mathcal{B}_{H})$ of ${\mathcal{B}}^{\mathop{\rm r}}_{/H}$ is nuclear. ∎
Corollary 5.8.
If $H$ is normal in $G$, $C^{*}(\mathcal{B}_{H})$ (resp. $C^{*}_{\mathop{\rm r}}(\mathcal{B}_{H})$) is nuclear and ${\mathcal{B}}_{/H}$ (resp. ${\mathcal{B}}^{\mathop{\rm r}}_{/H}$) is $\mathrm{C}^{*}$-amenable, then $C^{*}(\mathcal{B})$ (resp. $C^{*}_{\mathop{\rm r}}(\mathcal{B})$) is nuclear.
Proof.
The fiber $C^{*}(\mathcal{B}_{H})$ of ${\mathcal{B}}_{/H}$ over the unit $H$ is nuclear and ${\mathcal{B}}_{/H}$ is $\mathrm{C}^{*}$-amenable. By Theorem 3.20 and (5.4), $C^{*}(\mathcal{B})$ is nuclear. On the other hand, $C^{*}_{\mathop{\rm r}}(\mathcal{B}_{H})$ is the fiber of ${\mathcal{B}}^{\mathop{\rm r}}_{/H}$ over $H$ and, by (5.5), $C^{*}_{\mathop{\rm r}}(\mathcal{B})$ is isomorphic to the nuclear algebra $C^{*}_{\mathop{\rm r}}({\mathcal{B}}^{\mathop{\rm r}}_{/H})$. ∎
Theorem 5.9.
If $H$ is normal in $G$, then the following claims are equivalent
  1. (1)
    $\mathcal{B}$ has the wcp.
  2. (2)
    ${\mathcal{B}}_{/H}$ and $\mathcal{B}_{H}$ have the wcp.
  3. (3)
    ${\mathcal{B}}^{\mathop{\rm r}}_{/H}$ and $\mathcal{B}_{H}$ have the wcp.
Proof.
The first condition implies that $\mathcal{B}_{H}$ has the wcp, so we may continue the proof assuming that $\mathcal{B}_{H}$ has the wcp. In this case, there is no difference in the processes of constructing ${\mathcal{B}}_{/H}$ and ${\mathcal{B}}^{\mathop{\rm r}}_{/H}$ out of $\mathcal{L}$ because, for all $f\in\mathcal{L}$, we have $\|S_{f}\|^{2}=\|S_{f^{*}*f}\|=\|f^{*}*f\|_{C^{*}_{\mathop{\rm r}}(\mathcal{B})}=\|f^{*}*f\|_{C^{*}(\mathcal{B}_{H})}=\|\rho(f)\|^{2}.$ Hence, ${\mathcal{B}}_{/H}={\mathcal{B}}^{\mathop{\rm r}}_{/H}$ and the last two claims of the statement are equivalent.
Restricting the isomorphisms $\Omega\colon C^{*}(\mathcal{B})\to C^{*}({\mathcal{B}}_{/H})$ and $\Omega^{\mathop{\rm r}}\colon C^{*}_{\mathop{\rm r}}(\mathcal{B})\to C^{*}_{\mathop{\rm r}}({\mathcal{B}}^{\mathop{\rm r}}_{/H})$ of (5.4) and (5.5) to $C_{c}(\mathcal{B})$ we get $f\mapsto\rho\circ\tilde{f}$. Because ${\mathcal{B}}_{/H}={\mathcal{B}}^{\mathop{\rm r}}_{/H}$, we have a commutative diagram
[…]
and it becomes clear that $\mathcal{B}$ has the wcp if and only if ${\mathcal{B}}_{/H}$ has the wcp. ∎
The preceding results naturally leads to ask if, in the conditions of the Theorem above, the following are equivalent
  1. (1)
    $\mathcal{B}$ is $\mathrm{C}^{*}$-amenable.
  2. (2)
    $\mathcal{B}_{H}$ and ${\mathcal{B}}_{/H}$ are $\mathrm{C}^{*}$-amenable.
  3. (3)
    $\mathcal{B}_{H}$ and ${\mathcal{B}}^{\mathop{\rm r}}_{/H}$ are $\mathrm{C}^{*}$-amenable.
[…]
Proposition 5.10.
If $\mathcal{B}$ is $\mathrm{C}^{*}$-amenable and $H$ normal in $G$, then $\mathcal{B}_{H}$ and ${\mathcal{B}}_{/H}={\mathcal{B}}^{\mathop{\rm r}}_{/H}$ are $\mathrm{C}^{*}$-amenable.
Proof.
Corollary 4.18 implies that $\mathcal{B}_{H}$ is $\mathrm{C}^{*}$-amenable.
Forget about amenability for a moment and take a ccp map $\Phi\colon\mathcal{B}\to\mathcal{B}$. We claim that there exists a unique ccp map ${\Phi}_{/H}\colon{\mathcal{B}}_{/H}\to{\mathcal{B}}_{/H}$ such that for all $u\in G/H$ and $f\in C_{c}(\mathcal{B}_{u})\subseteq L^{1}(\mathcal{B}_{u})$, ${\Phi}_{/H}\circ\rho(f)=\rho(\Phi\circ f)$.
Take a nondegenerate *-representation $T\colon C^{*}(\mathcal{B})\to\mathbb{B}(X)$ with faithful integrated form. Via the isomorphism $L^{1}(\mathcal{L})\cong L^{1}(\mathcal{B})$, $T$ induces a *-representation $T^{1}\colon\mathcal{L}\to\mathbb{B}(X)$, see [16, VIII 15.9], and $T^{1}$ determines a unique *-representation $\dot{T}\colon{\mathcal{B}}_{/H}\to\mathbb{B}(X)$ such that $\dot{T}\circ\rho=T^{1}$. Moreover, the integrated forms of $\dot{T}$ and $T$ are the same up to the isomorphism of (5.4). Since the integrated form of $\dot{T}$ is faithful, the restrictions of $\dot{T}$ to each one of the fibers is isometric. Hence, for all $f\in\uplus_{u\in G/H}C_{c}(\mathcal{B}_{u})$, $\|\rho(\Phi\circ f)\|=\|\dot{T}_{\Phi\circ f}\|$.
Take a nondegenerate dilation of $T\circ\Phi$, $S\colon\mathcal{B}\to\mathbb{B}(Y)$, and $V\in\mathbb{B}(X,Y)$ such that $T\circ\Phi(b)=V^{*}S_{b}V$ and $\|V\|\leq 1$. Using the explicit descriptions $T^{1}$ and $S^{1}$ of [16, VIII 15.9], a straightforward computation reveals that for all $f\in\uplus_{u\in G/H}C_{c}(\mathcal{B}_{u})$, $T^{1}_{\Phi\circ f}=V^{*}S^{1}_{f}V=V^{*}\dot{S}_{\rho(f)}V$ and it follows that $\|\rho(\Phi\circ f)\|\leq\|\rho(f)\|$. Since $\rho(C_{c}(\mathcal{B}))$ is dense in $\dot{B}_{u}$, there exists a unique family of linear contractions $\{\dot{\Phi}_{u}\colon\dot{B}_{u}\to\dot{B}_{u}\}_{u\in G/H}$ whose union ${\Phi}_{/H}\colon{\mathcal{B}}_{/H}\to{\mathcal{B}}_{/H}$ satisfies ${\Phi}_{/H}\circ\rho(f)=\rho(\Phi\circ f)$, for all $f\in\uplus_{u\in G/H}C_{c}(\mathcal{B}_{u})$. For all $f\in C_{c}(\mathcal{B})$ we have ${\Phi}_{/H}\circ(\rho\circ\tilde{f})(u)={\Phi}_{/H}\circ(\rho(f|_{u}))=\rho((\Phi\circ f)|_{u})=\widetilde{\Phi\circ f}(u)$. Considering the family $\Gamma:=\{\rho\circ\tilde{f}\colon f\in C_{c}(\mathcal{B})\}$ and using [16, II 13.16] we get that ${\Phi}_{/H}$ is continuous and it is ccp because $T\circ{\Phi}_{/H}(z)=V^{*}\dot{S}_{z}V$. Notice that ${\mathop{\rm supp}}({\Phi}_{/H})$ is contained in the projection of ${\mathop{\rm supp}}(\Phi)$ into $G/H$.
Let $\{\Phi_{i}\colon\mathcal{B}\to\mathcal{B}\}_{i\in I}$ be a net of ccp maps with compact support converging uniformly to $\operatorname{id}_{\mathcal{B}}$ on compact slices. Then, $\{{\Phi_{i}}_{/H}\colon{\mathcal{B}}_{/H}\to{\mathcal{B}}_{/H}\}_{i\in I}$ is a net of ccp maps with compact supports. To complete the proof we use Lemma 3.4 and the family $\Gamma$ defined above in this proof. For all $f\in C_{c}(\mathcal{B})$ and $t\in G$, setting $g:=\rho\circ\tilde{f}$, we get
$\|{\Phi_{i}}_{/H}\circ g(tH)-g(tH)\|\leq\|T^{1}_{\Phi_{i}\circ(f|_{tH})-f|_{tH}}\|\leq\|\Phi_{i}\circ(f|_{tH})-f|_{tH}\|_{1}.$
and this yields
$\|{\Phi_{i}}_{/H}\circ g-g\|_{1}\leq\int_{G/H}\|{\Phi_{i}}_{/H}\circ g(tH)-g(tH)\|dtH\leq\int_{G/H}\int_{H}\|\Phi_{i}(f(ts))-f(ts)\|dsdtH=\|\Phi_{i}\circ f-f\|_{1}.$
Thus, $\lim_{i}\|{\Phi_{i}}_{/H}\circ g-g\|_{1}=0$. ∎
Appendix A Tensor products of equivariant cp maps between $\mathrm{W}^{*}$-algebras
Let $\gamma$ and $\delta$ be $\mathrm{W}^{*}$-actions of a group $G$ on $M$ and $N,$ respectively, and $\pi\colon M\to N$ an equivariant ccp map. Fix unitary implementations $(M\subseteq\mathbb{B}(X),U)$ and $(N\subseteq\mathbb{B}(Y),V)$ and a Hilbert space $Z$.
Remark A.1 (The coefficient function).
The function $Z\times Z\times\mathbb{B}(Z\otimes X)\to\mathbb{B}(X)$ given by $(z,w,T)\mapsto T_{z,w},$ where $\langle T_{z,w}u,v\rangle=\langle T(w\otimes u),z\otimes v\rangle$, is conjugate linear the first variable and linear in the rest. Moreover, it is a contraction in the sense that $\|T_{z,w}\|\leq\|z\|\|w\|\|T\|$. Hence, $(z,w,T)\mapsto T_{z,w}$ is continuous with respect to the product of the norm topologies.
Recall that given any Hilbert base $S$ of $Z$ we have
$\mathbb{B}(Z)\bar{\otimes}M=\{T\in\mathbb{B}(Z\otimes X)\colon\forall\ z,w\in S,\ T_{z,w}\in M\}.$
Thinking of each $T\in\mathbb{B}(Z\otimes X)$ as the matrix $(T_{z,w})_{z,w\in S}$ one can construct a function
$\operatorname{id}\bar{\otimes}_{S}\pi\colon\mathbb{B}(Z)\bar{\otimes}M\to\mathbb{B}(Z)\bar{\otimes}N$
that in matrix form can be expressed as
(A.2) $(\operatorname{id}\bar{\otimes}_{S}\pi)((T_{z,w})_{z,w\in S})=(\pi(T_{z,w}))_{z,w\in S}.$
In particular, $(\operatorname{id}\bar{\otimes}_{S}\pi)(T\otimes m)=T\otimes\pi(m)$ for all $T\in\mathbb{B}(Z)$ and $m\in M$.
To justify the existence of $\operatorname{id}\bar{\otimes}_{S}\pi$ one can proceed as follows. For each finite $F\subseteq S$ let $P_{F}\in\mathbb{B}(Z)$ be the orthogonal projection onto ${\mathop{\rm span}}(F)$. Since $P_{F}\mathbb{B}(Z)P_{F}$ is a finite dimensional $\mathrm{C}^{*}$-algebra,
$(P_{F}\otimes 1_{X})\mathbb{B}(Z)\bar{\otimes}M(P_{F}\otimes 1_{X})=P_{F}\mathbb{B}(Z)P_{F}{\otimes}M$
and there exists a unique ccp map $\operatorname{id}\otimes\pi\colon P_{F}\mathbb{B}(Z)P_{F}{\otimes}M\to\mathbb{B}(Z)\bar{\otimes}N$ sending $T\otimes m$ to $T\otimes\pi(m)$. For all $T\in\mathbb{B}(Z)\bar{\otimes}M,$ $\|(\operatorname{id}\otimes\pi)((P_{F}\otimes 1_{X})T(P_{F}\otimes 1_{X}))\|\leq\|T\|$. Taking $F$ in the directed set of finite subsets of $S$, some subnet of $\{(\operatorname{id}\otimes\pi)((P_{F}\otimes 1_{X})T(P_{F}\otimes 1_{X}))\}_{F}$ has a limit in the weak operator topology (wot). Expressed in matrix form, that limit must be $(\pi(T_{z,w}))_{z,w\in S}$. This implies that $\{(\operatorname{id}\otimes\pi)((P_{F}\otimes 1_{X})T(P_{F}\otimes 1_{X}))\}_{F}$ wot-converges to $(\pi(T_{z,w}))_{z,w\in S}$ (because all of it subnets have a subnet converging to that operator). Then we can define
$(\operatorname{id}\bar{\otimes}_{S}\pi)(T):=\texttt{wot}\lim_{F}(\operatorname{id}\otimes\pi)((P_{F}\otimes 1_{X})T(P_{F}\otimes 1_{X}))$
and it follows that $(\operatorname{id}\bar{\otimes}_{S}\pi)$ is ccp. Thinking in matrix form, it follows that $\pi$ is normal if and only if $\operatorname{id}\bar{\otimes}_{S}\pi$ is normal. In addition, $\operatorname{id}\bar{\otimes}_{S}\pi=\operatorname{id}$ (the identity map of the domain of $\operatorname{id}\bar{\otimes}_{S}\pi$) if and only if $\pi=\operatorname{id}$.
Lemma A.3.
If $S_{1}$ and $S_{2}$ are Hilbert bases of $Z$, then $\operatorname{id}\bar{\otimes}_{S_{1}}\pi=\operatorname{id}\bar{\otimes}_{S_{2}}\pi$.
Proof.
It suffices to show that for all $T\in\mathbb{B}(Z)\bar{\otimes}M$ and $z,w\in S_{1}$ we have $(\operatorname{id}\bar{\otimes}_{S_{2}}\pi)(T)_{z,w}=(\operatorname{id}\bar{\otimes}_{S_{1}}\pi)(T)_{z,w}.$ Taking limit over the finite sets $F$ of $S_{2}$ and using Remark A.1 we get
$(\operatorname{id}\bar{\otimes}_{S_{2}}\pi)(T)_{z,w}$ $=\lim_{F}(\operatorname{id}\bar{\otimes}_{S_{2}}\pi)(T)_{P_{F}z,P_{F}w}=\lim_{F}\sum_{x,y\in F}\langle z,x\rangle\langle y,w\rangle(\operatorname{id}\bar{\otimes}_{S_{2}}\pi)(T)_{x,y}$
$=\lim_{F}\sum_{x,y\in F}\langle z,x\rangle\langle y,w\rangle\pi(T_{x,y})=\lim_{F}\pi(T_{P_{F}z,P_{F}w})=\pi(T_{z,w})=(\operatorname{id}\bar{\otimes}_{S_{1}}\pi)(T)_{z,w},$
which yields the desired result. ∎
From now on we can safely write $\operatorname{id}\bar{\otimes}\pi$ instead of $\operatorname{id}\bar{\otimes}_{S}\pi.$ The tensor products $\mathbb{B}(Z)\bar{\otimes}M$ and $\mathbb{B}(Z)\bar{\otimes}N$ have canonical structures of $\mathbb{B}(Z)$-bimodules, it turns out that $\operatorname{id}\bar{\otimes}\pi$ is a bimodule map:
Lemma A.4.
For all $R,S\in\mathbb{B}(Z)$ and $T\in\mathbb{B}(Z)\bar{\otimes}M,$
$(\operatorname{id}\bar{\otimes}\pi)((R\otimes 1_{X})T(S\otimes 1_{X}))=(R\otimes 1_{Y})(\operatorname{id}\bar{\otimes}\pi)(T)(R\otimes 1_{Y}).$
Proof.
Since $\operatorname{id}\bar{\otimes}\pi$ preserves the involution, it suffices to consider the case $R=1_{X}$. Moreover, using the norm continuity of $\operatorname{id}\bar{\otimes}\pi$ and Borel Functional Calculus, we can further reduce the proof to the case where $S$ is a projection.
Let $Z^{\prime}$ be the union of Hilbert bases of $SZ$ and $(SZ)^{\perp}$. Take $z,w\in Z^{\prime}$ and $u,v\in Y$. We have
$\langle(\operatorname{id}\bar{\otimes}\pi)(T)(S\otimes 1_{Y})(z\otimes u),w\otimes v\rangle$ $=\langle Sz,z\rangle\langle(\operatorname{id}\bar{\otimes}\pi)(T)(z\otimes u),w\otimes v\rangle$
$=\langle Sz,z\rangle\langle\pi(T_{w,z})u,v\rangle=\langle Sz,z\rangle\langle\pi(T_{w,z})u,v\rangle.$
Once again, replacing $\pi$ with the identity of $M$ we get that $T(S\otimes 1_{X})]_{w,z}=\langle Sz,z\rangle T_{w,z}$ and we obtain
$\langle(\operatorname{id}\bar{\otimes}\pi)(T)(S\otimes 1_{Y})(z\otimes u),w\otimes v\rangle=\langle(\operatorname{id}\bar{\otimes}\pi)(T(S\otimes 1_{X}))(z\otimes u),w\otimes v\rangle;$
which implies $(\operatorname{id}\bar{\otimes}\pi)(T)(S\otimes 1_{Y})=(\operatorname{id}\bar{\otimes}\pi)(T(S\otimes 1_{X}))$. ∎
Corollary A.5.
If $P\subseteq\mathbb{B}(Z)$ is a $\mathrm{W}^{*}$-subalgebra, then $(\operatorname{id}\bar{\otimes}\pi)(P\bar{\otimes}M)\subseteq P\bar{\otimes}N.$
Proof.
Assume that $1_{X}\in P$. We know that $(\operatorname{id}\bar{\otimes}\pi)(P\bar{\otimes}M)\subseteq\mathbb{B}(Z)\bar{\otimes}N$. Combining Lemma A.4 with [22, Corollary 5.9] we obtain $(\operatorname{id}\bar{\otimes}\pi)(P\bar{\otimes}M)\subseteq(P^{\prime}\bar{\otimes}\mathbb{C}1_{Y})^{\prime}=P\bar{\otimes}\mathbb{B}(Y)$. Now that we know
$(\operatorname{id}\bar{\otimes}\pi)(P\bar{\otimes}M)\subseteq(P\bar{\otimes}\mathbb{B}(Y))\cap(\mathbb{B}(Z)\bar{\otimes}N),$
we can use [22, Corollary 5.10] to get $(\operatorname{id}\bar{\otimes}\pi)(P\bar{\otimes}M)\subseteq P\bar{\otimes}N$.
Now assume that the unit $p$ of $P$ is not $1_{Z}$ and let $P\to\tilde{P}\subseteq\mathbb{B}(pZ)$ be given by $m\mapsto\tilde{m}:=m|_{pZ}$. The decomposition $Z=(1_{Z}-p)Z\oplus pZ$ induces $P^{\prime}=\mathbb{B}((1_{Z}-p)Z)\oplus\tilde{P}^{\prime}$ and $P^{\prime\prime}=\mathbb{C}(1_{Z}-p)\oplus\tilde{P}$ is a $\mathrm{W}^{*}$-subalgebra of $\mathbb{B}(Z)$ containing $1_{X}$. Hence, $(\operatorname{id}\bar{\otimes}\pi)(P^{\prime\prime}\bar{\otimes}M)\subseteq P^{\prime\prime}\bar{\otimes}N.$ We know $P=0\oplus\tilde{P}$ and $p=0\oplus\tilde{p}$. Then,
$(\operatorname{id}\bar{\otimes}\pi)(P\bar{\otimes}M)=(\operatorname{id}\bar{\otimes}\pi)((p\otimes 1_{X})P^{\prime\prime}(p\otimes 1_{X}))=(p\otimes 1_{Y})(P^{\prime\prime}\bar{\otimes}N)(p\otimes 1_{X})=(0\oplus\tilde{P})\bar{\otimes}M=P\bar{\otimes}M,$
which completes the proof in the general case. ∎
Notice that
$\operatorname{id}\bar{\otimes}\gamma$ $\colon G\times(\mathbb{B}(Z)\bar{\otimes}M)\to\mathbb{B}(Z)\bar{\otimes}M$ $(t,T)$ $\mapsto(\operatorname{id}\bar{\otimes}\gamma_{t})(T)=(1\otimes U_{t})T(1\otimes U_{t})^{*}$
is a $\mathrm{W}^{*}$-action of $G$ on $\mathbb{B}(Z)\bar{\otimes}M$ with unitary implementation $(\mathbb{B}(Z)\bar{\otimes}M\subseteq\mathbb{B}(Z\otimes X),1\otimes U).$
Lemma A.6.
The map $\operatorname{id}\bar{\otimes}\pi$ is equivariant with respect to $\operatorname{id}\bar{\otimes}\gamma$ and $\operatorname{id}\bar{\otimes}\delta$.
Proof.
Fortunately, $\operatorname{id}\bar{\otimes}\gamma$ and $\operatorname{id}\bar{\otimes}\delta$ are implemented by the unitary representations $t\mapsto 1\otimes U_{t}$ and $t\mapsto 1\otimes V_{t}$. Using (A.2) we get that for all $T\in\mathbb{B}(Z)\bar{\otimes}M$ and $t\in G$
$(\operatorname{id}\bar{\otimes}\pi)((1\otimes U_{t})T(1\otimes U_{t})^{*})$ $=(\operatorname{id}\bar{\otimes}\pi)((U_{t}T_{z,w}U_{t}^{*})_{z,w\in S})=(\pi(U_{t}T_{z,w}U_{t}^{*})_{z,w\in S})$
$=(V_{t}\pi(T_{z,w})V_{t}^{*})_{z,w\in S}=(1\otimes V_{t})\pi(T)(1\otimes V_{t})^{*},$
so $\operatorname{id}\bar{\otimes}_{S}\pi\equiv\operatorname{id}\bar{\otimes}\pi$ is equivariant. ∎
All we have done so far can be repeated flipping the tensors, we mean using $M\bar{\otimes}\mathbb{B}(Z)$ instead of $\mathbb{B}(Z)\bar{\otimes}M$, so we get a ccp map $\pi\bar{\otimes}\operatorname{id}\colon M\bar{\otimes}\mathbb{B}(Z)\to N\bar{\otimes}\mathbb{B}(Z)$. At this point we have all the necessary ingredients to prove equivariant versions of some results by Nagisa and Tomiyama [20, Theorem 3], [23, Theorem 4].
Theorem A.7.
Let (for $i=1,2$) $G_{i}$ be a group and $\pi_{i}\colon M_{i}\to N_{i}$ be a ccp equivariant map with respect to the $\mathrm{W}^{*}$-actions $\gamma^{i}$ and $\delta^{i}$ on $M_{i}$ and $N_{i}$, respectively. Then
  1. (1)
    The (usual) von Neumann tensor product action $\gamma^{1}\bar{\otimes}\gamma^{2}$ of $G_{1}\times G_{2}$ on $M_{1}\bar{\otimes}M_{2}$ is
    $G_{1}\times G_{2}$ $\times(M_{1}\bar{\otimes}M_{2})\to M_{1}\bar{\otimes}M_{2}$ $(t_{1},t_{2},T)$ $\mapsto(\gamma^{1}_{t_{1}}\bar{\otimes}\gamma^{2}_{t_{2}})(T):=(\gamma^{1}_{t_{1}}\bar{\otimes}\operatorname{id})((\operatorname{id}\bar{\otimes}\gamma^{2}_{t_{2}})(T)).$
  2. (2)
    There exists a ccp map $\rho\colon M_{1}\bar{\otimes}M_{2}\to N_{1}\bar{\otimes}N_{2}$ which is $\gamma^{1}\bar{\otimes}\gamma^{2}-\delta^{1}\bar{\otimes}\delta^{2}$ equivariant and for all $m_{1}\in M_{1}$ and $m_{2}\in M_{2}$, $\rho(m_{1}\otimes m_{2})=\pi_{1}(m_{1})\otimes\pi_{2}(m_{2})$.
  3. (3)
    If $\pi_{1}$ and $\pi_{2}$ are normal, then $\rho$ can be made normal.
  4. (4)
    If $\pi_{1}$ and $\pi_{2}$ are (normal) conditional expectations onto $\mathrm{W}^{*}$-subalgebras, then $\rho$ can be made a (normal) conditional expectation.
Proof.
Take unitary implementations $(M_{i}\subseteq\mathbb{B}(X_{i}),U^{i})$ and $(N_{i}\subseteq\mathbb{B}(Y_{i}),V^{i})$. By Corollary A.5, $(\operatorname{id}\bar{\otimes}\pi_{2})(M_{1}\bar{\otimes}M_{2})\subseteq M_{1}\bar{\otimes}N_{2}$ and $(\pi_{2}\bar{\otimes}1)(M_{1}\bar{\otimes}N_{2})\subseteq N_{1}\bar{\otimes}N_{2}.$ Then $\rho\colon M_{1}\bar{\otimes}M_{2}\to N_{1}\bar{\otimes}N_{2},$ $T\mapsto(\pi_{2}\bar{\otimes}1)((\operatorname{id}\bar{\otimes}\pi_{2})(T)),$ is ccp. In addtion, for all $m_{i}\in M_{i}$, $\rho(m_{1}\otimes m_{2})=\pi_{1}(m_{1})\otimes\pi_{2}(m_{2})$. Regarding the normality claims, recall that $\operatorname{id}\bar{\otimes}\pi_{2}$ (respectively, $\pi_{1}\bar{\otimes}1$) is normal if and only if $\pi_{2}$ ($\pi_{1}$) is normal, so $\rho$ is normal if both $\pi_{1}$ and $\pi_{2}$ are normal
After recalling that each $\gamma^{i}_{t}$ is a $\operatorname{\rm w}^{*}$}-homeomorphism (because they are $\mathrm{C}^{*}$-isomorphism) one gets that for all $(t_{1},t_{2},T)\in G_{1}\times G_{2}\times M_{2}\bar{\otimes}M_{2}$, $(\gamma^{1}_{t_{1}}\bar{\otimes}\operatorname{id})((\operatorname{id}\bar{\otimes}\gamma^{2}_{t_{2}})(T))=(U^{1}_{t_{1}}\otimes U^{2}_{t_{2}})T(U^{1}_{t_{1}}\otimes U^{2}_{t_{2}})^{*}.$ Then $(t_{1},t_{2},T)\mapsto(\gamma^{1}_{t_{1}}\bar{\otimes}\operatorname{id})((\operatorname{id}\bar{\otimes}\gamma^{2}_{t_{2}})(T))$ is the action with unitary implementation $(M_{1}\bar{\otimes}M_{2}\subseteq\mathbb{B}(X_{1}\otimes X_{2}),U^{1}\otimes U^{2}),$ i.e. the usual action of $G_{1}\times G_{2}$ on $M_{1}\bar{\otimes}M_{2}$.
To prove $\rho$ is equivariant we take $T\in M_{1}\bar{\otimes}M_{2}$, $t_{1}\in G_{1}$ and $t_{2}\in G_{2}$. Using Lemma A.6 and Corollary A.5 we get
$\rho((\gamma^{1}_{t_{1}}\bar{\otimes}\gamma^{2}_{t_{2}})(T))$ $=(\pi_{2}\bar{\otimes}1)((\operatorname{id}\bar{\otimes}\pi_{2})((U^{1}_{t_{1}}\otimes U^{2}_{t_{2}})T(U^{1}_{t_{1}}\otimes U^{2}_{t_{2}})^{*}))$
$=(\pi_{2}\bar{\otimes}1)((\operatorname{id}\bar{\otimes}\pi_{2})((U^{1}_{t_{1}}\otimes 1)(1\otimes U^{2}_{t_{2}})T(1\otimes U^{2}_{t_{2}})^{*}(U^{1}_{t_{1}}\otimes 1)^{*}))$
$=(\pi_{2}\bar{\otimes}1)((U^{1}_{t_{1}}\otimes 1)(\operatorname{id}\bar{\otimes}\pi_{2})((1\otimes U^{2}_{t_{2}})T(1\otimes U^{2}_{t_{2}})^{*})(U^{1}_{t_{1}}\otimes 1)^{*})$
$=(V^{1}_{t_{1}}\otimes 1)(\pi_{2}\bar{\otimes}1)((1\otimes V^{2}_{t_{2}})(\operatorname{id}\bar{\otimes}\pi_{2})(T)(1\otimes V^{2}_{t_{2}})^{*})(V^{1}_{t_{1}}\otimes 1)^{*}$
$=(V^{1}_{t_{1}}\otimes 1)(1\otimes V^{2}_{t_{2}})(\pi_{2}\bar{\otimes}1)((\operatorname{id}\bar{\otimes}\pi_{2})(T))(1\otimes V^{2}_{t_{2}})^{*}(V^{1}_{t_{1}}\otimes 1)^{*}$
$=(\delta^{1}_{t_{1}}\bar{\otimes}\delta^{2}_{t_{2}})(\rho(T)),$
so $\rho$ is equivariant.
In case $\pi_{1}$ and $\pi_{2}$ are conditional expectations, let $p_{i}$ be the unit of $N_{i}$. Then we may take $Y_{i}=p_{i}X_{i}$ and use the implementations $(N_{i}\subseteq\mathbb{B}(pX_{i}),t\mapsto pU^{i}_{t}p)$. We are implicitly considering two faithful normal representations of $N_{1}\bar{\otimes}N_{2},$ one in $\mathbb{B}(p_{1}X_{1}\bar{\otimes}p_{2}X_{2})$ and one in $\mathbb{B}(X_{1}\otimes X_{2})$. Formally, (A.2) implies that for every $T\in N_{1}\bar{\otimes}N_{2}\subseteq\mathbb{B}(X_{1}\otimes X_{2})$, $\rho(T)=(\pi_{1}\bar{\otimes}1)(\operatorname{id}\bar{\otimes}\pi_{2}(T))=(\pi_{1}\bar{\otimes}1)(T)=T\in\mathbb{B}(p_{1}X_{1}\bar{\otimes}p_{2}X_{2})$. Then $\rho$ is a conditional expectation. ∎
Lemma 2.1 of [7] follows from the Theorem above by taking $G_{1}=G_{2}$ and identifying $G_{1}\cong\{(t,t)\colon t\in G_{1}\}\subseteq G_{1}\times G_{1}$. After this one gets [7, Proposition 3.9], which can be stated as
Corollary A.8.
If, for $i=1,2$, $\gamma^{i}$ is a $\mathrm{W}^{*}$-action of $G$ on $M_{i}$ and at least one of them is $\mathrm{W}^{*}$-amenable, then the diagonal action $\gamma^{1}\bar{\otimes}^{d}\gamma^{2}:=(\gamma^{1}\bar{\otimes}\gamma^{2})|_{G}$ of $G$ on $M_{1}\bar{\otimes}M_{2}$ is $\mathrm{W}^{*}$-amenable.
The “non-diagonal” version would be
Corollary A.9.
If, for $i=1,2$, $\gamma^{i}$ is a $\mathrm{W}^{*}$-amenable action of $G_{i}$ on $M_{i}$, then action $\gamma^{1}\bar{\otimes}\gamma^{2}$ of $G_{1}\times G_{2}$ on $M_{1}\bar{\otimes}M_{2}$ is $\mathrm{W}^{*}$-amenable.
Proof.
Use Theorem A.7 and the canonical isomorphism
$(L^{\infty}(G_{1})\bar{\otimes}M_{1})\bar{\otimes}(L^{\infty}(G_{2})\bar{\otimes}M_{2})\cong L^{\infty}(G_{1}\times G_{2})\bar{\otimes}(M_{1}\bar{\otimes}M_{2}),$
the details are left to the reader. ∎
Appendix B An extension Lemma for the continuous part of a $\mathrm{W}^{*}$-action
Let $\gamma$ be a $\mathrm{W}^{*}$-action of $G$ on $M$ and $(M\subseteq\mathbb{B}(X),U)$ a unitary implementation. We briefly recall some facts about the continuous part of $\gamma$.
As in [8] we use the bilinear function $L^{1}(G)\times M\to M,\ (f,m)\mapsto\gamma_{f}(m),$ such that for all $\varphi\in M_{*},\ f\in L^{1}(G)$ and $m\in M$, $\varphi(\gamma_{f}(m))=\int_{G}\varphi(\gamma_{t}(m))f(t)dt$. One way of proving the existence of such function is by using Riesz’s representation Theorem to show that for each $(f,m)\in L^{1}(G)\times M$ there exists a unique operator $I^{U}_{f}(m)\in\mathbb{B}(X)$ such that $\langle x,I^{U}_{f}(m)y\rangle=\int_{G}\langle U_{t}^{*}x,mU_{t}^{*}y\rangle f(t)dt=\int_{G}\langle x,\gamma_{t}(m)y\rangle f(t)dt$. Since $\|\langle x,\gamma_{t}(m)y\rangle\|\leq\|m\|\|x\|\|y\|,$ we have $\|I^{U}_{f}(m)\|\leq\|m\|\|f\|_{1}$. If $\omega$ is an ultra weak continuous functional of $\mathbb{B}(X)$ such that $\omega(M)=\{0\},$ then $\omega(T)=\sum_{n\in\mathbb{N}}\langle x_{n},Ty_{n}\rangle$ for sequences $\{x_{n}\}_{n\in\mathbb{N}},\{y_{n}\}_{n\in\mathbb{N}}\subset X$ such that $\sum_{n\in\mathbb{N}}\|x_{n}\|^{2}+\|y_{n}\|^{2}<\infty$; and $\omega(I^{U}_{f}(m))=\int_{G}\omega(\gamma_{t}(m))f(t)dt=0$ and it follows that $I^{U}_{f}(m)\in M$. So we can set $\varphi_{f}(m):=I^{U}_{f}(m)$.
For $f,g\in L^{1}(G)$ and $s\in G$ we define $sf,fs\in L^{1}(G)$ by $sf(t)=f({s^{-1}}t)$ and $fs(t)=f(t{s^{-1}})\Delta(s)^{-1}$; while fgf*g is the convolution product. With this notation, for all $m\in M$ one has
$\|\gamma_{f}(m)\|$ $\leq\|m\|\|f\|_{1}$ $\gamma_{s}(\gamma_{f}(m))=$ $\gamma_{sf}(m)$ $\gamma_{f}(\gamma_{s}(m))$ $=\gamma_{fs}(m)$ $\gamma_{f}(\gamma_{g}(m))$ $=\gamma_{f*g}(m).$
In particular, $\|\gamma_{r}(\gamma_{f}(m))-\gamma_{s}(\gamma_{f}(m))\|=\|\gamma_{rf-sf}(m)\|\leq\|m\|\|rf-sf\|_{1}.$ For $f\in C_{c}(G),$ $r\mapsto rf$, is $\|\ \|_{1}$-continuous and it follows that $\{\gamma_{f}(m)\colon f\in C_{c}(G),\ m\in M\}\subseteq M^{c}.$ Using an approximate unit of $L^{1}(G)$ contained in $C_{c}(G)$ one can show that $M^{c}$ is $\operatorname{\rm w}^{*}$}-dense in $M.$ For each $f\in C_{c}(G)^{+},$ $\gamma_{f}\colon M\to M^{c}$ is completely positive because for all $m_{1},\ldots,m_{n}\in M$ and $x_{1},\ldots,x_{n}\in X$ we have $\sum_{i,j=1}^{n}\langle x_{i},\gamma_{f}(m_{i}^{*}m_{j})x_{j}\rangle=\int_{G}\|\sum_{i=1}^{n}m_{i}U_{t}^{*}x_{i}\|^{2}f(t)dt\geq 0$.
The idea of the proof below comes from that of [8, Lemme 2.1].
Lemma B.1.
If $(Y,\rho_{c},V)$ is a unital covariant representation of $\gamma^{c}$, then there exists an equivariant ucp extension $\rho\colon M\to\mathbb{B}(X)$ of $\rho_{c}$ with $\rho(M)^{\prime}=\rho_{c}(M^{c})^{\prime}$.
Proof.
Let $I$ be the set of compact neighborhoods of $e\in G$ with the order $i\leq j\Leftrightarrow j\subseteq i$. Fix a net $\{f_{i}\}_{i\in I}\subseteq C_{c}(G)$ such that $f_{i}\geq 0$; $\int_{G}f_{i}(t)dt=1$ and ${\mathop{\rm supp}}(f_{i})\subseteq i$. The net $\{\rho_{c}\circ\gamma_{f_{i}}\}_{i\in I}$ is contained in the set of ucp maps from $M$ to $\mathbb{B}(X)$ and this set is compact with the topology of pointwise $\operatorname{\rm w}^{*}$}-convergence. Hence, there exists a subnet $\{i_{j}\}_{j\in J}$ and a ucp map $\rho\colon M\to\mathbb{B}(X)$ such that for all $m\in M$, $\rho(m)={\operatorname{\rm w}^{*}}\lim_{j}\rho_{c}(\gamma_{f_{i_{j}}}(m))$. For all $m\in M^{c}$ and $x,y\in X$ we have $\gamma_{f_{i}}(m)=\int_{G}\gamma_{t}(m)f_{i}(t)dt,$ because $t\mapsto\gamma_{t}(m)f_{i}(t)$ is norm-continuous and has compact support. This implies
$\langle\rho(m)x,y\rangle=\lim_{j}\int_{G}\langle\rho_{c}(m)U_{t}^{*}x,U_{t}^{*}y\rangle f_{i_{j}}(t)dt=\langle\rho_{c}(m)U_{e}^{*}x,U_{e}^{*}y\rangle=\langle\rho_{c}(m)x,y\rangle;$
and it follows that $\rho$ is an extension of $\rho_{c}$.
To prove that $\rho$ is equivariant we take $m\in M$, $t\in G$ and $x,y\in X$. We have $\langle\rho(\gamma_{t}(m))x,y\rangle=\lim_{j}\langle\rho_{c}(\gamma_{f_{i_{j}}}(\gamma_{t}(m)))x,y\rangle=\lim_{j}\langle\rho_{c}(\gamma_{f_{i_{j}}t}(m))x,y\rangle$. The net $\{f_{i_{j}}t*f_{i_{k}}\}_{k\in J}$ converges to $f_{i_{j}}t$ in $L^{1}(G)$, so $\gamma_{f_{i_{j}}t*f_{i_{k}}}(m)=\gamma_{f_{i_{j}}t}(\gamma_{f_{i_{k}}}(m))$ converges to $\gamma_{f_{i_{j}}t}(m)\in M^{c}$ in norm. Besides, $s\mapsto\gamma_{s}(\gamma_{f_{i_{k}}}(m))f_{i_{j}}t(s)$ is norm continuous and has compact support, so it is integrable and its norm-integral is $\gamma_{f_{i_{j}}t}(\gamma_{f_{i_{k}}}(m)).$ Combining these facts we get
$\langle\rho(\gamma_{t}(m))x,y\rangle$ $=\lim_{j}\lim_{k}\langle\rho_{c}(\gamma_{f_{i_{j}}t}(\gamma_{f_{i_{k}}}(m)))x,y\rangle=\lim_{j}\lim_{k}\langle\rho_{c}(\int_{G}\gamma_{s}(\gamma_{f_{i_{k}}}(m))f_{i_{j}}t(s)ds)x,y\rangle$
$=\lim_{j}\lim_{k}\int_{G}\langle U_{s}\rho_{c}(\gamma_{f_{i_{k}}}(m)))U_{s}^{*}x,y\rangle f_{i_{j}}(s{t^{-1}})\Delta(t)^{-1}ds$
$=\lim_{j}\lim_{k}\int_{G}\langle U_{s}U_{t}\rho_{c}(\gamma_{f_{i_{k}}}(m)))U_{t}^{*}U_{s}^{*}x,y\rangle f_{i_{j}}(s)ds.$
Notice that $\{U_{t}\rho_{c}(\gamma_{f_{i_{k}}}(m)))U_{t}^{*}\}_{k\in K}$ is a bounded net with wot-limit $U_{t}\rho(m)U_{t}^{*}$. Besides, for all $j\in J$, the support of $f_{i_{j}}$ is compact and $s\mapsto U_{s}^{*}x$ is continuous. Combining these facts we can prove that the net $\{s\mapsto\langle U_{st}\rho_{c}(\gamma_{f_{i_{k}}}(m)))U_{st}^{*}x,y\rangle f_{i_{j}}(s)\}_{k\in K}\subset C_{c}(G)$ converges uniformly to $s\mapsto\langle U_{st}\rho(m)U_{st}^{*}x,y\rangle f_{i_{j}}(s)$. Thus, $\rho$ is equivariant because
$\langle\rho(\gamma_{t}(m))x,y\rangle=\lim_{j}\int_{G}\langle U_{s}U_{t}\rho(m)U_{t}^{*}U_{s}^{*}x,y\rangle f_{i_{j}}(s)ds=\langle U_{t}\rho(m)U_{t}^{*}x,y\rangle.$
Clearly, $\rho_{c}(M^{c})\subseteq\rho(M),$ so $\rho(M)^{\prime}\subseteq\rho_{c}(M^{c})^{\prime}.$ To prove the converse inclusion we notice that for all $m\in M,$ $\rho(m)=\texttt{wot}\lim_{j}\rho_{c}(\gamma_{f_{i_{j}}}(m))\in\overline{\rho_{c}(M)}^{\texttt{wot}}=\rho_{c}(M)^{\prime\prime}.$ Then $\rho(M)\subseteq\rho_{c}(M)^{\prime\prime}$ and it follows that $\rho_{c}(M)^{\prime}=\rho_{c}(M)^{\prime\prime\prime}\subseteq\rho(M)^{\prime}.$

### v2

@@
 Abstract.
-We prove that for a locally compact group $G$, a $\mathrm{C}^{*}$-dynamical system $(A,G,\alpha)$ is amenable in the sense of Anantharaman-Delaroche if and only if, for every other system $(B,G,\beta)$, the diagonal system $(A\otimes_{\max}B,G,\alpha\otimes^{d}_{\max}\beta)$ has the weak containment property (wcp).
+We prove that for a locally compact group $G$, a $\mathrm{C}^{*}$-dynamical system $(A,G,\alpha)$ is amenable if and only if, for every other system $(B,G,\beta)$, the diagonal system $(A\otimes_{\max}B,G,\alpha\otimes^{d}_{\max}\beta)$ has the weak containment property (wcp).
 For Fell bundles over $G$, we construct a diagonal tensor product $\otimes^{d}_{\max}$ and show that a Fell bundle $\mathcal{A}$ has the positive approximation property of Exel and Ng (AP) precisely when $\mathcal{A}\otimes^{d}_{\max}\mathcal{B}$ has the wcp for every Fell bundle $\mathcal{B}$ over $G$.
@@
 Amenability is a central notion connecting group theory, operator algebras and noncommutative dynamics.
-Since the foundational work of Anantharaman-Delaroche on amenable actions of locally compact groups on $\mathrm{C}^{*}$- and $\mathrm{W}^{*}$-algebras [7, 8], a rich theory has developed relating amenability to approximation properties, tensorial behaviour and structural properties of crossed products.
+In the setting of actions of locally compact groups on von Neumann algebras, amenability originates in Zimmer’s work on measurable dynamics and was studied systematically by Anantharaman-Delaroche; see [7, 8].
+In this context, amenability is often referred to as Zimmer amenability.
+For actions on $\mathrm{C}^{*}$-algebras, Anantharaman-Delaroche introduced an amenability notion for discrete groups via the bidual [9].
+However, for a general locally compact group $G$ this approach does not extend directly, since the induced action on the bidual $A^{**}$ of a $G$-$\mathrm{C}^{*}$algebra AA need not be $w^{*}$-continuous.
+This conceptual difficulty was resolved more recently by Buss-Echterhoff-Willett [12] and, independently, by Ozawa-Suzuki [23], building on Ikunishi’s theory [19] of the enveloping $G$-$\mathrm{W}^{*}$-algebra $A^{\prime\prime}_{\alpha}$.
+These works show that several natural formulations of amenability for $\mathrm{C}^{*}$-dynamical systems of locally compact groups – including definitions via the enveloping $\mathrm{W}^{*}$-algebra and approximation properties – are equivalent.
+Since these foundational developments, a rich theory has emerged relating amenability to approximation properties, tensorial behaviour, and structural features of crossed products.
 The goal of this paper is to provide new, unified characterizations of amenability both for $\mathrm{C}^{*}$-dynamical systems and for Fell bundles over locally compact groups, and to exploit those characterizations to obtain permanence and nuclearity results.
+Throughout the paper, the term amenable for $\mathrm{C}^{*}$-actions is used in this modern sense, and we freely rely on the equivalent characterizations established in the works cited above.
 Our strategy is twofold.
@@
   1. (1)
-    $\alpha$ is amenable in the sense of Anantharaman-Delaroche (equivalently, the enveloping $\mathrm{W}^{*}$-action $\alpha^{\prime\prime}$ on $A^{\prime\prime}_{\alpha}$ is $\mathrm{W}^{*}$-amenable).
+    $\alpha$ is amenable (i.e., the enveloping $\mathrm{W}^{*}$-action $\alpha^{\prime\prime}$ on $A^{\prime\prime}_{\alpha}$ is $\mathrm{W}^{*}$-amenable in the sense of Anantharaman-Delaroche).
   2. (2)
@@
 It was first noticed in [7, 8] that the amenability of $\gamma$ is related to that of the restriction of $\gamma$ to the center of $M$, $\gamma|_{Z(M)}$.
-Anantharaman-Delaroche proved they are equivalent under separability conditions and the general case is a consequence of [11, Theorem 3.17].
+Anantharaman-Delaroche proved they are equivalent under separability conditions and the general case is a consequence of [12, Theorem 3.17].
 The continuous part of $\gamma$, denoted $\gamma^{c}$, is the restriction of $\gamma$ to the $\mathrm{C}^{*}$-subalgebra $M^{c}\subseteq M$ formed by those $m\in M$ such that $t\mapsto\gamma_{t}(m)$ is norm continuous.
@@
 In Appendix A, we address the problem of constructing the (von Neumann) tensor product of possibly non-normal, equivariant, contractive completely positive (ccp) maps between $\mathrm{W}^{*}$-algebras.
-Our approach combines techniques from [8, Lemme 2.1], [20, Theorem 3], and [23, Theorem 4].
+Our approach combines techniques from [8, Lemme 2.1], [22, Theorem 3], and [25, Theorem 4].
 While these methods are likely familiar to experts, we were unable to find a precise reference in the literature.
@@
 Let $\iota^{c}\colon M^{c}\to\mathbb{B}(X)$ be the inclusion and $\psi_{0}\colon C\to\mathbb{B}(X)$ the unique *-homomorphism such that $\psi_{0}\circ(1\rho\rtimes\lambda V)=(\pi\times\iota^{c})\rtimes U$; where $\pi\times\iota^{c}(a\otimes m)=\pi(a)m$.
-Fix a ccp extension $\psi\colon\mathbb{B}(L^{2}(G)\otimes Y)\to\mathbb{B}(X)$ of $\psi_{0}$ and name $D$ the multiplicative domain of $\psi$ [10, Proposition 1.5.7 & Definition 1.5.8]. Recall that
+Fix a ccp extension $\psi\colon\mathbb{B}(L^{2}(G)\otimes Y)\to\mathbb{B}(X)$ of $\psi_{0}$ and name $D$ the multiplicative domain of $\psi$ [11, Proposition 1.5.7 & Definition 1.5.8]. Recall that
 $D$ $=\{T\in\mathbb{B}(L^{2}(G)\otimes Y)\colon\psi(T^{*}T)=\psi(T)^{*}\psi(T),\ \psi(TT^{*})=\psi(T)\psi(T)^{*}\}$
@@
 Proof.
-The forward implication is already known [11, Theorem 7.2], but we provide a proof here for the convenience of the reader.
+The forward implication is already known [12, Theorem 7.2], but we provide a proof here for the convenience of the reader.
 Take any $\mathrm{C}^{*}$-algebra $B$ and equip it with the trivial action $\beta$ of $G$.
@@
 3. Amenability of Fell bundles and $\mathrm{C}^{*}$-partial actions
-When we say that $\mathcal{B}=\{B_{t}\}_{t\in G}$ is a Fell bundle we mean that the disjoint union $\bigsqcup_{t\in G}B_{t}$ forms a $\mathrm{C}^{*}$-algebraic bundle over $G$ in the sense of [16, Ch.
+When we say that $\mathcal{B}=\{B_{t}\}_{t\in G}$ is a Fell bundle we mean that the disjoint union $\bigsqcup_{t\in G}B_{t}$ forms a $\mathrm{C}^{*}$-algebraic bundle over $G$ in the sense of [17, Ch.
 VIII, 16.2], with fibers $B_{t}$.
@@
 $\lambda^{\mathcal{B}}\colon C^{*}(\mathcal{B})\longrightarrow C^{*}_{\mathop{\rm r}}(\mathcal{B})$
-the canonical surjective -homomorphism given by the reduced representation [15, Definition 2.7].
+the canonical surjective -homomorphism given by the reduced representation [16, Definition 2.7].
 We say that $\mathcal{B}$ has the weak containment property (wcp) if $\lambda^{\mathcal{B}}$ is injective, that is, if the full and reduced cross-sectional $\mathrm{C}^{*}$-algebras of $\mathcal{B}$ coincide.
 Exel and Ng referred to this property as amenability of the Fell bundle, but in order to avoid confusion with other notions of amenability used in this article, we adopt the terminology wcp instead (see Definition 3.10 and Remark 3.11).
-Whenever we say that a Fell bundle has an approximation property in the sense of Exel and Ng, we mean one of those introduced in [15, Definition 3.6]; by [15, Theorem 3.9], each of them implies the wcp.
+Whenever we say that a Fell bundle has an approximation property in the sense of Exel and Ng, we mean one of those introduced in [16, Definition 3.6]; by [16, Theorem 3.9], each of them implies the wcp.
 The $L^{1}$-approximation property, introduced by Abadie in [1, Definition 4.9(3)], also implies the wcp [1, Theorem 4.11], and is formally weaker than the Exel–Ng approximation properties [1, Proposition 4.10].
@@
 Before stating our main theorem for Fell bundles, we recall the relevant approximation properties.
-We use [16, VIII 3.8] to regard the multiplier algebra $\mathcal{M}(B_{e})$ of the fiber over the unit $e\in G$ as the $\mathrm{C}^{*}$-algebra of multipliers of $\mathcal{B}$ of order $e$.
+We use [17, VIII 3.8] to regard the multiplier algebra $\mathcal{M}(B_{e})$ of the fiber over the unit $e\in G$ as the $\mathrm{C}^{*}$-algebra of multipliers of $\mathcal{B}$ of order $e$.
 As shown in [1], the natural actions $\mathcal{M}(B_{e})\times\mathcal{B}\to\mathcal{B}$ and $\mathcal{B}\times\mathcal{M}(B_{e})\to\mathcal{B}$ are continuous.
@@
 $\|\Phi\|:=\sup\{\,\|\Phi(b)\|\colon b\in\mathcal{B},\,\|b\|\leq 1\}<\infty.$
-A net $\{\Phi_{i}\}_{i\in I}\subset\mathrm{CLB}(\mathcal{B})$ converges to $\Phi\in\mathrm{CLB}(\mathcal{B})$ on compact slices [15, Definition 3.4] if
+A net $\{\Phi_{i}\}_{i\in I}\subset\mathrm{CLB}(\mathcal{B})$ converges to $\Phi\in\mathrm{CLB}(\mathcal{B})$ on compact slices [16, Definition 3.4] if
 $\lim_{i}\|\Phi_{i}\circ f-\Phi\circ f\|_{\infty}=0\qquad\text{for all }f\in C_{c}(\mathcal{B}).$
@@
 Proof.
-By [16, II 14.6], $\Gamma^{\prime}$ is dense in $C_{c}(\mathcal{B})$ for the inductive limit topology, hence also dense for both $\|\cdot\|_{\infty}$ and $\|\cdot\|_{1}$.
+By [17, II 14.6], $\Gamma^{\prime}$ is dense in $C_{c}(\mathcal{B})$ for the inductive limit topology, hence also dense for both $\|\cdot\|_{\infty}$ and $\|\cdot\|_{1}$.
 Let $\|\cdot\|$ denote either of these norms.
@@
 In particular, if $\alpha$ is already a $\mathrm{C}^{*}$-action, then $\beta$ is Morita equivalent to $\alpha$.
-Since $\mathrm{C}^{*}$-amenability is invariant under Morita equivalence of $\mathrm{C}^{*}$-actions [11, Proposition 3.20], Definition 3.10 is consistent with the usual notion of amenability for $\mathrm{C}^{*}$-actions.
+Since $\mathrm{C}^{*}$-amenability is invariant under Morita equivalence of $\mathrm{C}^{*}$-actions [12, Proposition 3.20], Definition 3.10 is consistent with the usual notion of amenability for $\mathrm{C}^{*}$-actions.
 More generally, a $\mathrm{C}^{*}$-partial action $\alpha$ is $\mathrm{C}^{*}$-amenable if and only if any (equivalently, all) of its Morita enveloping actions are $\mathrm{C}^{*}$-amenable.
 Each of the equivalent conditions in Theorem 3.2 can be reformulated in terms of $\alpha$.
-For instance, for $\mathrm{C}^{*}$-actions the positive approximation property of Exel and Ng coincides with the AP of [11, Definition 4.3].
-(To get the equivalence, one may apply Lemma 3.4 with $\Gamma\subset C_{c}(\mathcal{B})$ consisting of sections of the form $t\mapsto f(t)a\delta_{t}$, where $f\in C_{c}(G)$ and $a\in A$.) On the other hand, some other approximation properties – such as the QAP of [11, Definition 4.1] or the weak versions of the AP and the QAP – do not correspond to the translated versions of the properties in Theorem 3.2.
+For instance, for $\mathrm{C}^{*}$-actions the positive approximation property of Exel and Ng coincides with the AP of [12, Definition 4.3].
+(To get the equivalence, one may apply Lemma 3.4 with $\Gamma\subset C_{c}(\mathcal{B})$ consisting of sections of the form $t\mapsto f(t)a\delta_{t}$, where $f\in C_{c}(G)$ and $a\in A$.) On the other hand, some other approximation properties – such as the QAP of [12, Definition 4.1] or the weak versions of the AP and the QAP – do not correspond to the translated versions of the properties in Theorem 3.2.
 Remark 3.13.
@@
 By [5, Theorem 4.2], $\alpha$ is Morita equivalent to $\beta$.
-Combining this with [11, Proposition 3.20] and Theorem 3.2, we obtain the chain of equivalences
+Combining this with [12, Proposition 3.20] and Theorem 3.2, we obtain the chain of equivalences
 […]
@@
 By [3, Corollary 5.2], $\mathbb{k}(\mathcal{B})$ is nuclear.
-Since $\beta$ is $\mathrm{C}^{*}$-amenable, [11, Theorem 7.2] implies that $C^{*}(\mathcal{C})$ is nuclear.
+Since $\beta$ is $\mathrm{C}^{*}$-amenable, [12, Theorem 7.2] implies that $C^{*}(\mathcal{C})$ is nuclear.
 Using [6, Theorem 4.5] we get that $C^{*}(\mathcal{B})$ is nuclear.
@@
 By [6, Proposition 4.13], $C^{*}_{\mathop{\rm r}}(\mathcal{B})$ is Morita equivalent to $C^{*}_{\mathop{\rm r}}(\mathcal{C})=\mathbb{k}(\mathcal{B})\rtimes_{{\mathop{\rm r}}\beta}G$.
-Hence, $\mathbb{k}(\mathcal{B})\rtimes_{{\mathop{\rm r}}\beta}G$ is nuclear and [18, Theorem 3.4] implies that $\mathbb{k}(\mathcal{B})$ is nuclear (unfortunately, in that article nuclear $\mathrm{C}^{*}$-algebras are called amenable).
+Hence, $\mathbb{k}(\mathcal{B})\rtimes_{{\mathop{\rm r}}\beta}G$ is nuclear and [20, Theorem 3.4] implies that $\mathbb{k}(\mathcal{B})$ is nuclear (unfortunately, in that article nuclear $\mathrm{C}^{*}$-algebras are called amenable).
 Corollary 5.2 of [3] implies $B_{e}$ is nuclear.
-Finally, if $G$ is inner amenable, then [19, Corollary 6.6] implies $\beta$ is $\mathrm{C}^{*}$-amenable.
+Finally, if $G$ is inner amenable, then [21, Corollary 6.6] implies $\beta$ is $\mathrm{C}^{*}$-amenable.
 ∎
@@
 Moreover, if $T|_{B_{e}}$ is faithful, then by the $\mathrm{C}^{*}$-identity each restriction $T|_{B_{t}}$ is isometric, and hence $T^{t}$ is faithful.
-The following definition extends [12, Definition A.2] to Fell bundles over locally compact groups.
+The following definition extends [13, Definition A.2] to Fell bundles over locally compact groups.
 Definition 4.1.
@@
 As sets, $\mathcal{B}$ and $\mathcal{B}^{d}$ coincide.
-By [12, Theorem A.4], there exist a *-representation $T^{d}\colon\mathcal{B}^{d}\to\mathbb{B}(Y)$ and an operator $V\in\mathbb{B}(X,Y)$ such that $Y=\overline{\mathop{\rm span}}\,T^{d}(\mathcal{B}^{d})VX$ and $\phi(b)=V^{*}T^{d}_{b}V$ for all $b\in\mathcal{B}^{d}$.
+By [13, Theorem A.4], there exist a *-representation $T^{d}\colon\mathcal{B}^{d}\to\mathbb{B}(Y)$ and an operator $V\in\mathbb{B}(X,Y)$ such that $Y=\overline{\mathop{\rm span}}\,T^{d}(\mathcal{B}^{d})VX$ and $\phi(b)=V^{*}T^{d}_{b}V$ for all $b\in\mathcal{B}^{d}$.
 Define $T\colon\mathcal{B}\to\mathbb{B}(Y)$ by $T_{b}:=T^{d}_{b}$.
@@
 Naturally, we name $S\colon\mathcal{A}\bigotimes_{\max}\mathcal{C}\to\mathbb{B}(Y)$ the disintegrated form of $\tilde{S}$.
-Using [16, VIII 11.8] and the fact that $\Psi(f\oslash g)=(\Phi\circ f)\oslash g$ get get $V^{*}S_{f(s)\otimes g(t)}V=T_{\Phi(f(s))\otimes g(t)}$, for all $f\in C_{c}(\mathcal{A})$, $g\in C_{c}(\mathcal{C})$ and $(s,t)\in G\times H$.
+Using [17, VIII 11.8] and the fact that $\Psi(f\oslash g)=(\Phi\circ f)\oslash g$ get get $V^{*}S_{f(s)\otimes g(t)}V=T_{\Phi(f(s))\otimes g(t)}$, for all $f\in C_{c}(\mathcal{A})$, $g\in C_{c}(\mathcal{C})$ and $(s,t)\in G\times H$.
 This implies $V^{*}S(A_{s}\otimes_{\max}C_{t})V\subseteq T(B_{s}\otimes_{\max}C_{t})$, then there exists a function $\Phi\bigotimes_{\max}\operatorname{id}_{\mathcal{C}}\colon\mathcal{A}\bigotimes_{\max}\mathcal{C}\to\mathcal{B}\bigotimes_{\max}\mathcal{C}$ such that $(\Phi\bigotimes_{\max}\operatorname{id}_{\mathcal{C}})(A_{s}\otimes_{\max}C_{t})\subseteq B_{s}\otimes_{\max}C_{t}$ and $T\circ\Phi(z)=V^{*}S_{z}V,$ for all $(s,t)\in G\times H$ and $z\in\mathcal{A}\bigotimes_{\max}\mathcal{C}$.
 This implies $(\Phi\bigotimes_{\max}\operatorname{id}_{\mathcal{C}})(a\otimes c)=\Phi(a)\otimes c$ and $\|\Phi(z)\|\leq\|V\|^{2}\|z\|=\|\tilde{\Phi}\|\|z\|=\|\Phi|_{A_{e}}\|\|z\|$.
-We have $(\Phi\bigotimes_{\max}\operatorname{id}_{\mathcal{C}})(f\oslash g)=(\Phi\circ f)\oslash g$, so [16, II 13.16] implies $\Phi\bigotimes_{\max}\operatorname{id}_{\mathcal{C}}$ is continuous and it is cp because $T\circ\Phi(z)=V^{*}S_{z}V$ (see Remark 4.8).
+We have $(\Phi\bigotimes_{\max}\operatorname{id}_{\mathcal{C}})(f\oslash g)=(\Phi\circ f)\oslash g$, so [17, II 13.16] implies $\Phi\bigotimes_{\max}\operatorname{id}_{\mathcal{C}}$ is continuous and it is cp because $T\circ\Phi(z)=V^{*}S_{z}V$ (see Remark 4.8).
 Finally, $\|(\Phi\bigotimes_{\max}\operatorname{id}_{\mathcal{C}})|_{A_{e}\otimes_{\max}C_{e}}\|=\|\Phi|_{A_{e}}\otimes_{\max}\operatorname{id}_{C_{e}}\|=\|\Phi|_{A_{e}}\|$.
@@
 Notice that $\Phi_{i}(\mathcal{B}_{H})\subseteq\mathcal{B}_{H}$ and that for all $n\in\mathbb{N}$ and $t\in H^{n},$ $\mathbb{M}_{t}(\mathcal{B})=\mathbb{M}_{t}(\mathcal{B}_{H})$.
-By [16, II 14.8], $\{f|_{H}\colon f\in C_{c}(\mathcal{B})\}=C_{c}(\mathcal{B}_{H})$.
+By [17, II 14.8], $\{f|_{H}\colon f\in C_{c}(\mathcal{B})\}=C_{c}(\mathcal{B}_{H})$.
 Then $\{\mathcal{B}_{H}\to\mathcal{B}_{H},b\mapsto\Phi_{i}(b)\}_{i\in I}$ is a net witnessing the BCAP of $\mathcal{B}_{H}$.
@@
 Assume $H$ is open in $G$ and think $C_{c}(\mathcal{B}_{H})\subseteq C_{c}(\mathcal{B})$ by extending each $f\in C_{c}(\mathcal{B}_{H})$ as 0 outside $H.$ This gives $L^{1}(\mathcal{B}_{H})\subseteq L^{1}(\mathcal{B})$ and, passing to enveloping $\mathrm{C}^{*}$-algebras, we get a *-homomorphism $\iota\colon C^{*}(\mathcal{B}_{H})\to C^{*}(\mathcal{B})$.
-To prove $\iota$ is injective we take a nondegenerate *-representation $T\colon C^{*}(\mathcal{B}_{H})\to\mathbb{B}(X)$ with faithful integrated form $\tilde{T}\colon C^{*}(\mathcal{B}_{H})\to\mathbb{B}(X).$ By [17, Theorem 2.1], $T$ is $\mathcal{B}$-positive and it can be induced to a *-representation $S\colon\mathcal{B}\to\mathbb{B}(Y)$ by the process of [16, Ch.
+To prove $\iota$ is injective we take a nondegenerate *-representation $T\colon C^{*}(\mathcal{B}_{H})\to\mathbb{B}(X)$ with faithful integrated form $\tilde{T}\colon C^{*}(\mathcal{B}_{H})\to\mathbb{B}(X).$ By [18, Theorem 2.1], $T$ is $\mathcal{B}$-positive and it can be induced to a *-representation $S\colon\mathcal{B}\to\mathbb{B}(Y)$ by the process of [17, Ch.
 XI].
-Since $G/H$ is discrete, we can use [16, XI 14.21] to deduce that $T$ is a subrepresentation of $S|_{\mathcal{B}_{H}}$.
+Since $G/H$ is discrete, we can use [17, XI 14.21] to deduce that $T$ is a subrepresentation of $S|_{\mathcal{B}_{H}}$.
 Consequently, $\tilde{T}|_{C_{c}(\mathcal{B}_{H})}$ is a subrepresentation of $\tilde{S}\circ\iota|_{C_{c}(\mathcal{B}_{H})}$ and this implies that $\iota$ is an isometry because, for all $f\in C_{c}(\mathcal{B}_{H}),$ $\|\iota(f)\|\leq\|f\|=\|\tilde{T}_{f}\|\leq\|\tilde{S}_{\iota(f)}\|\leq\|\iota(f)\|.$ From now on we think $C^{*}(\mathcal{B}_{H})\subseteq C^{*}(\mathcal{B})$ and omit any mention of $\iota$.
 Notice that for all $f,g\in C_{c}(\mathcal{B}_{H})$ and $h\in C_{c}(\mathcal{B})$ we have $f*h*g=f*h|_{H}*g.$ Since $C_{c}(\mathcal{B}_{H})$ contains an approximate unit of $C^{*}(\mathcal{B}_{H}),$ the restriction map $E_{c}\colon C_{c}(\mathcal{B})\to C_{c}(\mathcal{B}_{H}),$ $E_{c}(h):=h|_{H},$ is a contraction with respect to the $\mathrm{C}^{*}$-norm.
-By Tomiyama’s Theorem [10, Theorem 1.5.10], the continuous extension $E\colon C^{*}(\mathcal{B})\to C^{*}(\mathcal{B}_{H})$ of $E_{c}$ is a conditional expectation.
+By Tomiyama’s Theorem [11, Theorem 1.5.10], the continuous extension $E\colon C^{*}(\mathcal{B})\to C^{*}(\mathcal{B}_{H})$ of $E_{c}$ is a conditional expectation.
 Remark 5.1.
@@
 To prove that $C_{c}(\mathcal{B}_{H})\subseteq C_{c}(\mathcal{B})$ induces and inclusion of $\mathrm{C}^{*}$-algebras $C^{*}_{\mathop{\rm r}}(\mathcal{B}_{H})\subseteq C^{*}_{\mathop{\rm r}}(\mathcal{B})$ we take a nondegenerate *-representation $T\colon\mathcal{B}\to\mathbb{B}(X)$ with $T|_{B_{e}}$ faithful.
-By [15, Corollary 2.15], there exists a faithful *-representation $\pi\colon C^{*}_{\mathop{\rm r}}(\mathcal{B})\to\mathbb{B}(L^{2}(G,X))$ such that, for all $f\in C_{c}(\mathcal{B}),$ $\pi(f)=[\lambda^{G}\tilde{\otimes}T]_{f}.$ For all $f\in C_{c}(\mathcal{B}_{H})\subseteq C_{c}(\mathcal{B}),$ $\pi(f)=[\lambda^{G}|_{H}\tilde{\otimes}T|_{\mathcal{B}_{H}}]_{f}$ and $\lambda^{G}|_{H}$ is a direct sum of copies of $\lambda^{H}$ because $G$ is the disjoint union of open cosets $Ht$.
+By [16, Corollary 2.15], there exists a faithful *-representation $\pi\colon C^{*}_{\mathop{\rm r}}(\mathcal{B})\to\mathbb{B}(L^{2}(G,X))$ such that, for all $f\in C_{c}(\mathcal{B}),$ $\pi(f)=[\lambda^{G}\tilde{\otimes}T]_{f}.$ For all $f\in C_{c}(\mathcal{B}_{H})\subseteq C_{c}(\mathcal{B}),$ $\pi(f)=[\lambda^{G}|_{H}\tilde{\otimes}T|_{\mathcal{B}_{H}}]_{f}$ and $\lambda^{G}|_{H}$ is a direct sum of copies of $\lambda^{H}$ because $G$ is the disjoint union of open cosets $Ht$.
 Then $\|f\|_{C^{*}_{\mathop{\rm r}}(\mathcal{B})}=\|[\lambda^{G}\tilde{\otimes}T]_{f}\|=\|[\lambda^{H}\tilde{\otimes}T|_{\mathcal{B}_{H}}]_{f}\|=\|f\|_{C^{*}_{\mathop{\rm r}}(\mathcal{B}_{H})},$ which implies that we may identify $C^{*}_{\mathop{\rm r}}(\mathcal{B}_{H})$ with the closure of $C_{c}(\mathcal{B}_{H})$ in $C^{*}_{\mathop{\rm r}}(\mathcal{B})$.
@@
 Assume that $H$ is a closed and normal subgroup of $G$.
-Let $\mathcal{L}=\{L^{1}(\mathcal{B}_{u})\}_{u\in G/H}$ be the partial cross-sectional bundle over $G/H$ derived from $\mathcal{B}$, and ${\mathcal{B}}_{/H}=\{\dot{B}_{u}\}_{u\in G/H}$ the bundle $\mathrm{C}^{*}$-completion of $\mathcal{L}$, see [16, VIII 6 & VIII 16.7].
+Let $\mathcal{L}=\{L^{1}(\mathcal{B}_{u})\}_{u\in G/H}$ be the partial cross-sectional bundle over $G/H$ derived from $\mathcal{B}$, and ${\mathcal{B}}_{/H}=\{\dot{B}_{u}\}_{u\in G/H}$ the bundle $\mathrm{C}^{*}$-completion of $\mathcal{L}$, see [17, VIII 6 & VIII 16.7].
 By construction, each fiber $L^{1}(\mathcal{B}_{u})$ is the $\|\ \|_{1}$-completion of $C_{c}(\mathcal{B}_{u})$, where $\mathcal{B}_{u}=\{B_{t}\}_{t\in u}$ is the reduction of $\mathcal{B}$ to the coset $u\in G/H$.
-The natural map $C_{c}(\mathcal{B})\to C_{c}(\mathcal{L}),$ $f\mapsto\tilde{f}$ with $\tilde{f}(u)=f|_{u}$, extends to a *-isomorphim $\Phi\colon L^{1}(\mathcal{B})\to L^{1}(\mathcal{L})$ (see [16, VIII 6.7]).
+The natural map $C_{c}(\mathcal{B})\to C_{c}(\mathcal{L}),$ $f\mapsto\tilde{f}$ with $\tilde{f}(u)=f|_{u}$, extends to a *-isomorphim $\Phi\colon L^{1}(\mathcal{B})\to L^{1}(\mathcal{L})$ (see [17, VIII 6.7]).
 The cross-sectional $\mathrm{C}^{*}$-algebras of $\mathcal{B}$ and $\mathcal{L}$ are the envelping $\mathrm{C}^{*}$-algebras of $L^{1}(\mathcal{B})$ and $L^{1}(\mathcal{L})$, so we may think $C^{*}(\mathcal{B})=C^{*}(\mathcal{L})$.
-In addition, the map $\rho\colon\mathcal{L}\to{\mathcal{B}}_{/H}$ of [16, VIII 17.7] is a dense embedding in the sense of [16, XI 12.6], so $\{\rho\circ f\colon f\in C_{c}(\mathcal{L})\}$ is dense in $C_{c}({\mathcal{B}}_{/H})$ with respect to the inductive limit topology (use [16, II 14.6]).
+In addition, the map $\rho\colon\mathcal{L}\to{\mathcal{B}}_{/H}$ of [17, VIII 17.7] is a dense embedding in the sense of [17, XI 12.6], so $\{\rho\circ f\colon f\in C_{c}(\mathcal{L})\}$ is dense in $C_{c}({\mathcal{B}}_{/H})$ with respect to the inductive limit topology (use [17, II 14.6]).
 The construction of the bundle $\mathrm{C}^{*}$-completion implies that all *-representations of $\mathcal{L}$ have the form $T\circ\rho$ for a *-representation $T$ of ${\mathcal{B}}_{/H}.$ Then the continuous *-homomorphism $L^{1}(\mathcal{L})\to L^{1}({\mathcal{B}}_{/H})$ mapping $f\in C_{c}(\mathcal{L})$ to $\rho\circ f$ induces an isomorphism between $C^{*}(\mathcal{L})$ and $C^{*}({\mathcal{B}}_{/H})$.
@@
 We claim that the fiber $\dot{B}_{H}$ of ${\mathcal{B}}_{/H}$ over the unit $H$ of $G/H$ is $C^{*}(\mathcal{B}_{H})$.
-For the bundle $\mathcal{L}$, [16, XI 11.6] says that $C^{*}(\mathcal{B}_{H})=\dot{B}_{H}$ if and only if every *-representation of $L^{1}(\mathcal{B}_{H})$ is $\mathcal{L}$-positive.
-Using [16, XI 12.7] we get that this last claim is equivalent to say that every *-representation of $\mathcal{B}_{H}$ is $\mathcal{B}$-positive, which is the case [17, Theorem 2.1].
+For the bundle $\mathcal{L}$, [17, XI 11.6] says that $C^{*}(\mathcal{B}_{H})=\dot{B}_{H}$ if and only if every *-representation of $L^{1}(\mathcal{B}_{H})$ is $\mathcal{L}$-positive.
+Using [17, XI 12.7] we get that this last claim is equivalent to say that every *-representation of $\mathcal{B}_{H}$ is $\mathcal{B}$-positive, which is the case [18, Theorem 2.1].
 5.4. The reduced completion of the partial cross-sectional bundle
 We want to construct a reduced version ${\mathcal{B}}^{\mathop{\rm r}}_{/H}$ of ${\mathcal{B}}_{/H}$ so that $C^{*}_{\mathop{\rm r}}(\mathcal{B})$ is isomophic to $C^{*}_{\mathop{\rm r}}({\mathcal{B}}^{\mathop{\rm r}}_{/H})$ ($H$ is normal and closed in $G$).
-To do this we take a nondegenerate *-representation $T\colon\mathcal{B}\to\mathbb{B}(X)$ with $\phi:=T|_{B_{e}}$ faithful, and let $S\colon\mathcal{L}\to\mathbb{B}(Y)$ be the *-representation corresponding to $\operatorname{Ind}_{B_{e}\uparrow\mathcal{B}}(\phi)$ via the process of [16, VIII 15.9].
-Using $S$ and the construction of [16, VIII 16.7] we can produce a Fell bundle ${\mathcal{B}}^{\mathop{\rm r}}_{/H}=\{\dot{B}^{\mathop{\rm r}}_{u}\}_{u\in G/H}$ and a dense embedding $\rho^{\mathop{\rm r}}\colon\mathcal{L}\to{\mathcal{B}}^{\mathop{\rm r}}_{/H}$ such that, for all $f\in\mathcal{L},$ $\|\rho^{\mathop{\rm r}}(f)\|=\|S_{f}\|$.
+To do this we take a nondegenerate *-representation $T\colon\mathcal{B}\to\mathbb{B}(X)$ with $\phi:=T|_{B_{e}}$ faithful, and let $S\colon\mathcal{L}\to\mathbb{B}(Y)$ be the *-representation corresponding to $\operatorname{Ind}_{B_{e}\uparrow\mathcal{B}}(\phi)$ via the process of [17, VIII 15.9].
+Using $S$ and the construction of [17, VIII 16.7] we can produce a Fell bundle ${\mathcal{B}}^{\mathop{\rm r}}_{/H}=\{\dot{B}^{\mathop{\rm r}}_{u}\}_{u\in G/H}$ and a dense embedding $\rho^{\mathop{\rm r}}\colon\mathcal{L}\to{\mathcal{B}}^{\mathop{\rm r}}_{/H}$ such that, for all $f\in\mathcal{L},$ $\|\rho^{\mathop{\rm r}}(f)\|=\|S_{f}\|$.
 By construction, for all $r\in G$; $f\in C_{c}(\mathcal{B}_{rH})$ and $\xi\in Y,$ $S_{f}\xi=\int_{N}\operatorname{Ind}_{B_{e}\uparrow\mathcal{B}}(\phi)_{f(rt)}\xi dt$.
@@
 This ultimately implies that $\tilde{S}\circ\Phi|_{C_{c}(\mathcal{B})}=\widetilde{\operatorname{Ind}}_{B_{e}\uparrow\mathcal{B}}(\phi)|_{C_{c}(\mathcal{B})}$.
-The fiber $\dot{B}^{\mathop{\rm r}}_{H}$ of ${\mathcal{B}}^{\mathop{\rm r}}_{/H}$ over the unit $H$ of $G/H$ is the completion of $L^{1}(\mathcal{B}_{H})$ with respect to the seminorm $f\mapsto\|S_{f}\|.$ Hence, to prove $\dot{B}^{\mathop{\rm r}}_{H}=C^{*}_{\mathop{\rm r}}(\mathcal{B}_{H})$ it suffices to show that for all $f\in C_{c}(\mathcal{B}_{H})$ we have $\|f\|_{C^{*}_{\mathop{\rm r}}(\mathcal{B}_{H})}=\|S_{f}\|.$ By Fell’s absorption principle [15, 17], $\operatorname{Ind}_{B_{e}\uparrow\mathcal{B}}(\phi)\equiv\operatorname{Ind}_{B_{e}\uparrow\mathcal{B}}(T|_{B_{e}})$ is weakly equivalent ($\sim$) to $\lambda^{G}\otimes T.$ The process of restricting *-representations is continuous with respect to the regional topology [16, VIII 21.20] and, by definition, it is a homeomorphism with respect to the operation of taking integrated forms.
+The fiber $\dot{B}^{\mathop{\rm r}}_{H}$ of ${\mathcal{B}}^{\mathop{\rm r}}_{/H}$ over the unit $H$ of $G/H$ is the completion of $L^{1}(\mathcal{B}_{H})$ with respect to the seminorm $f\mapsto\|S_{f}\|.$ Hence, to prove $\dot{B}^{\mathop{\rm r}}_{H}=C^{*}_{\mathop{\rm r}}(\mathcal{B}_{H})$ it suffices to show that for all $f\in C_{c}(\mathcal{B}_{H})$ we have $\|f\|_{C^{*}_{\mathop{\rm r}}(\mathcal{B}_{H})}=\|S_{f}\|.$ By Fell’s absorption principle [16, 18], $\operatorname{Ind}_{B_{e}\uparrow\mathcal{B}}(\phi)\equiv\operatorname{Ind}_{B_{e}\uparrow\mathcal{B}}(T|_{B_{e}})$ is weakly equivalent ($\sim$) to $\lambda^{G}\otimes T.$ The process of restricting *-representations is continuous with respect to the regional topology [17, VIII 21.20] and, by definition, it is a homeomorphism with respect to the operation of taking integrated forms.
 This implies
@@
 where the last equivalence holds because $\lambda^{G}|_{H}\sim\lambda^{H}$ and because the process of taking tensor products with representations of the base group is continuous with respect to the regional topology.
-By [15, Corollary 2.15], for all $f\in L^{1}(\mathcal{B}_{H})$ we have $\|S_{f}\|=\|\lambda^{H}\tilde{\otimes}(T|_{\mathcal{B}_{H}})(f)\|=\|f\|_{C^{*}_{\mathop{\rm r}}(\mathcal{B}_{H})}.$ Then $\dot{B}^{\mathop{\rm r}}_{H}=C^{*}_{\mathop{\rm r}}(\mathcal{B}_{H})$.
+By [16, Corollary 2.15], for all $f\in L^{1}(\mathcal{B}_{H})$ we have $\|S_{f}\|=\|\lambda^{H}\tilde{\otimes}(T|_{\mathcal{B}_{H}})(f)\|=\|f\|_{C^{*}_{\mathop{\rm r}}(\mathcal{B}_{H})}.$ Then $\dot{B}^{\mathop{\rm r}}_{H}=C^{*}_{\mathop{\rm r}}(\mathcal{B}_{H})$.
 Now we want to prove that $C^{*}_{\mathop{\rm r}}(\mathcal{B})=C^{*}_{\mathop{\rm r}}({\mathcal{B}}^{\mathop{\rm r}}_{/H}).$ We know $\lambda^{G}\tilde{\otimes}T\sim\operatorname{Ind}_{B_{e}\uparrow\mathcal{B}}(\phi),$ so there exists a faithful *-representation
@@
 $\operatorname{Ind}_{B_{e}\uparrow\mathcal{B}}(\phi)=\operatorname{Ind}_{\mathcal{B}_{H}\uparrow\mathcal{B}}(\operatorname{Ind}_{B_{e}\uparrow\mathcal{B}_{H}}(T|_{B_{e}})).$
-The integrated form of $\operatorname{Ind}_{B_{e}\uparrow\mathcal{B}_{H}}(T|_{B_{e}})\sim\lambda^{H}\otimes(T|_{\mathcal{B}_{H}})$ is a *-representation $R$ of $L^{1}(\mathcal{B}_{H})$ for which there exists a faithful *-representation $R^{\prime}$ of $\dot{B}^{\mathop{\rm r}}_{H}=C^{*}_{\mathop{\rm r}}(\mathcal{B}_{H})$ such that $R^{\prime}\circ\rho|_{L^{1}(\mathcal{B}_{H})}=R,$ see [15, Corollary 2.15].
-By [16, XI 12.6-12.7], $S=\operatorname{Ind}_{L^{1}(\mathcal{B}_{H})\uparrow\mathcal{L}}(R)$ and $\operatorname{Ind}_{\dot{B}^{\mathop{\rm r}}_{H}\uparrow{\mathcal{B}}^{\mathop{\rm r}}_{/H}}(R^{\prime})\circ\rho^{\mathop{\rm r}}=\operatorname{Ind}_{L^{1}(\mathcal{B}_{H})\uparrow\mathcal{L}}(R).$ From Proposition 3.1 of [17] we get the existence of a faithful *-representation
+The integrated form of $\operatorname{Ind}_{B_{e}\uparrow\mathcal{B}_{H}}(T|_{B_{e}})\sim\lambda^{H}\otimes(T|_{\mathcal{B}_{H}})$ is a *-representation $R$ of $L^{1}(\mathcal{B}_{H})$ for which there exists a faithful *-representation $R^{\prime}$ of $\dot{B}^{\mathop{\rm r}}_{H}=C^{*}_{\mathop{\rm r}}(\mathcal{B}_{H})$ such that $R^{\prime}\circ\rho|_{L^{1}(\mathcal{B}_{H})}=R,$ see [16, Corollary 2.15].
+By [17, XI 12.6-12.7], $S=\operatorname{Ind}_{L^{1}(\mathcal{B}_{H})\uparrow\mathcal{L}}(R)$ and $\operatorname{Ind}_{\dot{B}^{\mathop{\rm r}}_{H}\uparrow{\mathcal{B}}^{\mathop{\rm r}}_{/H}}(R^{\prime})\circ\rho^{\mathop{\rm r}}=\operatorname{Ind}_{L^{1}(\mathcal{B}_{H})\uparrow\mathcal{L}}(R).$ From Proposition 3.1 of [18] we get the existence of a faithful *-representation
 $\nu\colon C^{*}_{\mathop{\rm r}}({\mathcal{B}}^{\mathop{\rm r}}_{/H})\to\overline{\widetilde{\operatorname{Ind}}_{\dot{B}^{\mathop{\rm r}}_{H}\uparrow{\mathcal{B}}^{\mathop{\rm r}}_{/H}}(R)(C_{c}({\mathcal{B}}^{\mathop{\rm r}}_{/H}))}$
@@
 The function $\Omega^{\mathop{\rm r}}\colon C_{c}(\mathcal{B})\to C_{c}({\mathcal{B}}^{\mathop{\rm r}}_{/H}),$ $f\mapsto\rho^{\mathop{\rm r}}\circ\tilde{f}$, is a *-homomorphism because $\nu\circ\Omega^{\mathop{\rm r}}=\mu$.
-Using [16, II 14.6] one can show that the image of $\Omega^{\mathop{\rm r}}$ is dense with respect to the inductive limit topology, so $\Omega^{\mathop{\rm r}}(C_{c}(\mathcal{B}))$ is dense in $C^{*}({\mathcal{B}}^{\mathop{\rm r}}_{/H})$ and $\Omega^{\mathop{\rm r}}$ extends to an isomorphism
+Using [17, II 14.6] one can show that the image of $\Omega^{\mathop{\rm r}}$ is dense with respect to the inductive limit topology, so $\Omega^{\mathop{\rm r}}(C_{c}(\mathcal{B}))$ is dense in $C^{*}({\mathcal{B}}^{\mathop{\rm r}}_{/H})$ and $\Omega^{\mathop{\rm r}}$ extends to an isomorphism
 (5.5) $\Omega^{\mathop{\rm r}}\colon C^{*}_{\mathop{\rm r}}(\mathcal{B})\to C^{*}_{\mathop{\rm r}}({\mathcal{B}}^{\mathop{\rm r}}_{/H}).$
@@
 Proof.
-The first claim is [17, Corollary 5.6].
+The first claim is [18, Corollary 5.6].
 Since $(\mathcal{B}_{N})_{H}=\mathcal{B}_{H}$ and we have conditional expectations, $E\colon C^{*}(\mathcal{B})\to C^{*}(\mathcal{B}_{N})$ and $E_{\mathop{\rm r}}\colon C^{*}_{\mathop{\rm r}}(\mathcal{B})\to C^{*}_{\mathop{\rm r}}(\mathcal{B}_{N})$, we may assume that $N=G$ (i.e.
@@
 Take a nondegenerate *-representation $T\colon C^{*}(\mathcal{B})\to\mathbb{B}(X)$ with faithful integrated form.
-Via the isomorphism $L^{1}(\mathcal{L})\cong L^{1}(\mathcal{B})$, $T$ induces a *-representation $T^{1}\colon\mathcal{L}\to\mathbb{B}(X)$, see [16, VIII 15.9], and $T^{1}$ determines a unique *-representation $\dot{T}\colon{\mathcal{B}}_{/H}\to\mathbb{B}(X)$ such that $\dot{T}\circ\rho=T^{1}$.
+Via the isomorphism $L^{1}(\mathcal{L})\cong L^{1}(\mathcal{B})$, $T$ induces a *-representation $T^{1}\colon\mathcal{L}\to\mathbb{B}(X)$, see [17, VIII 15.9], and $T^{1}$ determines a unique *-representation $\dot{T}\colon{\mathcal{B}}_{/H}\to\mathbb{B}(X)$ such that $\dot{T}\circ\rho=T^{1}$.
 Moreover, the integrated forms of $\dot{T}$ and $T$ are the same up to the isomorphism of (5.4).
@@
 Take a nondegenerate dilation of $T\circ\Phi$, $S\colon\mathcal{B}\to\mathbb{B}(Y)$, and $V\in\mathbb{B}(X,Y)$ such that $T\circ\Phi(b)=V^{*}S_{b}V$ and $\|V\|\leq 1$.
-Using the explicit descriptions $T^{1}$ and $S^{1}$ of [16, VIII 15.9], a straightforward computation reveals that for all $f\in\uplus_{u\in G/H}C_{c}(\mathcal{B}_{u})$, $T^{1}_{\Phi\circ f}=V^{*}S^{1}_{f}V=V^{*}\dot{S}_{\rho(f)}V$ and it follows that $\|\rho(\Phi\circ f)\|\leq\|\rho(f)\|$.
+Using the explicit descriptions $T^{1}$ and $S^{1}$ of [17, VIII 15.9], a straightforward computation reveals that for all $f\in\uplus_{u\in G/H}C_{c}(\mathcal{B}_{u})$, $T^{1}_{\Phi\circ f}=V^{*}S^{1}_{f}V=V^{*}\dot{S}_{\rho(f)}V$ and it follows that $\|\rho(\Phi\circ f)\|\leq\|\rho(f)\|$.
 Since $\rho(C_{c}(\mathcal{B}))$ is dense in $\dot{B}_{u}$, there exists a unique family of linear contractions $\{\dot{\Phi}_{u}\colon\dot{B}_{u}\to\dot{B}_{u}\}_{u\in G/H}$ whose union ${\Phi}_{/H}\colon{\mathcal{B}}_{/H}\to{\mathcal{B}}_{/H}$ satisfies ${\Phi}_{/H}\circ\rho(f)=\rho(\Phi\circ f)$, for all $f\in\uplus_{u\in G/H}C_{c}(\mathcal{B}_{u})$.
 For all $f\in C_{c}(\mathcal{B})$ we have ${\Phi}_{/H}\circ(\rho\circ\tilde{f})(u)={\Phi}_{/H}\circ(\rho(f|_{u}))=\rho((\Phi\circ f)|_{u})=\widetilde{\Phi\circ f}(u)$.
-Considering the family $\Gamma:=\{\rho\circ\tilde{f}\colon f\in C_{c}(\mathcal{B})\}$ and using [16, II 13.16] we get that ${\Phi}_{/H}$ is continuous and it is ccp because $T\circ{\Phi}_{/H}(z)=V^{*}\dot{S}_{z}V$.
+Considering the family $\Gamma:=\{\rho\circ\tilde{f}\colon f\in C_{c}(\mathcal{B})\}$ and using [17, II 13.16] we get that ${\Phi}_{/H}$ is continuous and it is ccp because $T\circ{\Phi}_{/H}(z)=V^{*}\dot{S}_{z}V$.
 Notice that ${\mathop{\rm supp}}({\Phi}_{/H})$ is contained in the projection of ${\mathop{\rm supp}}(\Phi)$ into $G/H$.
@@
 We know that $(\operatorname{id}\bar{\otimes}\pi)(P\bar{\otimes}M)\subseteq\mathbb{B}(Z)\bar{\otimes}N$.
-Combining Lemma A.4 with [22, Corollary 5.9] we obtain $(\operatorname{id}\bar{\otimes}\pi)(P\bar{\otimes}M)\subseteq(P^{\prime}\bar{\otimes}\mathbb{C}1_{Y})^{\prime}=P\bar{\otimes}\mathbb{B}(Y)$.
+Combining Lemma A.4 with [24, Corollary 5.9] we obtain $(\operatorname{id}\bar{\otimes}\pi)(P\bar{\otimes}M)\subseteq(P^{\prime}\bar{\otimes}\mathbb{C}1_{Y})^{\prime}=P\bar{\otimes}\mathbb{B}(Y)$.
 Now that we know
 $(\operatorname{id}\bar{\otimes}\pi)(P\bar{\otimes}M)\subseteq(P\bar{\otimes}\mathbb{B}(Y))\cap(\mathbb{B}(Z)\bar{\otimes}N),$
-we can use [22, Corollary 5.10] to get $(\operatorname{id}\bar{\otimes}\pi)(P\bar{\otimes}M)\subseteq P\bar{\otimes}N$.
+we can use [24, Corollary 5.10] to get $(\operatorname{id}\bar{\otimes}\pi)(P\bar{\otimes}M)\subseteq P\bar{\otimes}N$.
 Now assume that the unit $p$ of $P$ is not $1_{Z}$ and let $P\to\tilde{P}\subseteq\mathbb{B}(pZ)$ be given by $m\mapsto\tilde{m}:=m|_{pZ}$.
@@
 All we have done so far can be repeated flipping the tensors, we mean using $M\bar{\otimes}\mathbb{B}(Z)$ instead of $\mathbb{B}(Z)\bar{\otimes}M$, so we get a ccp map $\pi\bar{\otimes}\operatorname{id}\colon M\bar{\otimes}\mathbb{B}(Z)\to N\bar{\otimes}\mathbb{B}(Z)$.
-At this point we have all the necessary ingredients to prove equivariant versions of some results by Nagisa and Tomiyama [20, Theorem 3], [23, Theorem 4].
+At this point we have all the necessary ingredients to prove equivariant versions of some results by Nagisa and Tomiyama [22, Theorem 3], [25, Theorem 4].
 Theorem A.7.